### v1

# natural language processing; greenwashing detection; corporate climate-related greenwashing; climate pretrained models; climate topic detection; green claim detection; climate QA; climate-related deceptive techniques; enviromental performance prediction copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language processing

Abstract.
Greenwashing is an effort to mislead the public about the carbon transition of an entity, such as a state or company. We provide a comprehensive survey of the scientific literature addressing natural language processing methods to identify potentially misleading climate-related corporate communications, indicative of greenwashing. We break the detection of greenwashing into intermediate tasks, and review the state-of-the-art approaches for each of them. We discuss datasets, methods, and results, as well as limitations and open challenges. We also provide an overview of how far the field has come as a whole, and point out future research directions.
natural language processing; greenwashing detection; corporate climate-related greenwashing; climate pretrained models; climate topic detection; green claim detection; climate QA; climate-related deceptive techniques; enviromental performance prediction
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language processing
1. Introduction
Since the first report of the Intergovernmental Panel on Climate Change in 1990  (Change, 1990), world leaders and citizens have increasingly recognized the significant adverse impacts of human activities on the environment, particularly on climate change (Peter Jackson, 2007). This increased awareness has translated into guidelines, laws, and investments, such as the European Green Deal (Publications Office of the European Union, 2019) or the Inflation Reduction Act in the US (The White House, 2023). Many companies have used the financial incentives offered by states, and the guidelines and legislation to make significant steps towards sustainability (Time, 2024). At the same time, this growing attention also generated an advertising opportunity for companies that aim to promote themselves as environmentally aware and responsible. Indeed, some companies have been found to deliberately manipulate their data and statistics to appear more environment-friendly. The Diesel Scandal around the Volkswagen car company is a prominent example (US Environmental Protection Agency, 2024). However, such cases are not the norm. More commonly, companies avoid outright data manipulation but present themselves in a misleadingly positive light regarding their environmental impact – a practice called greenwashing.
In addition to their broader communication strategies, most companies disseminate information about their environmental initiatives through dedicated sustainability 10-K and annual reports. In addition, corporations also leverage various media channels and traditional communication platforms to convey their environmental credentials. As a result, the corpus of documents used in textual analysis is diverse and heterogeneous.
In such documents, greenwashing can appear in multiple forms. The Green Claims Directive of the European Union (US Environmental Protection Agency, 2023) identifies four types of greenwashing:
  1. (1)
    using unreliable sustainability labels (e.g., H&M and Decathlon have used labels such as eco design and conscious and were prompted by the Netherlands Authority for Consumers and Markets to stop the practice in 2022 as the retailers did not clearly describe why the products received those labels (Netherlands Authority for Consumers and Markets (2022), ACM))
  2. (2)
    presenting legal requirements for the product as its distinctive features (e.g., McDonald’s advertises its reduction of plastic waste by using reusable cutlery (McDonald’s France, 2024), while this is a legal obligation since January 2023 in France (LegiFrance, 2020))
  3. (3)
    making green claims about the entire product when the claim applies only to a part/aspect of the product (e.g., Kohl’s and Walmart were fined by the U.S. Federal Trade Commission for the misleading claim that their products were made of bamboo, a sustainable material, when they were made out of rayon, which is derived from bamboo via a highly toxic process that removes any environmental benefits of bamboo (Lesley Fair, Federal Trade Commission, 2022)), and finally
  4. (4)
    making environmental claims for which the company cannot provide evidence (e.g., Keurig was fined by Canada’s Competition Bureau for misleading claims about the recycling process of its single-use coffee pods (Competition Bureau Canada, 2022)).
Greenwashing is a widely observed practice rather than a marginal occurrence: the Australian Competition & Consumer Commission performed a sweep analysis on 247 businesses in October 2022 and found that $57\%$ of them made misleading claims about their environmental performance (Australian Competition and Consumer Commission, 2023). The International Consumer Protection Enforcement Network, too, found in November 2020 that $40\%$ of around 500 websites of businesses covering different sectors contain misleading green claims (UK Competition and Markets Authority, 2021).
Addressing greenwashing is not only in the broader public interest but also in the company’s best interest, as the practice is often subject to legal sanctions prohibiting misleading or deceptive conduct, with businesses facing fines or lawsuits as a consequence (Mary Foley, Forbes, 2024). In addition, political awareness and public awareness of the issue via traditional and social media (Thomas Wagner, Bonpote, 2021) can harm the reputation of a company involved in the practice. These factors make greenwashing a risk factor that can affect investments and that should be monitored at the company and portfolio levels.
The question thus arises to what extent greenwashing can be detected automatically. Such automated detection could help investors, journalists, prosecutors, and activists spot potentially illicit behavior, but it could also help companies avoid unintentionally misleading communication. In this paper, we provide a comprehensive survey of natural language processing (NLP) methods to identify potentially misleading climate-related corporate communication that can be indicative of greenwashing. We have identified 61 works that propose ways to check one or more aspects of greenwashing in text. We cover the intermediary steps involved in a natural language processing (NLP) analysis of the environmental disclosure of a company, ranging from detecting environmental or climate-related text to greenwashing detection. We structure these works in a unified categorization and discuss tasks, datasets, methods, and results for each of them. Unlike other surveys on the topic (Moodaley and Telukdarie, 2023b; Bernini et al., 2023), which compile and discuss findings reported in the scientific literature, our approach dives into the actual methodologies of these works.
Our main contributions and findings are:
  • We provide the first survey on NLP-based detection of greenwashing that describes and systemizes the methods used in the literature. Our survey focuses on tasks and datasets but also contextualizes the performance of the models and provides future research directions.
  • We find no datasets of positive and negative examples of greenwashing. Hence, it is not possible to train or evaluate models on the task of identifying greenwashing.
  • While there are no datasets explicitly labeled for greenwashing, many works have studied intermediary tasks, such as detecting misleading claims, vague language, or environmental reporting inconsistencies. Researchers have used these characteristics to develop theoretical definitions of greenwashing and proposed measurable indicators to identify potential cases. Although not a direct substitute for labeled greenwashing datasets, these indicators provide tools for analyzing corporate disclosures and assessing the likelihood of greenwashing.
  • Open challenges remain: While many models have been proposed for greenwashing-related tasks, it is difficult to compare approaches and solutions as many studies do not report measures of uncertainty (e.g., confidence interval, multiple runs, hyper-parameters tuning), nor do they report simple baselines that help assess the difficulty of datasets such as random baseline or TF-IDF baselines. Even the intermediary tasks thus remain active areas of research.
Our survey is structured as follows. We start with a discussion of the related work on greenwashing surveys (Section 2), followed by the legal context around climate actions and the definition of greenwashing (Section 3). We then survey intermediate tasks that have been studied in the literature (Section 4): identifying climate-related statements; identifying topics aligned with frameworks such as the Task Force on Climate-related Financial Disclosures (TCFD) or Environmental, Social, and Governance (ESG) factors; assessing the presence of green claims; and examining sentiment, argumentation quality, deceptive practices, and stance to discern the tone and authenticity of a message. We then review approaches that aim to identify greenwashing as a whole (Section 5). We conclude by discussing challenges and opportunities in greenwashing detection (Sections 6 and 7). We note that we have an Appendix section that will be part of the supplemental online-only material.
2. Related Work
Moodaley and Telukdarie (2023b) analyze the literature on greenwashing, sustainability reporting, and the intersection with research in Artificial Intelligence (AI) and Machine Learning (ML). The authors provide an overview of how the literature has changed over time by examining the number of publications and the different keywords covered by such publications. Unlike our work, this survey does not describe the actual approaches of the publications. Bernini et al. (2023) review the definitions of greenwashing in business, management, and accounting. They conclude that there is no agreed-upon definition of greenwashing, except that it is defined as a gap between the company’s actions and its disclosure. In contrast to our study, the review provides only a brief overview of the techniques for detecting greenwashing. While our review focuses on the semantic and linguistic aspects of greenwashing identification, other works focus on quantitative approaches that can serve the same objective (Dao et al., 2024).
Other works are less directly related to our survey. For example, a growing community is focused on using AI to tackle climate change [^1] See for example https://www.climatechange.ai/. Rolnick et al. (2022) review how machine learning has been used and could be used to tackle climate change. They explore different sectors of the economy, such as electricity, transportation, industry, and agriculture, among others. The survey discusses climate investments and the effects of climate change on finance, but not greenwashing. Alonso et al. (2023) provide a survey on ML approaches for climate finance, also known in the literature as green finance or carbon finance. Climate finance is a subfield in economics dedicated to providing tools from financial economics to tackle climate change. Finally, several works study the impact of AI on climate change (Hershcovich et al., 2022; Kaack et al., 2022; Rohde et al., 2023; Verdecchia et al., 2023). Our work, in contrast, surveys approaches that use AI to detect greenwashing.
3. Preliminaries
3.1. Legal Context
In this section, we provide the legal context for the definition of greenwashing. We first present international and national laws, followed by how they are applied in the corporate context.
International and Nation Level Legislation and Guidelines on Climate Change Mitigation
The first international summit on the effect of humans on the environment was the United Nations Conference on the Human Environment, held in 1972 in Sweden (United Nations, European Commission, 2024). The Intergovernmental Panel on Climate Change (IPCC) was created in 1988 and released its first report in 1990  (Change, 1990). It was only in 1992 that the first treaty between countries was signed, the United Nations Framework Convention on Climate Change (UNFCCC) (United Nations, European Commission, 2024). The treaty only encouraged countries to reduce their emissions, and in 1997, the Kyoto Protocol set commitments for the countries to follow (United Nations Framework Convention on Climate Change, 2011). The 36 countries that participated in the Protocol reduced their emissions in 2008–2012 by a large margin in respect to the levels in 1990; however, the global emissions increased by $32\%$ in the same period. Since the creation of UNFCCC, the nations, i.e., the parties, have met yearly at the “Conference of the Parties” (COP).
The three main directions of climate change policy are reducing greenhouse emissions, promoting renewable energy, and improving energy efficiency. One major milestone was the introduction in 2005 of the European Union Emissions Trading System, the world’s first large-scale emissions trading scheme. The trading system limits the amount of CO2 emitted by European industries and covers $46\%$ of the EU’s CO2 emissions. It is estimated that the trading system reduced CO2 emissions in the EU by $3.8\%$ between 2008 and 2016 (Bayer and Aklin, 2020). At COP21, in 2015, 194 nations plus the European Union signed the Paris Agreement, a treaty by which the nations commit to keep the rise in global surface temperature below 2 °C (3.6 °F) above pre-industrial levels. To achieve this goal, each country sends a national climate action plan every five years, and the parties assess the collective progress made towards achieving the climate goals. The first such assessment took place at COP28, where it was established that an important direction to achieve the long-term goals of countries was to transition away from fossil fuels to renewable energy. In 2021, the European Climate Law was adopted: the EU commits to reducing its emissions by at least $55\%$ by 2030 with respect to the 1990 levels and becoming climate neutral by 2050 (Council of the EU Press release, 2021). Similarly, countries such as Canada (Government of Canada, 2024), Taiwan (The Republic of China Ministry of Environment(2023), Taiwan), South Korea (The Korean Law Information Center, 2024), and Australia (Government of New South Wales, 2023), among others, aim to achieve carbon neutrality by 2050 and have passed laws setting this goal.
Corporate Level Laws and Guidelines
Laws and regulations at international and national levels have repercussions on companies that must comply or risk fines. However, assessing if a company has taken the necessary steps is not trivial, given that some laws refer to how a company will operate in the future, for example, by polluting less. Even without laws, investors can be concerned about environmental or social issues; hence, companies have used voluntary disclosures for a long time.
One of the first organizations to provide standards for reporting on climate change or social aspects is the Global Reporting Initiative (GRI), with its first guidelines published in 2000 (GRI, 2023). In 2015, the Financial Stability Board and the Group of 20 created the Task Force for Climate-related Financial Disclosure (TCFD) guidelines on disclosure in response to shortcomings of COP21, in particular the lack of standards climate-related disclosure (Wikipedia contributors, 2024). In 2021, the International Sustainability Standards Board (ISSB) was created to establish standards for climate-related disclosure, and starting in 2024, the standard released by this board will be applied worldwide (IFRS Foundation, 2024). ISSB standards were aligned with GRI disclosure standards to make them complementary and interoperable (Global Reporting Initiative, 2023), while the ISSB standards are expected to take over TCFD (IFRS Foundation, 2024). Differently from the voluntary reporting standards of GRI, TCFD, and ISSB, the Corporate sustainability reporting law in the EU required the creation of the European Sustainability Reporting Standards (ESRS), which are mandatory for companies subject to EU law (Directorate-General for Financial Stability, Financial Services and Capital Markets Union, 2023). ESRS has high interoperability with GRI and ISSB. GRI, TCFD, ESRS and ISSB fall under the umbrella term of environmental, social, and governance (ESG) guidelines for company disclosure. The European Union created its first law obliging companies to provide non-financial disclosure reporting in 2014, the Non-Financial Reporting Directive, which focused disclosure on environmental and social aspects. In 2023, the EU expanded this legislation via the Corporate Sustainability Reporting Law. Some European countries anticipated this legislation with their own, for example, the 2001 New Economic Regulations Act in France. Switzerland, which is not part of the EU, has imposed a mandatory TCFD disclosure for large public companies, banks, and insurance companies starting from January 2024 (The Swiss Federal Council, 2022), and similar laws exist in New Zealand (New Zealand Ministry of Environment, 2023). While not compulsory, initiatives like the Carbon Disclosure Project (CDP) have played a pivotal role by standardizing responses related to the climate disclosure of a company via structured questionnaires, thereby facilitating more systematic and comparable reporting.
3.2. Definition of Greenwashing
A widely cited and comprehensive definition, synthesizing those commonly found in the literature, is provided by the Oxford English Dictionary (OED, 2023).
Definition 0.
Greenwashing: To mislead the public (or to counter public or media concerns) by falsely representing a person, company, product, etc., as environmentally responsible.
While individuals, companies, or countries can all engage in greenwashing, we will focus on climate-related greenwashing by companies in this survey, with the following definition:
Definition 0.
Corporate climate-related greenwashing: To mislead the public into falsely representing the effort made by a company to achieve its carbon transition.
As climate-related disclosures face increasing regulation and greenwashing poses significant risks to a company’s reputation, some companies adopt a strategy of silence, avoiding discussions about their environmental impact. This deliberate lack of communication is known as greenhushing (Letzing, 2022). These definitions imply that greenwashing is a deliberate act; however, in many cases, it results from an error or miscommunication by companies genuinely trying to showcase their sustainability efforts, and that are trying to best follow disclosure standards.
As shown by Definitions 3.1 and 3.2, greenwashing is not defined by easily identifiable properties but as a general concept. Since the concept is so unspecific, researchers focused on components indicative of potentially misleading communications but easier to define. Because of this, in this review we are mentioning “greenwashing” explicitly, but also paraphrasing it as “misleading communications”, “misrepresentation of the company’s environmental impact/stance/performance”, or mentioning only components of it such as “cheap talk”, “selective disclosure/transparency”, “deceptive techniques”, “biased narrative”. They should all be understood in the context of climate-related misleading communications, as components associated or indicative of potential greenwashing even if they are not synonymous.
4. Intermediary tasks
4.1. Pretraining Models on Climate-Related Text
The first step in applying a language model for a given task is typically the pretraining process, which involves training the model on relevant corpora. Although general-purpose models such as BERT (Devlin et al., 2019) and LLaMA (Touvron et al., 2023) have demonstrated strong performance in various tasks, they are inherently limited by the knowledge and vocabulary present in their training corpora. This limitation can lead to suboptimal results when these models are applied to highly specialized domains with unique terminology and concepts. Hence, domain-specific models are usually trained on targeted datasets to improve their performance in niche areas.
Task Description
A language model is pretrained on climate-related text, usually on the classical tasks of masked token prediction and next sentence prediction. The goal is to produce a domain-specific pre-trained language model that will be subsequently fine-tuned to specific tasks. There are two approaches: (1) training a model with a domain-specific corpus from scratch, or (2) further-training [^2] Further-training refers to the process of training a pretrained model on its original task (e.g., next-token prediction or masked-token prediction) using additional domain-specific data to specialize its knowledge for that domain. a generalist pretrained model on a domain-specific corpus.
Annotated Datasets
There have been several efforts to create a climate-related corpus. Webersinke et al. (2022) gathered climate-related news articles, research abstracts, and corporate reports. Vaghefi et al. (2022) introduced Deep Climate Change, a dataset composed of abstracts of articles from climate scientists, and built a corpus specific to climate research texts. Schimanski et al. (2023b) introduced a dataset that focuses on text related to Environment Social and Governance (ESG). Similarly, Mehra et al. (2022) built a dataset using text from the Knowledge Hub of Accounting for Sustainability for an ESG domain-specific corpus. More recently, Thulke et al. (2024) proposed a dataset comprising news, publications (abstracts and articles), books, patents, the English Wikipedia, policy and finance-related texts, Environmental Protection Agency documents, and ESG, and IPCC reports. Mullappilly et al. (2023) proposed a climate-specific multilingual dataset. Yu et al. (2024) published a pretraining dataset constructed with annual and sustainable reports from EU banks. Unfortunately, to the best of our knowledge, only Yu et al. (2024)’s ClimateBUG-data dataset is publicly available.
Supervised Solutions
Based on the above datasets, the models ClimateBERT, ClimateGPT-2, climateBUG-LM and EnvironmentalBERT, ESGBERT were proposed (Webersinke et al., 2022; Vaghefi et al., 2022; Yu et al., 2024; Schimanski et al., 2023b; Mehra et al., 2022). More recently, generative domain-specific models such as Llama-2 for ClimateGPT (Thulke et al., 2024) or Vicuna for Arabic Mini-ClimateGPT (Mullappilly et al., 2023) have also been proposed.
Performance of Models
The domain-specific models drastically improve the performance on domain-specific masked-language modeling (Webersinke et al., 2022; Yu et al., 2024) and next token prediction (Vaghefi et al., 2022). They were also evaluated on domain-specific downstream tasks. These tasks are either based on pre-existing datasets such as ClimateFEVER (Leippold and Diggelmann, 2020) or introduced by the authors, as in ClimateBERT’s climate detection (Webersinke et al., 2022). At this stage, we evaluate whether fine-tuning enhances downstream performance, with a detailed analysis of the tasks presented in subsequent sections.
Cl.BERT DisRoBERTa NB
Climatext (Spokoyny et al., 2023; Leippold and Varini, 2020) 85.14 86.06 83.39
Climate Detection (Webersinke et al., 2022) 99.1±$\pm 1$ 98.6±$\pm 0.8$
Sentiment (Webersinke et al., 2022) 83.8±$\pm 3.6$ 82.5±$\pm 4.6$
Net-zero/Reduction (Schimanski et al., 2023a) 96.2±$\pm 0.4$ 94.4±$\pm 0.6$
Cl.BERT DisRoBERTa TF-IDF
TCFD classification (Sampson et al., 2022) 0.852 0.819 0.867
Cl.BERT DisRoBERTa SVM
SciDCC (Spokoyny et al., 2023; Mishra and Mittal, 2021) 52.97 51.13 48.02
ClimaTOPIC (Spokoyny et al., 2023) 64.24 63.61 58.34
Cl.FEVER (claim) (Xiang and Fujii, 2023; Leippold and Diggelmann, 2020) 76.8 72
Cl.FEVER (claim) (Webersinke et al., 2022; Leippold and Diggelmann, 2020) 75.7±$\pm 4.4$ 74.8±$\pm 3.6$
Cl.FEVER (evid.) (Spokoyny et al., 2023; Leippold and Diggelmann, 2020) 61.54 61.54
Cl.Stance (Spokoyny et al., 2023; Vaid et al., 2022) 52.84 52.51 42.92
ClimateEng (Spokoyny et al., 2023; Vaid et al., 2022) 71.83 72.33 51.81
ClimaINS (Spokoyny et al., 2023) 84.80 84.38 86.00
ClimaBENCH (Spokoyny et al., 2023) 69.44 69.27
Nature (Schimanski et al., 2024a) 94.11 94.03
Cl.BERT SVM+ELMo SVM+BoW
Commitment&Actions (Bingler et al., 2023) 81 79 76
Specificity (Bingler et al., 2023) 77 76 75
Cl.BERT DisBERT SVM
Env. Claim (Stammbach et al., 2023) 83.8 83.7 70.9
Cl.BERT DisBERT LSTM
DARE sentiment (Xiang and Fujii, 2023) 87.4 89.9 88.2
(a) F1-scores of climateBERT when compared to a similarly sized models. ROC-AUC for TCFD classification.
RoBERTa EnvRoBERTa DisRoBERTa EnvDisRoB.
Environment (Schimanski et al., 2023b) 92.35±$\pm 2.29$ 93.19±$\pm 1.65$ 90.97±$\pm 2$ $92.35\pm 1.65$
Social (Schimanski et al., 2023b) 89.87±$\pm 1.35$ 91.90±$\pm 1.79$ 90.59±$\pm 1.03$ $91.24\pm 1.86$
Governance (Schimanski et al., 2023b) 77.03±$\pm 1.82$ 78.48±$\pm 2.62$ 76.65±$\pm 2.39$ 78.86±$\pm 1.59$
EnvDisRoB. ClimateBERT RoBERTa DisRoBERTa
Water (Schimanski et al., 2024a) 94.47±$\pm 1.37$ 95.10±$\pm 1.13$ 94.55±$\pm 0.86$ $94.98\pm 1.16$
Forest (Schimanski et al., 2024a) 95.37±$\pm 0.92$ 95.34±$\pm 0.94$ 94.78±$\pm 0.48$ 95.29±$\pm 0.65$
Biodiversity (Schimanski et al., 2024a) 92.76±$\pm 1.01$ 92.49±$\pm 1.03$ 92.46±$\pm 1.54$ 92.29±$\pm 1.23$
Nature (Schimanski et al., 2024a) 94.19±$\pm 0.81$ 93.50±$\pm 0.64$ 93.97±$\pm 0.26$ 93.55±$\pm 0.72$
(b) F1-scores of multiple models compared to EnvironmentalBERT (based on RoBERTa and distilRoBERTa).
GPT-2 climateGPT-2
ClimateFEVER (Vaghefi et al., 2022) 62 72
LLama-2-Chat (7B) ClimateGPT (7B)
Multiple Benchmarks (Thulke et al., 2024) 71.4 77.1
(c) F1-scores of ClimateGPT-2 and custom performances (average Accuracy and F1-score over multiple benchmarks) of ClimateGPT compared to similarly sized models.
Table 1. Performance of domain-specific models on domain-specific tasks. The figure displayed in the tables are the values reported by authors in the corresponding studies; with the following abbreviations: CL for Climate, Dis for Distil, EnvDisRob for EnvDistilRoBERTa, Evid. for evidences. Each row reports the performances of models fine-tuned in the same experimental setting. Detailed performances are reported in Appendix C.
As shown in Table 1(c), both ClimateGPT-2 and ClimateGPT significantly outperform the baseline model in their designated tasks. These models benefit from additional training on climate-related data, highlighting the advantages of domain adaptation. However, when trained from scratch, Thulke et al. (2024) report substantially lower performance, underscoring the importance of large-scale pretraining data. For smaller models (climateBERT, climateBUG-LM and EnvironmentalBERT), the domain-adaptation also enhanced the performance on domain-specific tasks, though the significance of this improvement is less pronounced. As detailed in Table 1(a), out of the 15 evaluations of fine-tuned domain-specific climateBERT (Webersinke et al., 2022) on climate-related tasks compared to the fine-tuned base model (distilRoBERTa), 4 reported an improvement compared to the base model (¿1%), 8 reported a marginal improvement (¡1%) and 3 indicate no measurable difference (¡=0%). When compared to DistilBERT in the last two tasks from Table 1(a), climateBERT exhibits a slight improvement in one case, while performing comparably in the other. Additionally, climateBERT achieves performance similar to the simple baselines in certain instances (Table 1(a)). For example, in Specificity Classification (Bingler et al., 2023), climateBERT outperforms an SVM with bag-of-word by 2%, and on the TCFD classification (Sampson et al., 2022) climateBERT performs slightly below the TF-IDF baseline model. Likewise, climateBUG-LM (Yu et al., 2024) demonstrates a slight improvement of less than 1% over BERT and FinBERT on a climate-related detection task. As shown in Table 1(c), the domain-adapted models shows a small improvement in the performance on the Environmental task, by 1.38% for distilRoBERTa and 0.84% for RoBERTa.
Insights
Domain-specific language models, particularly those developed for environmental and climate-related texts, offer an interesting avenue for improving model performance through enhanced knowledge within specialized domains. Theoretically, such models should outperform general models by leveraging their tailored vocabulary and contextual understanding. However, our findings suggest that the improvements provided by the domain-specific models on the proposed domain-specific downstream tasks are limited. Indeed, we found that fine-tuning general models on domain-specific datasets consistently yields competitive performance. However, these findings do not undermine the significance of existing works that have been fundamental in demonstrating the potential of domain-specific models, particularly in generating domain-specific text (MLM and next token prediction). Moreover, ClimateGPT (Thulke et al., 2024) demonstrated significant improvement for the domain-adapted Llama-2, showing that domain-adaptation is highly relevant. Future research could focus on designing tasks where a lack of domain understanding—whether in terms of vocabulary or knowledge—significantly hinders performance.
4.2. Climate-Related Topic Detection
The first step toward detecting climate-related greenwashing is identifying text addressing climate-related topics.
Task Description
Given an input sentence or a paragraph, output a binary label, “climate-related” or “not climate-related”.
Dataset Input Labels Positive Label Details
ClimateBug-data (Yu et al., 2024) sentences from Banks’ reports relevant/ irrelevant Climate change and sustainability (including ESG, SDGs related to the environment, recycling and more)
ClimateBERT’s climate detection (Bingler et al., 2023) paragraphs from reports 1/0 Climate policy, climate change or an environmental topic
Climatext (Wikipedia, 10-K, claims) (Leippold and Varini, 2020) sentences from Wikipedia, 10-Ks or web scraping 1/0 Directly related to climate-change
Climatext (wiki-doc) (Leippold and Varini, 2020) sentences from a Wikipedia page 1/0 Extracted from a Wikipedia page related to climate-change
Sustainable signals’s reviews (Lin et al., 2023b) online product reviews (user comments) relevant/ irrelevant Contains terms related to sustainability
Table 2. Label definitions of the datasets related to climate change and sustainability topic detection.
Unsupervised Solutions
The earliest work on identifying corporate climate-related text was conducted by Doran and Quinn (2009), where the authors collected a large corpus of 10-K filings from 1995 to 2008, and filtered climate-related reports using hand-selected keywords. While Doran and Quinn (2009) focused on identifying climate-related text within company reports to assess how businesses addressed climate change, other early efforts took a different approach by analyzing the external impacts of media attention toward climate change on companies. For instance, Engle et al. (2020) developed the WSJ Climate Change News Index based on identifying climate-related news articles in media sources using climate-change vocabulary frequency. This index served as a tool to measure the influence of climate-change attention in the media on companies. Gorovaia and Makrominas (2025) proposed measuring the environmental content of CSR (corporate social responsibility) reports using a lexicon-based approach, with the goal of comparing CSR reports of environmental violators to companies with clean records.
Annotated Datasets
More recent studies have introduced several annotated datasets aimed at climate-related corporate text classification. One of the most prominent is ClimaText (Leippold and Varini, 2020), a large dataset from diverse sources such as Wikipedia, U.S. Security and Exchange Commission 10-K filings, and web-scraped content. The dataset is designed to cover a broad range of topics and document types to help assess climate-related discussions across different domains. It is divided into multiple subsets: wiki-doc, Wikipedia, 10-k, claims. Climatext (Wiki-doc) is annotated automatically, while Climatext (Wikipedia, 10-k and claims) is annotated by humans. In contrast, ClimateBERT’s authors (Webersinke et al., 2022) provide a smaller, more focused dataset consisting of paragraphs from companies’ annual and sustainability reports. This dataset was created as a downstream task for evaluating the ClimateBERT model, focusing on corporate disclosures. The dataset was later extended by Bingler et al. (2023), increasing its size and refining the annotation to capture both specificity and sentiment related to corporate climate discourse (see Sections 4.44.6). Another large dataset, ClimateBUG-data (Yu et al., 2024), also focuses on corporate communications but in EU banks. Beyond corporate disclosures, Sustainable Signals (Lin et al., 2023b) introduces a dataset of product reviews, each classified based on its relevance to sustainability. In addition to focusing on consumer perspectives through reviews, this dataset also examines sustainability aspects in product descriptions, providing a broader understanding of how sustainability is communicated in the context of consumer products.
Labels and Guidelines
As shown in Table 2, although these datasets address similar tasks, they differ significantly in their label definitions. The scope of the labels varies, ranging from a narrow focus on climate change to broader topics such as sustainability and environmental impact. While the task of the Climatext (wiki-doc) dataset is to detect whether a sentence is related to climate change or not, the weak label actually means that the sentence originates from a Wikipedia page related to climate change. Which is not the same task, as it focuses on the source of the statement rather than the actual content or its relevance to climate-related topics. However, the weakly labeled dataset serves as a useful filtering mechanism, helping to identify sentences that are potentially interesting. It provides an initial pool of data that can be refined with human annotations. Human annotations ensure that the labels align with the task’s goals, something weak labels alone cannot achieve.
Supervised Solutions
The solutions proposed to tackle that task range from keyword-based models (Leippold and Varini, 2020; Bingler et al., 2023) to fine-tuning BERT-like models (Leippold and Varini, 2020; Webersinke et al., 2022; Garrido-Merchán et al., 2023; Bingler et al., 2023; Yu et al., 2024) (see Table LABEL:tab:models in the appendix for details). The performance of the best models is above 90% (Garrido-Merchán et al., 2023; Bingler et al., 2023; Yu et al., 2024) (see Table 12 in the appendix for details).
Insights
The task of determining whether a statement is climate-related has been extensively studied, with numerous datasets available to support research in this area. State-of-the-art fine-tuned models now achieve near-perfect performance on these datasets, indicating that, under current conditions, the task is effectively solved. However, the labels used across the studies vary in scope. This highlights the inherent challenge of isolating climate-related discourse, given its close association with broader environmental and sustainability topics. For example, whether to include only explicit references to climate change or incorporate broader environmental and sustainability discussions, can profoundly influence the results and interpretations of such analyses. Moreover, relying only on weak labels might result in tasks that fail to capture the complexities of the subject.
4.3. Thematic Analysis
Once a text is known to be related to climate, one aims to know the exact topic of the text. Companies may prioritize specific categories while underreporting others, potentially signaling selective transparency practices such as greenwashing or greenhushing.
Task Description
Given an input sentence or a paragraph, output a subtopic related to climate change. This is a multiclass classification task that can be either supervised or unsupervised, depending on the availability of labeled data. Alternatively, it can be framed as a clustering task with the goal of discovering latent subtopic structures.
Dataset Input Labels Label details
TCFD rec. (Bingler et al., 2022) paragraphs from corporate annual reports TCFD 4 main categories: Metrics and Targets, Risk Management, Strategy, Governance and General (see appendix A for details)
TCFD (Sampson et al., 2022) paragraphs from regulatory and discretionary reports TCFD 11 recommendations: Metrics and Targets (board’s oversight, management’s role), Risk Management (identified risk, impact, resilience), … (see Appendix A for details)
FineBERT’s ESG (Huang et al., 2020) Sentences from 10-K Environmental, Social, Governance, General Environmental - e.g., climate change, natural capital, pollution and waste, and environmental opportunities
ESGBERT’s ESG (Schimanski et al., 2023b) Sentences from reports and corporate news Environment, Social, Governance and None Environmental criteria comprise a company’s energy use, waste management, pollution, […] as well as compliance with governmental regulations. Special areas of interest are climate change and environmental sustainability.
Multilingual ESG classification (Lee and Kim, 2023) Sentences from Korean reports (English and Korean) Environment, Social, Governance, Neutral, and Irrelevant Environmental factors include the reduction of hazardous substances, eco-friendly management, climate change, carbon emission, natural resources, […]
SciDCC (Mishra and Mittal, 2021) News articles (Title, Summary, Body) Environment, Geology, Animals, Ozone Layer, Climate, etc. Category in which the article was published (Automatic Label)
Transaction Ledger (Jain et al., 2023) Transaction ledger entry (description of transaction) Accounting Adjustments, Administration, Advertising, Benefits & Insurance, etc. Standardized commodity classes
ClimateEng (Vaid et al., 2022) Tweets posted during COP25 filtered by keywords (relevant to climate-change) Ocean/Water, Politics, Disaster, Agriculture/Forestry, General Sub-categories of climate-change
ESGBERT’s Nature (Schimanski et al., 2024a) Paragraphs from reports General, Nature, Biodiversity, Forest, Water Multi-label Nature-related topics
ClimaTOPIC (Spokoyny et al., 2023) CDP responses (short texts) Adaptation, Buildings, Climate Hazards, Emissions, Water, etc. Category of the question (Automatic Label)
Table 3. Summary of datasets, labels, scopes, and sources for climate-related text classification tasks.
4.3.1. Classifying Climate-Related Financial Disclosure (TCFD)
The Task Force on Climate-Related Financial Disclosures (TCFD) proposes four main categories of climate-related disclosure (Governance, Strategy, Risk Management, and Metrics and Targets) and 11 recommendations for climate-related disclosures (see Appendix A for details). Finding texts associated to these categories in company communication can help identify gaps and inconsistencies in reporting.
Unsupervised Solutions
Ding et al. (2023) introduced a score to quantify the extent of climate-related content in corporate disclosures, as well as 4 TCFD-category-based similarity scores. The score is based on sentence similarity with sentences from TCFD referential documents. They identified discrepancies that could signal greenwashing by comparing these scores with actual carbon emission data. Auzepy et al. (2023) examined reports from banks endorsing the TCFD recommendations. They proposed a fine-grained analysis of TCFD categories using a zero-shot entailment classification method on sentences.
Annotated Datasets
Although unsupervised methods can be employed for thematic analysis, recent studies have demonstrated the value of curated datasets for categorizing corporate disclosures. For instance, Bingler et al. (2022) leveraged report headings to annotate paragraphs into the four main TCFD categories. Similarly, Sampson et al. (2022) assessed the quality of climate-related disclosures using a dataset of 162k sentences, where each sentence was labeled according to the TCFD recommendations by experts through a question answering (QA) process.
4.3.2. Classifying Environmental, Social, and Governance Disclosure (ESG)
Beyond climate-specific reporting, Environmental, Social, and Governance (ESG) disclosures represent a broader set of non-financial metrics that evaluate a company’s sustainability performance. ESG reporting has become a global standard, encompassing climate-related initiatives and other dimensions of corporate responsibility. The standardized nature of ESG metrics makes them valuable for computational models that aim to detect inconsistencies in corporate claims, such as overstating environmental impacts through greenwashing. However, as noted in Berg et al. (2022)’s aggregated analysis of ESG ratings, discrepancies persist across different evaluation frameworks. These inconsistencies raise the question of whether artificial intelligence can be harnessed to improve the consistency and reliability of ESG assessments.
Unsupervised Solutions
Rouen et al. (2023) proposed an ML algorithm to describe the content of ESG reports. They used industry-topic dictionaries to compute the topic frequency using a TF-IDF-based algorithm. They used that classification to identify selective disclosure. Mehra et al. (2022) experimented with sentence similarity to extract sentences that were the most relevant to environmental factors, and (Bronzini et al., 2023) selected sentences related to ESG using INSTRUCTOR-xl (Su et al., 2023). They both subsequently used the extracted sentences to predict ESG scores.
Annotated Datasets
Multiple works (Huang et al., 2020; Schimanski et al., 2023b; Lee and Kim, 2023) proposed datasets for topic classification with the labels: Environment, Social, Governance and General. Huang et al. (2020) focused on reports while Schimanski et al. (2023b) added corporate news. Schimanski et al. (2023b) introduced a larger dataset for pre-training domain-specific models and an evaluation dataset of 2,000 sentences. While most ESG-related reports are typically available in English, some local companies provide them only in their native language. To partially address this, Lee and Kim (2023) proposed a multilingual dataset constructed from Korean corporate reports.
4.3.3. Other Topics
Although TCFD and ESG are established frameworks for analyzing corporate communications, many other topics may be of interest. These topics can follow other frameworks such as CDP, but the thematic analysis can also be unsupervised, effectively discovering the topics mentioned.
Unsupervised Solutions
Kang and Kim (2022) analyzed the content of sustainable reports with thematic analysis. It was conducted using sentence similarity between the reports and content from the SDGs website. Yu et al. (2024) examined the themes covered in the climate-related sections of bank reports by employing a dual approach: using hand-selected keywords to target anticipated topics and applying clustering techniques to group climate-related sentences and uncover underlying themes. Brié et al. (2022) leveraged ClimateBERT to extract climate-related paragraphs from corporate documents. They then applied a Structural Topic Model (STM) (Margaret E. Roberts and Airoldi, 2016) to identify thematic clusters within the data. This STM-based approach was also used by Fortes and Le Guenedal (2020) to identify influential topics for EUR/USD exchange rate. Their findings highlighted a significant discrepancy between the European Central Bank (ECB) and the Federal Reserve (FED), with the ECB attributing greater importance to climate change in the context of financial stability than the FED. Such thematic analysis can be used to identify selective disclosure.
Annotated Datasets
Spokoyny et al. (2023) introduced ClimaTOPIC, a topic classification dataset derived from responses to the CDP questionnaires, organized into 12 categories that align with the thematic structure of the CDP questions. This can be used to identify sections relevant to a particular CDP topic (e.g., Emissions, Adaptation, or Energy), enabling the extraction of structured information from unstructured documents. Schimanski et al. (2024a) presents a dataset focused on nature-related topics (such as Water, Forest, and Biodiversity) extracted from annual reports, sustainability reports, and earnings call transcripts (ESGBERT Nature). While less explicitly linked to climate change, these topics often reflect the consequences of environmental degradation associated with climate issues. While NLP is often used to analyze corporate report texts, it can also be used to analyze tables. For example, Jain et al. (2023) proposes estimating Scope 3 emissions  [^3] Scope 3 emissions refer to all indirect greenhouse gas (GHG) emissions that occur in a company’s value chain, excluding emissions from the company’s own operations (Scope 1) and its purchased electricity (Scope 2). based on corporate transactions (see Section 4.10 for details). The first step involves grouping expenses into standardized commodity classes, for which they introduced a dataset of 8K examples of corporate expenses categorized accordingly. As previously discussed, social media and news articles offer a broader perspective for cross-referencing corporate communications. For example, Vaid et al. (2022) introduced ClimateEng, a climate-related topic classification dataset composed of tweets grouped into five categories: Disaster, Ocean/Water, Agriculture/Forestry, Politics, and General. This dataset can be used to detect greenwashing by comparing the public’s concerns with the narrative presented in corporate sustainability reports. Similarly, Mishra and Mittal (2021) published the Science Daily Climate Change dataset (SciDCC), consisting of 11,000 news articles grouped into 20 climate-related categories. This dataset offers another avenue for identifying greenwashing, as companies’ communications can be cross-referenced with external news coverage, potentially highlighting contradictions or omissions in corporate reports.
4.3.4. Models and Conclusion
Supervised Solutions
The most common solution proposed for these tasks is fine-tuning a Transformer model (Huang et al., 2020; Schimanski et al., 2023b; Lee and Kim, 2023; Bingler et al., 2022; Sampson et al., 2022; Vaid et al., 2022; Spokoyny et al., 2023; Schimanski et al., 2024a; Jain et al., 2023). To contextualize the performance of Transformer models, some studies also report performances of classical baselines (Huang et al., 2020; Spokoyny et al., 2023; Schimanski et al., 2024a; Jain et al., 2023; Bingler et al., 2022; Sampson et al., 2022). Bingler et al. (2022) proposed a custom approach combining logistic regression with features from a fine-tuned language model, while Sampson et al. (2022) also evaluated clustering techniques and stacked models. The performance of the best-performing model is systematically high (above 80%), except on SciDCC and ClimaTOPIC. Classical approaches (SVM, NB) also reach good performances, under-performing by less than 20% on all datasets (except on Bingler et al. (2022)’s dataset). The TF-IDF baseline even outperformed fine-tuned Transformers on Sampson et al. (2022)’s dataset. This shows that those topics have distinguishable vocabularies.
Automatic Labels
The performance gap observed in Bingler et al. (2022) may be attributed to using paragraph-level inputs rather than sentence-level inputs aggregated for classification, as fine-tuned models also achieve around 20% precision at the paragraph level. The lower performances on SciDCC(Mishra and Mittal, 2021) and ClimaTOPIC(Spokoyny et al., 2023) are likely caused by the automation of the labeling process. The labels are therefore not designed to be labels. In SciDCC there are labels that are highly similar (e.g. Endangered Animals and Extinctions), or labels that include other labels (e.g. Environment, Climate, Pollution). As they are categories from a journal, the categories changed in time. For ClimaTOPIC, the labels are question categories, which are designed to group questions not to precisely identify them. Therefore a question about ”the emission of a building” might fit in either Emissions or Building, yet it is assigned only one label.
Insights
Topic classification is a well-established area of research in natural language processing (NLP) and has been extensively applied to climate-related topics. Fine-tuned Transformer models proposed in the literature consistently demonstrate near-perfect performance on human-annotated datasets, indicating that the task is largely solved. When reported, classical keyword baselines also perform relatively well, achieving significantly better than random performance and approximately 80% of the performance of fine-tuned models, suggesting the topics have distinct vocabularies. Future research could investigate topics that are lexically similar but distinct in their subject or focus. Topic detection serves as a valuable tool for structuring documents and has been utilized to analyze the extent to which companies address specific topics, enabling the detection of selective disclosure (Bingler et al., 2023, 2022). However, it is critical to ensure that labels are well-defined and data is accurately annotated. Lower performance observed on automatically generated labels may reflect challenges in predictability, potentially due to difficulties caused by the automatic annotation. These works might be revisited to conduct human annotations.
4.4. In-depth Disclosure: Climate Risk Classification
Climate change can bring both risks and opportunities for companies. Potential risks are, e.g., reputational risks (e.g. environmental controversies), regulatory risks (e.g. new regulations on emissions), and physical risks (e.g. droughts impacting production). Opportunities are, e.g., financial opportunities (e.g. benefiting from grants that aim to support less polluting industries), market opportunities (e.g. electric cars becoming more popular with environmentally conscious clients), etc. If a company systematically avoids discussing climate-related risks or disproportionately emphasizes opportunities, it creates a biased narrative that can serve as an indicator of greenwashing.
Task Description
Given an input sentence or a paragraph, output “opportunity” or “risk” label. Some works focus only on risks, classifying them into types of risks (e.g., physical risk, reputational risk, regulatory risk, or transition risk)
Dataset Input Labels
Ask BERT’s Climate Risk (Kölbel et al., 2021) Sentences from TCFD’s example reports and non-climate-related sentences Risk type: Transition risk, physical risk, and general risk (no guidelines)
Climate Risk (Friederich et al., 2021) Paragraphs from European companies annual reports Risk type:Acute, Chronic, Policy & legal, Tech & Market, Reputational, and Negative (no guidelines)
ClimateBERT’s Sentiment (Bingler et al., 2023) Paragraphs from companies’ annual reports Risk or threat that negatively impacts an entity of interest (negative sentiment); or Opportunity arising due to climate change (positive sentiment); Neutral otherwise.
Sentiment Analysis (Xiang and Fujii, 2023) paragraphs from academic texts on climate change and health published between 2013 and 2020 Risk (negative) if it discusses climate change causing public health issues, serious consequences, or worsening trends, greenwashing. Opportunity (positive): highlights potential benefits, positive actions, or research addressing gaps. Neutral otherwise.
Table 4. Summary of datasets, labels, scopes, and sources for climate-related Risk classification.
Unsupervised Solutions
Similarly to climate-related detection, climate risk classification has been tackled with keyword-based approaches. Li et al. (2020), Kheradmand et al. (2021) and Chou et al. (2023) proposed using dictionaries of words related to climate risk to identify paragraphs dealing with climate risk. Li et al. (2020) constructed risk measures based on the frequency of the terms in the risk dictionaries. Chou et al. (2023) analyzed the topics mentioned in conjunction with physical and transition risks. Sautner et al. (2023) proposed using one dictionary for risk and one for opportunity classification.
Annotated Datasets
Several studies have introduced specialized annotated datasets for risk/opportunity classification. For instance, Kölbel et al. (2021) created a dataset for climate-related risk classification with three categories: physical risk, transition risk, and general risk, using active learning for annotation. Similarly, Friederich et al. (2021) developed an annotated dataset for risk classification with five labels, covering acute physical risk, chronic physical risk, policy and legal risks, technology and market risks, and reputational transition risks. Bingler et al. (2023) released a dataset as part of the ClimateBERT downstream tasks, focusing on classifying paragraphs from corporate reports into three categories: opportunity, neutral, or risk. Extending beyond corporate disclosures, Xiang and Fujii (2023) compiled a climate-related risk/opportunity classification dataset of academic texts from the Web of Science and Scopus.
Supervised Solutions
The solutions proposed are similar to the other tasks: fine-tuned Transformer models (Kang and Kim, 2022; Kölbel et al., 2021; Friederich et al., 2021; Bingler et al., 2023; Webersinke et al., 2022; Xiang and Fujii, 2023) and keyword-based-features with simple models (Kölbel et al., 2021; Friederich et al., 2021; Bingler et al., 2023) (see Table LABEL:tab:models in the appendix for details). The notable exception is Xiang and Fujii (2023), who evaluated LSTM-based solutions alongside the Transformers models. Most studies (Kang and Kim, 2022; Kölbel et al., 2021; Bingler et al., 2023; Webersinke et al., 2022; Xiang and Fujii, 2023) reported high performances, above 80%. Kölbel et al. (2021) and Bingler et al. (2023) also reported good performances for the keyword-based baselines (72% and 84%). The performance reported by Friederich et al. (2021) on climate risk tasks highlights distinct challenges across subtasks. When tasked with identifying the specific type of risk in sentences already known to be about climate risks, word-based models outperformed fine-tuned Transformers. This suggests that each risk type had a distinct vocabulary. In contrast, when working with a dataset that reflects real-world proportions of risk and non-risk examples—where the data is heavily imbalanced—classical models struggled due to the overlap in vocabulary between general and climate risk-related text.
Difference between sentiment and risk
It is important to distinguish risk/opportunity classification from sentiment analysis. While risk/opportunity classification is inspired by sentiment analysis (Bingler et al., 2023), they play a complementary role. A statement such as “With proactive climate risk management, we are ready to tackle extreme weather disruptions, ensuring resilience” is classified as positive using a sentiment analysis model (Pérez et al., 2021), and as mentioning a risk using a risk/opportunity classification model (Bingler et al., 2023). Sentiment analysis focuses on the form of the statement, while risk is about the content. As the literature on sentiment analysis (Wankhade et al., 2022) is broad, the analysis of the tone of the statement can be done using existing models (Polignano et al., 2022; Kang and Kim, 2022) complementing approaches on risk classification.
Insights
The performance of fine-tuned models might indicate that the task of identifying texts talking about risks is solved. However, the work by Friederich et al. (2021) shows limitations when experimenting with heavily imbalanced datasets, which actually correspond to real-world settings. Future research could focus on such skewed settings.
4.5. Green Claim Detection
Identifying greenwashing involves more than just finding text related to climate change, as a company might simply state factual information about the climate without making any commitments on it. Hence, we now focus on detecting green claims, i.e., claims that a product, service, or corporate practice either contributes positively to environmental sustainability or is less harmful to the environment than alternatives (Stammbach et al., 2023; Woloszyn et al., 2021). The European Commission calls these claims “environmental claims”:
Dataset Input Positive Label description
Green Claims (Woloszyn et al., 2021) Marketing Tweets Environmental (or green) advertisements refer to all appeals that include ecological, environmental sustainability, or nature-friendly messages that target the needs and desires of environmentally concerned stakeholders.
Environmental Claims (Stammbach et al., 2023) Paragraph from reports Environmental claims refer to the practice of suggesting or otherwise creating the impression […] that a product or a service is environmentally friendly (i.e., it has a positive impact on the environment) or is less damaging to the environment than competing goods or services […] In our case, claims relate to products, services, or specific corporate environmental performance.
Table 5. Label definitions of the datasets on green claims detection.
Task Description
Given an input sentence or a paragraph, output a binary label, “green claim” or “not green claim”.
Annotated Datasets
The most notable approaches aiming at identifying climate-related claims introduced annotated datasets (Woloszyn et al., 2021; Stammbach et al., 2023). Woloszyn et al. (2021) proposed to focus on social media marketing through the detection of green claims in tweets. Stammbach et al. (2023) proposed a dataset for environmental claim detection in paragraphs from corporate reports.
Supervised Solutions
Both Stammbach et al. (2023) and Woloszyn et al. (2021) evaluated fine-tuned Transformer models on the datasets. Stammbach et al. (2023) also evaluated classical approaches and experimented with a model fine-tuned on general claim detection (Arslan et al., 2020) applied to their domain-specific task (see Table LABEL:tab:models in the Appendix for details). Woloszyn et al. (2021) and Stammbach et al. (2023) demonstrated good performances with fine-tuned Transformer (above 84%), yet they have very different Inter-Annotator Agreement (IAA), with Krippendorff’s $\alpha=0.8223$ for Woloszyn et al. (2021) and $\alpha=0.47$ for Stammbach et al. (2023). This might be explained by the difference in context: marketing tweets might be easier to understand for humans and more self-sufficient compared to paragraphs extracted from reports. Additionally, Woloszyn et al. (2021) experimented with adversarial attack, showing the sensibility to character-swap and word-swap. This shows that the models tend to rely heavily on particular words, showing only a superficial understanding. While it is difficult to generalize that conclusion given the small size of the dataset, it still indicates that BERT-like architecture can over-fit or rely on superficial cues instead of building an accurate representation of the paragraph. It is essential to assess model performance in challenging scenarios, where the presence of nonsensical or noisy inputs can reveal the fragility of model comprehension.
Insights
Green claims have been analyzed in company reporting (Stammbach et al., 2023) and social media communication (Woloszyn et al., 2021). Fine-tuned models can solve the task rather well, even if challenges remain:
  • Inter annotator agreement: The annotator agreement remains low for green claim detection in reports (Stammbach et al., 2023).
  • Integration with existing claim detection literature: While claim detection is a well-established field with extensive literature, studies on environmental claims do not fully connect with existing research on claim detection. In particular, they do not distinguish between Claims, Verifiable Claims (claims that can be checked), and Check-worthy Claims (claims that are interesting to verify) (Panchendrarajan and Zubiaga, 2024).
  • Evaluation Sensitivity and Real-World Robustness: The literature shows that fine-tuned models are sensitive to adversarial attacks, meaning small perturbations of the text influence greatly the performance of the classifier. As for all tasks, it is also important evaluate models robustness in real-world settings. Poor data quality might induce perturbations in the texts, reducing the performance of models.
4.6. Green Claim Characteristics
Once we have established that a sentence is climate-related (Section 4.2) and that it is a claim about the company (Section 4.5), we can endeavor to further classify the claim into fine-grained categories. Table 6 shows various characteristics of claims that have been studied.
Dataset Input Labels
Implicit/Explicit Green Claims (Woloszyn et al., 2021) Marketing Tweets Implicit green claims raise the same ecological and environmental concerns as explicit green claims (see definition in Section 4.5), but without showing any commitment from the company. If the tweet does not contain a green claim then No Claim.
Specificity (Bingler et al., 2023) Paragraph from reports A paragraph is Specific if it includes clear, tangible, and firm-specific details about events, goals, actions, or explanations that directly impact or clarify the firm’s operations, strategy, or objectives. Non-specific otherwise.
Commitments & Actions(Bingler et al., 2023) Paragraph from reports A paragraph is a commitment or an action if it contains targets for the future or actions already taken in the past.
Net Zero/Reduction (Schimanski et al., 2023a) Paragraph from Net Zero Tracker (Lang et al., 2023) The paragraph contains either a Net-Zero target, a Reduction target or no target (None)
Table 6. Label definitions the datasets related to characterization of green claims.
Task Description
Given an input sentence or a paragraph labeled as a green claim, output a more fine-grained characterization of the claim. This is a multi-label classification task; the labels can be about the form (e.g. specificity) or the substance (e.g. action, targets, facts).
Annotated Datasets
Existing works characterized both the content and the form of claims. On the content dimension, there are datasets for identifying if the claim is about an action or a commitment (Bingler et al., 2023; Schimanski et al., 2023a). Bingler et al. (2023) introduce a dataset for the identification of commitments and actions (Yes/No), and Schimanski et al. (2023a) released one for the identification of reduction targets (net zero, reduction, general). On the form dimension, Woloszyn et al. (2021) characterized claims as implicit or explicit, and Bingler et al. (2023) annotated their dataset on the specificity of claims (Specific/Not specific). Bingler et al. (2023) ultimately used the specificity and commitment/action characteristics to identify cheap talks related to climate disclosure.
Supervised Solutions
Woloszyn et al. (2021); Schimanski et al. (2023a) and Bingler et al. (2023) experimented with fine-tuned Transformers. Bingler et al. (2023) also evaluated classical baselines (SVM, NB), and Schimanski et al. (2023a) experimented with GPT-3.5-turbo (see Table LABEL:tab:models in the Appendix for details). All approaches reached good performances, in particular fine-tuned Transformers reaching performances above 80%. The exception remains the Specificity classification with a F1-score of 77% (close to the baselines). This low performance might be intrinsic to the task. Indeed, humans have disagreement when distinguishing specific and unspecific claims: Bingler et al. (2023) measured a low IAA on Specificity (Krippendorff’s $\alpha=0.1703$). This might indicate that the task is not well defined in the first place.
Insights
The high performances of fine-tuned models indicate that this type of task is solved. Furthermore, the demonstrated ability of GPT-3.5 as a zero-shot classifier (Schimanski et al., 2023a) highlights the potential of large language models to classify these characteristics without requiring extensive annotated datasets. These characteristics are particularly useful for identifying statements that may indicate greenwashing. For example, Bingler et al. (2023) utilized the proportion of non-specific commitments as a Cheap Talk Index demonstrating their practical applications. However, characteristics such as specificity are inherently ambiguous and subjective, as evidenced by the low inter-annotator agreement (IAA). Therefore, future research could focus on disambiguating what constitutes a specific statement from a non-specific one more objectively.
4.7. Green Stance Detection
Beyond making claims about the environmental impact of their products and processes, companies contribute to environmental discussions and can have an impact on regulatory frameworks through their communications and industry presence. It is thus helpful to understand the stance of an organization on the existence and gravity of climate change, on climate mitigation and adaptation efforts, and on climate-related regulations.
Dataset Input Labels Label details
ClimateFEVER (evidence)  (Leippold and Diggelmann, 2020) A claim and an evidence sentence from Wikipedia Support, Refutes, Not Enough Information Determines the relation between a claim and a single evidence sentence
LobbyMap (Stance) (Morio and Manning, 2023) Page from a company communications (report, press release, …) Strongly supporting, Supporting, No or mixed position, Not supporting, Opposing Given the policy and the page, classifies the stance
Global Warming Stance Detection (GWSD) (Luo et al., 2020) Sentences from news about global warming Stance of the evidence (Agree, Disagree, Neutral) toward the claim: Climate-Change is a serious concern.
ClimateStance (Vaid et al., 2022) Tweets posted during COP25 filtered by keywords (relevant to climate-change) Stance towards climate change prevention: Favor, Against, Ambiguous. (Stance used as a broad notion including sentiment, evaluation, appraisal, …)
Stance on Remediation Effort (Lai and Chen, 2023) Texts extracted from the TCFD sections of the financial reports The text indicates support for climate change remediation efforts or refutation of such efforts. (No guidelines)
Related subtask
ClimateFEVER (claim)  (Leippold and Diggelmann, 2020) A claim and multiple evidence sentences from Wikipedia Support, Refutes, Debated, Not Enough Information Determines if a claim is supported by a set of retrieved evidence sentences
LobbyMap (Page) (Morio and Manning, 2023) Page from a company communications (report, press release, …) 1/0 Contains a stance on a remediation policy
LobbyMap (Query) (Morio and Manning, 2023) GHG emission regulation, Renewable energy, Carbon tax, … Classifies the remediation policy
Table 7. Summary of datasets, labels, scopes, and sources for climate-related stance detection.
Task Description
Given two input sentences or paragraphs, one labeled as the claim and one as the evidence, predict the stance between the two: supports, refutes or neutral. Some studies fix the claim and only vary the evidence (e.g. the claim is always Climate change poses a severe threat), training a model to predict the stance of the evidence in respect to the fixed claim. Other studies train a model to predict the relation between varying claims/evidences.
Related subtask
The first subtask is collecting the evidences when the claims are already available (e.g. if they were collected manually). In order to build these datasets, research can rely on simple heuristics such as downloading all tweets published during the COP25 filtered with keywords (Vaid et al., 2022). However, other researchers performed a more elaborate procedure. Leippold and Diggelmann (2020) proposed a pipeline for collecting evidence. To retrieve relevant evidence from Wikipedia for a given claim, the pipeline involves three steps: document-level retrieval using entity-linking and BM25 to identify top articles, sentence-level retrieval using sentence embeddings trained on the FEVER dataset to extract relevant sentences, and sentence re-ranking using a pretrained ALBERT model to classify and rank evidence based on relevance. The second one encompasses the broader process: identifying claims and finding evidence, before predicting the relation. Wang et al. (2021) used a generic claim detection model trained on ClaimBuster (Arslan et al., 2020) for the claims and Google Search API to collect evidence. Finally, the last subtask is to train multiple models on each step: identifying evidences, identifying the claim and classifying the stance. This is the approach used by (Morio and Manning, 2023).
Annotated Datasets
As described previously, researchers built datasets of evidence related to one claim while focusing stance in tweets on seriousness of climate change such as Global Warming Stance Detection (GWSD) (Luo et al., 2020), ClimateStance (Vaid et al., 2022) stance on climate change of tweets posted during the COP25, or Lai and Chen (2023)’s dataset which focuses on the stance toward climate change remediation efforts. Another research initiative built a dataset of claim-evidence pairs, ClimateFEVER, and then trained a model on the stance classification (Leippold and Diggelmann, 2020; Wang et al., 2021). Finally, Morio and Manning (2023) proposed a dataset to assess corporate policy engagement built upon LobbyMap, which tackles the 3 steps: finding pages with evidence, identifying the claims targeted by that evidence, and classifying the stance.
Supervised Solutions
The solution proposed for to tackle stance classification are fine-tuned Transformer models (Vaid et al., 2022; Vaghefi et al., 2022; Xiang and Fujii, 2023; Morio and Manning, 2023; Webersinke et al., 2022; Wang et al., 2021; Spokoyny et al., 2023; Lai and Chen, 2023; Luo et al., 2020) and classical approaches (Spokoyny et al., 2023; Morio and Manning, 2023; Luo et al., 2020). Vaid et al. (2022) also experimented with FastText and Xiang and Fujii (2023) with LSTM-based models (See Table LABEL:tab:models in the Appendix for details). The performances are quite heterogeneous, Lai and Chen (2023) reaching F1-scores around 90% on classification of stance on remediation efforts, Luo et al. (2020) around 72% on GWSD, while performance on ClimateStance could not exceed 60%. Performances on ClimateFEVER are also quite low, however, when selecting only non-ambiguous examples, Xiang and Fujii (2023) could reach performances around 80%. Finally, performances on LobbyMap (Morio and Manning, 2023) are quite low (between 31% and 57.3% depending on the strictness of the metric). Overall, the performances show that the datasets are challenging (see Table 19 in the Appendix for details).
Exhaustivity
Datasets such as ClimateFEVER (Leippold and Diggelmann, 2020) and LobbyMap (Morio and Manning, 2023) rely on automated construction methods that may not ensure exhaustivity of evidence coverage. ClimateFEVER utilizes BM25 and Wikipedia for evidence retrieval, which could result in missing relevant information, particularly when compared to more advanced retrieval methods. LobbyMap relies on the LobbyMap.org website, which was not designed to be exhaustive. These limitations should be further investigated.
Insights
If an organization presents itself as environmentally friendly while simultaneously promoting climate-skeptic narratives or opposing climate remediation efforts, it probably is creating a misleading portrayal of its environmental stance. On the contrary, if they are aligned, it supports the authenticity of the organization’s efforts. Fortunately, stance detection has been extensively studied, yielding promising results; however, current performance levels leave room for improvement. Future research should focus on enhancing both methods and datasets to address these gaps. While existing datasets provide strong foundations, they may suffer from a lack of exhaustivity, particularly in evidence retrieval and coverage, which requires further investigation.
4.8. Question Answering
Question answering (QA) is a known task:
Task Description
Given an input question and a set of resources (paragraphs or documents), produce an answer to the question.
Dataset Input Labels
ClimaQA  (Spokoyny et al., 2023) The text from a response to one of the CDP questions; and one of the questions from the CDP questionnaire 1: the response answers this question 0: The response does not answer this question, but another one
ClimateQA (Luccioni et al., 2020) Sentence from reports and a question based on the TCFD recommendations 1: the sentence answers the question 0: The sentences does not answer the question
ClimaINS  (Spokoyny et al., 2023) The text from a response to one of the questions from the NAIC questionnaire MANAGE, RISK PLAN, MITIGATE, ENGAGE, ASSESS, RISKS The labels correspond the 8 questions asked in the NAIC questionnaires
Table 8. Summary of datasets, labels, scopes, and sources for climate-related QA datasets.
QA can be used for climate-specific applications such as structuring the information related to climate change from documents (Luccioni et al., 2020; Schimanski et al., 2023a), building chatbots with climate-related knowledge to make it more accessible to non-experts (Vaghefi et al., 2023; Rony et al., 2022), or helping identify potentially misleading information (Ni et al., 2023).
Solution
The QA process can be divided into two main steps (Kraus et al., 2023): the retrieval step and the answer generation step. The retrieval step involves locating the relevant information or answer to the user’s query in external documents. The answer generation step involves formulating a response to the user’s query from (1) the information retrieved from documents in the first step or (2) the pre-trained internal knowledge of a model.
Retriever
Retriever systems can be generic, using techniques such as sentence similarity, BM25, or document tags to filter documents and retrieve specific passages. Schimanski et al. (2024b) published ClimRetrieve, a benchmark for climate-related information retrieval in corporate reports. They found that simple embedding approaches are limited. Therefore, several works have specifically focused on improving the identification of answers within climate-related documents (Luccioni et al., 2020; Spokoyny et al., 2023). Luccioni et al. (2020) reformulated TCFD recommendations as questions and annotated reports to identify answers to the questions. The dataset includes questions paired with potential answers, each labeled to indicate whether the answer addresses the question or not. Based on the QA model trained on this dataset, they also proposed a tool called ClimateQA. Spokoyny et al. (2023) focused on questions from the CDP questionnaire and the NAIC Climate Risk Disclosure survey. As the responses to those questionnaires are publicly available, they constructed two datasets, ClimaINS and ClimaQA. ClimaQA is similar to ClimateQA (question/potential answer pairs). On the contrary, ClimaINS is a QA dataset framed as a classification task: it contains the responses from the NAIC survey, and each response is labeled as answering one of the 8 questions of the survey. Once a model has been trained on these datasets, it can be used to search for answers to the questions in other documents (effectively retrieving the information from an unstructured document). For ClimaINS, the authors experimented with multiple fine-tuned Transformers. As ClimaQA is framed as a retrieval task, the authors used BM25, sentence-BERT, and ClimateBERT to rank possible answers. Using the methodology evaluated on ClimaQA, they proposed a system to extract responses to CDP questions directly from unstructured reports and automatically fill the questionnaire (Spokoyny et al., 2023). Luccioni et al. (2020) trained and evaluated a RoBERTa-based model called ClimateQA on finding answers to TCFD-based questions. They all reached good performance but with room for improvement (see Table 21 in the Appendix for details), demonstrating the feasibility of the task but also the need for further research.
Answer generation
After the retrieval, the second step of question answering is to generate an answer for the question (given input resources or not). The generation can be simple such as finding a particular information in a paragraph. For example, Schimanski et al. (2023a) used a QA model (RoBERTa SQuaD v2 (Rajpurkar et al., 2016)) to extract the target year, the percentage of reduction, and the baseline year from the reduction target of a company from an input text that contains an emission reduction target. While it is possible to rely on generalist models, as they are proficient few-shot learners (Brown et al., 2020; Thulke et al., 2024), multiple studies proposed using domain-specific models. Climate Bot (Rony et al., 2022) built a dataset of climate-specific QA. Given a question and a document (scientific/news), the model should find the span answering the question. Mullappilly et al. (2023) proposed a dataset of question-answer pairs based on ClimaBench (Spokoyny et al., 2023) and CCMRC (Rony et al., 2022) to train a model to generate answers. Thulke et al. (2024) introduced a climate-specific corpus of prompt/completion pairs for Instruction Fine-Tuning created by experts and non-experts. (Rony et al., 2022) finetuned an ALBERT model on their climate-specific QA dataset. Mullappilly et al. (2023) trained a Vicuna-7B on their climate-specific answer generation dataset. And Thulke et al. (2024) trained LLama-2-based models (ClimateGPT) on their domain-specific prompt/completion dataset. They also evaluated their ClimateGPT model alongside multiple LLMs on climate-related benchmarks. They found that the domain-specific models outperformed generalist ones (see Table 21 in the Appendix for details).
Retrieval Augmented Generation
By combining the last two components (a retriever and a QA system), one can develop Retrieval-Augmented Generation (RAG) systems. Rony et al. (2022) provide the first example of a RAG system in climate-related tasks based on sentence-BERT to retrieve documents and AlBERT to identify the answer. More recently, Ni et al. (2023) proposed ChatReport, a methodology based on ChatGPT for analyzing corporate reports through the lens of TCFD questions using RAG. The reports, the answers to the TCFD questions, and a TCFD conformity assessment are chunked and stored in a vector store for easy retrieval. Each chunk and/or answer can be retrieved and injected into the prompt. They added in the prompt the notions of greenwashing and cheap talk to invite the model to provide a critical analysis of the retrieved answers. Similarly, Vaghefi et al. (2023) introduced ChatClimate, a RAG-based pipeline to augment GPT-4 with knowledge about IPCC reports. While the previous LLM-based approaches rely on closed-source models, Mullappilly et al. (2023) and Thulke et al. (2024) proposed RAG-based pipelines relying on open-source models. Rony et al. (2022) is the only study that evaluated their RAG pipeline using classical metrics (F1-score, BLEU, METEOR). Unfortunately, those metrics are not well aligned with human judgment for text generation (Chen et al., 2019). Therefore Vaghefi et al. (2023) and Mullappilly et al. (2023) relied on human or ChatGPT evaluations. They conclude that their approaches improve on non-domain-specific RAG systems.
Insights
The main application of Question-Answering (QA) systems is to analyze complex documents, such as corporate climate reports, that can be particularly useful for greenwashing detection. Modern LLMs achieve strong results in QA due to instruction-following fine-tuning, which improves adherence to specific instructions, and Retrieval-Augmented Generation (RAG), which bases its answers on retrieved documents. However, retrieval systems remain a performance bottleneck (Maekawa et al., 2024). Furthermore, answer generation is challenging to evaluate (Sai et al., 2022), often requiring human feedback or advanced model-based assessments.
4.9. Classification of Deceptive Techniques
In analyzing climate-related discourse, it can be useful to identify rhetorical strategies that might obscure or misrepresent an entity’s stance. For example, expressing support for climate remediation policies on one side (Morio and Manning, 2023), and promoting arguments downplaying the urgency of climate change on the other (Coan et al., 2021) could signal a lack of authenticity. Another example could be promoting a product through misleading rhetoric, by, for example, claiming that a product is better because it is natural, which would be an “appeal to nature” – a fallacious argument (Vaid et al., 2022; Jain et al., 2023). Detecting these deceptive techniques could help identify misleading communications.
Dataset Input Labels (Labels details)
LogicClimate (Jin et al., 2022) texts from climatefeedback.org Faulty Generalization, Ad Hominem, Ad Populum, False Causality, … Classifies fallacies (Multi-label)
Neutralization Techniques (Bhatia et al., 2021) paragraphs from other previous works on climate-change Denial of Responsibility, Denial of Injury, Denial of Victim, Condemnation of the Condemner, … Classifies neutralization techniques
Contrarian Claims (Coan et al., 2021) paragraphs from conservative think tank No Claim, Global Warming is not happening, Climate Solutions won’t work, Climate impacts are not bad, … Classifies arguments into super/sub-categories of climate science denier’s arguments
Table 9. Summary of datasets, labels, scopes, and sources for tasks related to deceptive techniques in climate-related context.
Task Description
The goal is to classify statements into argumentative categories: fallacies, types of arguments, or rhetorical techniques.
Unsupervised Solutions
Oppong-Tawiah and Webster (2023) frame greenwashing as fake news. They propose to tackle this identification through the form perspective. They developed a profile-deviation-based method to detect greenwashing in corporate tweets by comparing linguistic cues (e.g., quantity, specificity, complexity, diversity, hedging, affect, and vividness) to theoretically ideal profiles of truthful and deceptive communication. They compute a greenwashing score as Euclidean distances.
Annotated Datasets
An organization seeking to justify limited action might adopt a rhetoric that downplays the urgency of global warming or dismisses the impact of negligent behavior. These kinds of climate contrarian arguments mostly fall into a finite set of categories (e.g. “Solutions will not work”, or “Human influence is not demonstrated”). Therefore, Coan et al. (2021) constructed a taxonomy of such contrarian claims and published a large dataset annotated with such claims. Similarly, Bhatia et al. (2021) proposed a dataset to classify neutralization techniques, i.e., rationalizations that individuals use to justify deviant or unethical behavior (e.g. Denial of the victim). Finally, Jin et al. (2022) studied fallacious arguments and how they apply to climate change more broadly. They constructed LogicClimate, a climate-specific dataset of sentences annotated with fallacies, using articles from the Climate Feedback website.
Supervised Solutions
Fine-tuned Transformer architectures and classical approaches have been evaluated on each dataset (Jin et al., 2022; Bhatia et al., 2021; Coan et al., 2021). Jin et al. (2022) experimented with an ELECTRA model, fine-tuned to use the structure of the argument. The performance of fine-tuned approaches ranges from 58.77% to 79%. The models struggle with fallacy detection in LogicClimate, but perform well with generalist fallacy detection, contrarian claims, and neutralization techniques. Thulke et al. (2024) experimented with multiple zero-shot approaches using multiple LLM on the binary classification of contrarian claims using Coan et al. (2021)’s dataset. Their climateGPT-70B and LLama-2-Chat-70B models both reached an F1-score of 72.5% (see Table 23 in the Appendix for details).
Insights
The low performance of models on tasks such as fallacy detection and neutralization classification indicates the inherent complexity of these tasks. Fallacy detection, for one, is known to be inherently subjective (Helwe et al., 2024). The detection of neutralization techniques, too, appears to be subjective, as even human annotators achieve only a moderate performance level of 70% F1-score. Furthermore, both LogicClimate (Jin et al., 2022) and the neutralization dataset (Bhatia et al., 2021) are very small in size. To address these limitations, future research could explore several directions, including increasing the size of the datasets, defining the labels more precisely, or permitting multiple correct annotations (Helwe et al., 2024) to acknowledge viable disagreement. Additionally, the analyses on rhetorical techniques could be combined with other analyses (such as stance detection) to search for communication patterns.
4.10. Environmental Performance Prediction
Greenwashing can be interpreted as a misrepresentation of the company or product’s environmental performance. The environmental performance is often summarized by a quantitative metric such as the ESG score, the Finch score [^4] The Finch Score is a sustainability rating system designed to help consumers make eco-friendly choices by evaluating products on a scale from 1 to 10. See https://www.choosefinch.com/, or CO2 emissions. These scores are usually not directly mentioned in the company reports, but have to be inferred based on company communications:
Task Description
Given an input company report, output an environment-related quantitative value (such as the ESG score or the amount of Carbon Emission), even if that value is not mentioned in the report. In variants of this task, the goal is to predict not a value, but a range for that value, out of a set of possible ranges.
Unsupervised Solutions
Jain et al. (2023) proposed to predict the Scope 3 emissions from the list of financial transactions of the company. They first performed classification of all transaction descriptions to map them to their emission factors (emission per \$ spent), and then compute the emissions of each transaction. For ESG score prediction, Bronzini et al. (2023) proposes to use LLMs as few-shot learners to extract triples of ESG category, action, and company from sustainable reports, and to construct a graph representation with few-shot learning. They demonstrated the triplet generation using Alpaca, WizardLM, Flan-T5, and ChatGPT on a few examples. They could analyze disclosure choices and company similarities using the constructed graph. More importantly, they also used the graph to interpret ESG scores through the interpretability analysis of OLS predictions, effectively predicting the ESG score.
Annotated Datasets
Mehra et al. (2022) proposed to predict the changes in the ESG score instead of the actual value. They constructed their dataset by extracting the three sentences most relevant to the environment from financial reports and associated them with the E score [^5] ESG scores are usually aggregated scores along multiple dimensions. In this study, they are focusing on the ”Environment” part of the ESG score.’s change and direction of change. Instead of relying on scores, Clarkson et al. (2020) proposed to focus on good/bad CSR (corporate social responsibility) performers. Their approach evaluates CSR performance based on linguistic style rather than content, aiming to identify whether language patterns alone influence the perception of CSR quality. Sharma et al. (2024) introduced a multi-modal dataset of the tweets of companies to predict a company’s ESG unmanaged risk. Finally, focusing on products and not companies, Lin et al. (2023b) introduced a dataset of online product descriptions and reviews used to predict the Finch Score. All these scores can be used to identify companies and products that are actually environmentally friendly, helping users distinguish actual sustainability from greenwashing. However, they can also be used to find discrepancies between the communicated green-ness of a product or company (e.g. the percentage of climate-related texts in reports) and the actual green-ness (e.g. ESG score or quantity of emissions), which might indicate cases of greenwashing.
Supervised Solutions
Mehra et al. (2022) evaluated fine-tuned Transformers on ESG change prediction. Clarkson et al. (2020) experimented with hand-chosen features plugged into random forest and SVM on CSR performers prediction. Lin et al. (2023b) also experimented with traditional approaches (e.g. gradient boosting) and custom architectures based on Transformer models on Finch score prediction. Sharma et al. (2024) introduced baseline models, which are simple image and text embedding models (e.g., CLIP or sentence-BERT) with a classification head on ESG score prediction. A detailed list of models is reported in Table LABEL:tab:models in the Appendix. Bronzini et al. (2023); Clarkson et al. (2020); Sharma et al. (2024) did not report baselines, which makes the study difficult to analyze. (Mehra et al., 2022) reported good accuracy, demonstrating that a few sentences hold a significant amount of information to predict ESG score changes. However, there is definitely room for improvement, indicating that the information selected is not sufficient. Lin et al. (2023a) reported the performance of the baseline (average score) reaching already mean squared error (MSE) of 11.7% already quite low. Their Transformer-based approach reached MSE of 7.4% improving slightly on classical approaches (Gradient Boosting reaching MSE of 8.2%). Detailed performances can be found in Table 24 in the Appendix.
Insights
We expected the prediction of ESG and CSR metrics to be difficult and require an extensive understanding of a company. However, Bronzini et al. (2023); Lin et al. (2023b); Mehra et al. (2022) and Clarkson et al. (2020) showed that it is possible to build strong predictors relying only on textual elements. This can be explained because analysts reward transparency (Bronzini et al., 2023), so the quantity and complexity of disclosure have strong predictive power (Clarkson et al., 2020). Although existing studies have examined specific dimensions, future research should investigate the interaction between form (Clarkson et al., 2020) and content (Bronzini et al., 2023), as this could help uncover inconsistencies that may indicate greenwashing practices.
5. Greenwashing Detection
After having discussed different subtasks of greenwashing detection, we now come to the final, all-englobing task:
Task Description
Greenwashing detection is the task of predicting if a text contains greenwashing. This is a binary classification task. Greenwashing detection can also be done at an aggregated scale (such as by a yearly indicator).
Greenhushing and Selective Disclosure
There is a large body of work on disclosure (as described in Section 4.3.1). Quantifying climate-related disclosure can highlight greenhushing and selective disclosure. Bingler et al. (2022) identify a lack of disclosure in the Strategy and Metrics&Targets categories, which they describe as cherry-picking. They also highlight that merely announcing support for the TCFD does not lead to an increase in disclosure – on the contrary, it is a practice called cheap talk. Similarly, Auzepy et al. (2023) conducted a more fine-grained analysis of TCFD-related disclosure in the banking industry. They also highlighted large differences across TCFD categories. In particular, they found a low disclosure rate on the fossil-fuel industry-related topic despite large investments in that sector.
Climate Communication as an Image-Building Strategy
Ding et al. (2023) and Chou et al. (2023) found a correlation between climate-related disclosure and carbon emission: companies that emit more tend to disclose more climate-related information. Bingler et al. (2023) proposed a Cheap talk index based on claims specificity. They reached the same conclusion that larger emitters tend to disclose more, but also that negative news coverage is correlated with cheap talk. Polignano et al. (2022) showed that while disclosing more, reports mostly focus on positive disclosure. Kang and Kim (2022) proposed a sentiment ratio metric that highlights overly positive corporate reports. They identified periods of overly positive communications that followed negative environmental controversies; in other words, companies that tried to rebuild their image after a controversy. Gorovaia and Makrominas (2025) analyzed CSR reports of environmental violators and companies with clean records. They studied an ensemble of variables: quantification of environmental content, readability score, and sentiment analysis. They found that violators publish longer, more positive, and less readable CSR reports with more environment-related content. This suggests that companies use climate disclosure as a tool to mitigate controversies and/or to improve their image. Zhao et al. (2023) propose to use the discrepancies between internal disclosure and social media perception of a company to identify potential greenwashing.
The Role of Disclosure Style in Perceived Commitment
Clarkson et al. (2020) found that companies that use a more complex language in their CSR disclosure are associated with a higher CSR rating. Schimanski et al. (2023b) concluded that a higher quantity of ESG communication is associated with higher ESG rating. Rouen et al. (2023) also highlighted the relationship between disclosure quantity and complexity with the ESG score. This might indicate that the linguistic style is a good predictor of the substance of the discourse, or that analysts are rewarding companies that communicate on ESG-related issues.
ESG Score in Greenwashing Detection
Lee and Raschke (2023) proposed a greenwashing index based on the difference between the ESG score and the ESG score, weighted by the quantity of communication on each topic (E, S and G). Sharma et al. (2024) define two types of risks: managed risk (the company is addressing it), and unmanaged risk (the company is not currently addressing it but it could). Based on these, they propose an ESG unmanaged risk score as a measure for environmental performance, which can then be used as a signal for identifying greenwashing.
Contrasting Stances: A Method to Identify Greenwashing in Climate Communications
Morio and Manning (2023) proposed to evaluate a company’s stance towards climate change mitigation policies. They hypothesize that companies with a mixed stance might be engaging in greenwashing. The hypothesis was not tested and remains unconfirmed; still, such methodology helps understand the narratives of companies around climate change. Coan et al. (2021) were able to analyze the type of contrarian claims used by conservative think-tank (CTTs) websites and contrarian blogs. They showed a shift over time from climate change denying arguments to opposing climate change mitigation policies. Despite this shift, they found that the primary donors of these think tanks continue to support organizations promoting a climate-denier narrative aimed at discrediting scientific evidence and scientists. Discrepancies between the stance extracted from official communications (Morio and Manning, 2023) on the one hand, and the stance of third-party media organizations such as think tanks founded by the company (Coan et al., 2021) on the other, might indicate greenwashing.
Defining the linguistic features of greenwashing
Oppong-Tawiah and Webster (2023) we propose to identify greenwashing using the linguistic profile of tweets. They estimate the deceptiveness of the text using a keyword-based approach and linguistic indicators (e.g. word quantity, sentence quantity). They found a correlation between greenwashing and lower financial performances. A scoring system based solely on linguistic elements has significant limitations. For example, the following tweet gets a high greenwashing score: “Read about #[company’s] commitment to a #lowcarbon future http://[company website]”. However, it is merely an invitation to read a Web page.
Insights
Several indicators for greenwashing have been proposed: overly positive sentiment (Kang and Kim, 2022), the use of specific linguistic cues (Oppong-Tawiah and Webster, 2023), discrepancies between ESG scores and corporate disclosures (Lee and Raschke, 2023; Sharma et al., 2024), ambiguous stances (Morio and Manning, 2023), and inconsistencies between social media perception and official disclosures (Zhao et al., 2023). While these approaches laid the theoretical groundwork for understanding indicators of greenwashing, they have a significant limitation: they are not empirically evaluated (Bernini et al., 2023). Only a few studies, such as that by Kang and Kim (2022), have made attempts to validate their signals against actual cases, but even these efforts have not resulted in a comprehensive dataset of greenwashing examples. This gap between theory and practice highlights the necessity of developing datasets containing real-world instances of greenwashing. Without empirical evaluation, the prediction that a company engages in greenwashing is unfounded at best, and misleading or even defamatory at worst. However, building such a dataset comes with many challenges. First, one would have to identify suitable sources of documents that are openly accessible and likely to contain greenwashing. Second, one would have to overcome the inconspicuous and subjective nature of greenwashing itself. Finally, the publication of any such dataset exposes the authors to charges of defamation by the companies they accuse of greenwashing.
6. Open Challenges in Greenwashing Detection
We discuss open challenges in the evaluation (Section 6.1), model robustness (Section 6.2), and data sources (Section 6.3).
6.1. Evaluation Methodology
Ambiguous Labels
Many tasks in greenwashing detection are framed as classification tasks. However, in some cases, the labels are not sufficiently distinguishable. For example, specificity classification (Bingler et al., 2023) is the task of distinguishing specific claims from unspecific claims. It turns out that this task is itself too unspecific, as the low inter-annotator agreement shows. However, even for tasks that seem easy to distinguish, some edge cases remain ambiguous, e.g., the distinction between “climate-related” and “not climate-related”. The climate detection model climateBERT (Bingler et al., 2023) classifies the sentence “This year, we are producing more than 10,000 [bicycle/electric bicycle], demonstrating our continued growth and commitment to meeting market demand.” as “climate-related” for bicycle but not for electric bicycle. While electric bicycles are often associated with solutions to reduce emissions, the statement is not explicitly making the link. If we consider indirect links, then non-electric bicycles are also solutions. There lies the question of which product is inherently linked to climate-change, and which one requires contextualization.
When labels in classification tasks are not clearly distinguishable, the task is not well-defined. Consequently, it does not mean much if models perform well on such a task. It merely means that the model has learned whatever label distribution the dataset had. Work in this area should thus always first check human inter-annotator agreement on the task to make sure it is meaningful.
Generated Text Evaluation
Some tasks are not classification tasks, but text generation tasks. Evaluating generated text is a challenging task (Sai et al., 2022), and different techniques are being used: BLEU, METEOR, human-based evaluation, and LLM-based evaluation (Mullappilly et al., 2023; Rony et al., 2022; Vaghefi et al., 2023; Thulke et al., 2024). BLEU and METEOR are not well-suited for evaluating tasks such as question answering  (Chen et al., 2019), as they rely only on n-gram similarity. Human evaluation also has its shortcomings. First, it is limited to smaller sizes and variations in settings. Second, human evaluation of generated texts should be based on explicit guidelines (Thulke et al., 2024; Bulian et al., 2024), but sometimes resorts to comparisons (Mullappilly et al., 2023; Thulke et al., 2024) or simple user preferences (Mullappilly et al., 2023). Third, human evaluation can be influenced by factors other than the quality of the answer – for example, the observer-expectancy effect (Rosenthal, 1976). For instance, in ChatClimate (Vaghefi et al., 2023), humans evaluate the accuracy of text generated by a generic ChatGPT vs. the proposed ChatClimate chatbot. The text produced by the ChatClimate bot always gives specific references (down to page numbers), while the generic ChatGPT does not. This allows evaluators to easily identify the more precise answer based solely on the type of references used, raising the concern that this criterion may have disproportionately influenced assessments, potentially overlooking broader aspects such as overall accuracy or quality. LLM-based evaluation has its own challenges: Language models are known to reproduce human biases (Gallegos et al., 2024). They are sensitive to the way questions are formulated, to the order of answers in multiple-choice questions (Pezeshkpour and Hruschka, 2024), and to cues that can influence human evaluation such as using “Our Model”/“Their model” (which is the wording used in Mullappilly et al. (2023)). LLM-based evaluation should thus always adhere to the same rigorous standards that are applied to human evaluation.
Figure 1. Upset plot (Lex et al., 2014) of the methods used in the studies mentioned in this literature review. Each bar counts the number of studies reporting a given set of methods. The number of a papers using a given method is reported on the horizontal bars. We have highlighted in red the intersection sets that include both fine-tuned Transformers and keyword-based methods.
Baselines
Figure 1 shows the approaches that are commonly used in greenwashing detection tasks (more details in Table LABEL:tab:models). We see that the most common solution is fine-tuning transformer models; more recent approaches include using LLMs as few-shot learners. The figure also shows which baselines have been used:
  • Random/Majority Baseline: Any proposed approach should beat a random baseline. However, we find that only 9% of papers we studied compare their results to the random or majority baseline. For example, Friederich et al. (2021) reported performances under 50% for a binary classification task, which is worse than a theoretical random classifier on a balanced dataset. Their dataset is not balanced, but it is difficult to draw conclusions without the performance of a random classifier. Random classifiers cannot just suggest that a model underperforms, but also that a model has a comparatively good performance. For example, on SciDCC or ClimaTOPIC, the performance of models is around 50-60%. While this looks mediocre, it is in fact an excellent performance as the random classifier performed around 5-7% (Spokoyny et al., 2023). In case of a heavily imbalanced dataset, the majority baseline is more meaningful. If such a baseline is not provided, but the dataset is available, it can even be computed a posteriori. Unfortunately, neither is the case for the work of Friederich et al. (2021). Similarly, a constant value baseline can help understand the performance of a regression model. Sharma et al. (2024) and Bronzini et al. (2023) evaluated the prediction of the ESG score, but do not provide such a baseline. Thus, the reader is left to wonder whether a model that always predicts the mean value might not actually perform better.
  • Simple Baselines: If a keyword-based baseline performs well, or even better than a proposed method, then the proposed method is likely an overkill. For example, Rouen et al. (2023); Clarkson et al. (2020) showed that even for the prediction of an ESG score (which is a complex target), linguistic features such as the quantity and complexity of text are good predictors. This suggests that simple predictors could perform well even on complex tasks. Indeed, simple baselines can outperform finetuned approaches, for example on TCFD classification (Sampson et al., 2022) or on the ClimaINS dataset (Spokoyny et al., 2023). Beyond a validation (or invalidation) of the proposed approaches, a keyword-based baseline would also allow us to understand if models rely on an “understanding” of the text or merely on word cues. And yet, less than half of the works we studied report the performance of a keyword-based baseline.
  • Human Baselines: Using a human as a baseline can help determine whether a task is well-defined. If even a qualified human cannot achieve high accuracy, this most likely means that the task is subjective, ill-defined, or in disagreement with the dataset. Vice versa, if a human achieves high accuracy, but the model achieves even higher accuracy, this could mean that the model has acquired expert knowledge. However, with the exception of 3 studies, all studies discussed here not come with a human evaluation.
6.2. Model Robustness
When evaluating a model, it is important to know whether the performance on a dataset is due to an inherent capacity of the model or due to random factors that may be entirely different on another dataset.
Model Performance Evaluation
Many factors can influence the performance of a model (e.g. data sampling, hyper-parameters, random initialization, model architecture). Thus, a better performance in an experiment could be due also to these factors rather than due to the choice of the model. This is even more true when studies report differences in performance of less than 1%. For this reason, the uncertainty of the evaluation should always be quantified – by confidence intervals, p-values or at least the standard deviation. However, we found that only seven papers, out of the 61 we studied, quantify the uncertainty of the model performance (Webersinke et al., 2022; Leippold and Varini, 2020; Bingler et al., 2023; Schimanski et al., 2024a, 2023b; Lee and Kim, 2023; Schimanski et al., 2023a).
Noisy Text and Adversarial Attacks
Most studies mentioned in this survey trained and evaluated models on datasets that contain only clean examples. However, real-world data is usually noisy. Thus, the models may perform entirely differently on real-world datasets. Studies should quantify not just the performance of the models, but also their robustness, as fine-tuned models are prone to word-based adversarial attacks (Woloszyn et al., 2021).
Imbalanced Datasets
The class of interest is relatively rare in most tasks. For example, only 5% of text is in the positive class “climate-related” in the study of Yu et al. (2024). This means that (1) a simple baseline that always says “non-climate-related” will have a very high precision, and (2) a model trained on such a dataset would probably overfit. Therefore, most studies mentioned in this review trained and evaluated the models on relatively balanced datasets. On the other hand, a model evaluated only on a balanced dataset may fail when tested on real-world, unbalanced datasets (Friederich et al., 2021). For example, a model reaching an F1-score of 90% on a balanced dataset would significantly drop on an unbalanced dataset with 5% of text in the positive class (to a binary F1-score of 47%, and a macro F1-score of 71%). This is due to the large number of false positives. Thus, studies should consider both the balanced and the unbalanced case.
Precision-Recall Trade-off
In most methods (and most tasks), higher precision can be traded for smaller recall. In the case of greenwashing, both metrics can be useful: a higher recall is desirable when providing leads for human analysis, so as to avoid missing potential greenwashing cases. However, if the goal is to accurately measure greenwashing prevalence, minimizing false positives is critical, as false accusations can have serious repercussions. Most studies, however, are conducted without the use case in mind.
6.3. Data Sources
Close-source data and reproducibility
The studies mentioned in our review use various data sources, most of which are publicly available: annual and sustainable reports  (Doran and Quinn, 2009; Kölbel et al., 2021; Chou et al., 2023; Huang et al., 2020; Kraus et al., 2023; Kheradmand et al., 2021; Rouen et al., 2023; Lai and Chen, 2023; Kang and Kim, 2022; Auzepy et al., 2023; Brié et al., 2022; Webersinke et al., 2022; Luccioni et al., 2020; Kheradmand et al., 2021; Bronzini et al., 2023; Lee and Kim, 2023; Polignano et al., 2022; Kraus et al., 2023), social media (Woloszyn et al., 2021; Luo et al., 2020; Vaid et al., 2022; Coan et al., 2021; Oppong-Tawiah and Webster, 2023; Sharma et al., 2024), specialized websites (Morio and Manning, 2023; Schimanski et al., 2023a; Jin et al., 2022; Leippold and Varini, 2020; Coan et al., 2021). However, multiple works rely on closed-source data: ESG scores, carbon emissions, and machine-readable reports that are collected through private providers such as Refinitiv, Reuters, RavenPack, Trucost, Bloomberg, RobecoSAM, Compustat or Wharton (Sautner et al., 2023; Bingler et al., 2023; Stammbach et al., 2023; Friederich et al., 2021; Brié et al., 2022; Lee and Raschke, 2023; Webersinke et al., 2022; Li et al., 2020; Clarkson et al., 2020; Li et al., 2020; Sautner et al., 2023; Mehra et al., 2022; Sharma et al., 2024; Clarkson et al., 2020; Schimanski et al., 2023b; Vinella et al., 2023; Kraus et al., 2023). Relying on private providers limits the reproducibility of these works, for both academic work and applications.
No greenwashing datasets
The most fundamental shortcoming of current work on greenwashing is that there is no comprehensive real-world dataset with instances of greenwashing. This limits the conclusions that can be drawn from current studies drastically. The lack of a gold standard dataset can be explained by several factors:
  • The definitions of greenwashing are often vague and ambiguous, leaving room for interpretation, and making it difficult to annotate examples of greenwashing.
  • Green claims are often not self-contained, but spread across multiple paragraphs in a same document, across multiple documents, or even across different types of sources (such as tables, databases, etc.). Therefore, just collecting all the elements required to correctly identify greenwashing is a challenge on its own.
  • Understanding the claim and annotating greenwashing requires expertise, and sometimes even debate, hence producing greenwashing annotations is a difficult task.
  • Greenwashing is a strong accusation that might impact a company reputation, and that can also, if not substantiated by evidence, be considered a defamation.
To build a dataset of real-world greenwashing cases, several data collection strategies can be considered. To avoid relying on a subjective annotation, researchers can rely on a third-party judgment. For example, using reputable news articles that explicitly discuss greenwashing to find actual examples, drawing on prior research focused on detecting media mentions of greenwashing (Gourier and Mathurin, 2024). Another valuable resource are regulatory records from authorities such as the Advertising Standards Authority (ASA) or databases like greenwash.com.
Missing link to regulatory texts
Existing work often relies on its own definitions of greenwashing. This risks introducing subjectivity and incompatibility with other works. A safer choice is to use regulatory texts as a reference, as this is done, e.g., in Stammbach et al. (2023). Recently, the European Commission adopted the corporate sustainability reporting directive (CSRD) (Union, 2022), making environmental disclosure requirements explicit, and the directive on green claims (US Environmental Protection Agency, 2023), which defines requirements when making green claims. These could be good starting points for defining greenwashing in future research. Such regulations help not just with the definition of greenwashing, but also with its judgment. Instead of emitting a judgment, a method can link a statement to a regulatory text. For example, the works that predict TCFD simply link a statement to a recommendation (Ding et al., 2023; Auzepy et al., 2023; Brié et al., 2022; Bingler et al., 2022; Sampson et al., 2022). This approach makes the methods less prone to ad hoc definitions of scores and judgment strategies, and lets a human take the final decision.
Quantifying communication vs Quantifying information
With the exception of Bronzini et al. (2023), the studies reviewed in this work all measure how often a given text contains climate-related content, TCFD disclosures, or specific topics of interest. However, a company can repeatedly emphasize the same objective or repeatedly highlight a single initiative. With this, they skew the quantity of climate-related content in their favor, while they may still fail to implement any meaningful action. This phenomenon has been called “cheap talk” (Bingler et al., 2023). Future research should thus quantify not only how often some pieces of information are conveyed, but also how many different pieces there are.
7. Conclusion
In this survey, we have provided a comprehensive overview of current research initiatives on detection of greenwashing in textual corporate communications. Determining whether a communication contains greenwashing presents three fundamental challenges. First, the absence of a commonly accepted and actionable definition of greenwashing creates ambiguity, complicating any attempt at standardized assessment. Second, there are significant legal and reputational consequences of being accused of greenwashing. Finally, there are no explicit precedents that could serve as jurisprudence for systematic labeling. This makes the annotation of actual examples of greenwashing challenging, and indeed, there is currently no large-scale dataset of instances of greenwashing. For this reason, the detection of greenwashing has remained on a theoretical level without empirical validation (Moodaley and Telukdarie, 2023a). Most existing works target intermediate tasks such as identifying climate-relatedness or green claims. This work also serves its purposes: First, determining whether a text can contain greenwashing at all helps restricting attention to the relevant resources. Second, decomposing greenwashing into more manageable steps can lead to an overall better performance, as supported by the principles behind the Chain-of-Thought (Wei et al., 2022). Third, using frugal approaches helps reduce costs and energy consumption. Finally, intermediate tasks can help humans understand the final verdict, which is important for the explainability of decisions. Several challenges remain in these works:
  • Evaluation methodology. As models are capable of tackling more complex tasks, we need tools to correctly evaluate models on these complex tasks. This include building robust evaluation methods for generated texts or building classification benchmarks that are difficult to solve but simple to evaluate. Current results often lack hyper-parameter tuning, simple baselines, and inter-annotator agreement measures, making it difficult to draw definitive conclusions from them.
  • Model robustness. When evaluating models, we need to make sure to evaluate also their robustness in real-world applications that exhibit data quality issues or highly imbalanced classes. This includes quantifying the uncertainty of the performance measures. Current works often do not provide confidence intervals or robustness assessments.
  • Quantifying information. Most of the studies we reviewed quantify the amount of text related to a topic, rather than the amount of information. Future work should quantify also the information itself.
  • Relying on regulatory texts. Most existing works identify one or more supposed characteristics of a misleading claim, but do not confront the statements to regulatory texts. As regulatory texts are becoming more precise about the requirements of environmental communications, they will increasingly become the gold standard for identifying problematic claims.
Any step toward understanding the characteristics and formalizing greenwashing is a step toward developing AI systems capable of automatically identifying these misleading communications. While previous research has laid a foundation for the theoretical understanding of greenwashing, and for the detection of signals that can indicate misleading climate-related communications, future research should focus on building a representative and diverse corpus of real-world cases of potential greenwashing to first evaluate models and, secondly, to empirically analyze its mechanisms and patterns. We hope that future research will help identify and combat greenwashing more reliably, thereby fostering transparency and accountability in climate communication.
Limitations
This survey provides a comprehensive overview of studies on automatic detection of misleading climate-related corporate communications and its related tasks as of early 2025. A number of limitations remain: Given the extensive scope of the field, our review concentrated on NLP-based research, providing only limited discussion of topics like ESG prediction, which are predominantly explored through quantitative methods. Furthermore, our review relied primarily on keyword searches on arXiv, Google Scholar, and ResearchGate, complemented by references from identified papers. In this, we may have overlooked relevant studies that use alternative terminology or are published in less accessible venues. Finally, much of greenwashing happens not on the text level, but on the factual level (with claims that contradict data) or on the multi-modal level (across images or videos). These remain out of scope for our survey.
Acknowledgment
We would like to express our sincere gratitude to Nina Godart and AEF Info for their invaluable contributions and knowledge on the topic. We also extend our thanks to Takaya Sekine CFA for his insightful comments. This work was partially supported by the grant ANR-20-CHIA-0012-01 (“NoRDF”).
Appendix
Appendix A TCFD recommendations
The recommendations are directly taken from the TCFD website: https://www.fsb-tcfd.org/recommendations/
Governance
Disclose the organization’s governance around climate-related risks and opportunities.
  • Describe the board’s oversight of climate-related risks and opportunities.
  • Describe management’s role in assessing and managing climate-related risks and opportunities.
Strategy
Disclose the actual and potential impacts of climate-related risks and opportunities on the organization’s businesses, strategy, and financial planning where such information is material.
  • Describe the climate-related risks and opportunities the organization has identified over the short, medium, and long term.
  • Describe the impact of climate-related risks and opportunities on the organization’s businesses, strategy, and financial planning.
  • Describe the resilience of the organization’s strategy, taking into consideration different climate-related scenarios, including a 2°C or lower scenario.
Risk Management
Disclose how the organization identifies, assesses, and manages climate-related risks.
  • Describe the organization’s processes for identifying and assessing climate-related risks.
  • Describe the organization’s processes for managing climate-related risks.
  • Describe how processes for identifying, assessing, and managing climate-related risks are integrated into the organization’s overall risk management.
Metrics and Targets
Disclose the metrics and targets used to assess and manage relevant climate-related risks and opportunities where such information is material.
  • Disclose the metrics used by the organization to assess climate-related risks and opportunities in line with its strategy and risk management process.
  • Disclose Scope 1, Scope 2 and, if appropriate, Scope 3 greenhouse gas (GHG) emissions and the related risks.
  • Describe the targets used by the organization to manage climate-related risks and opportunities and performance against targets.
A.1. Data sources
Source Papers
10-K reports (Doran and Quinn, 2009; Kölbel et al., 2021; Chou et al., 2023; Huang et al., 2020; Kraus et al., 2023)
Manually collected reports (Kheradmand et al., 2021; Rouen et al., 2023; Lai and Chen, 2023; Kang and Kim, 2022; Auzepy et al., 2023; Brié et al., 2022; Webersinke et al., 2022; Luccioni et al., 2020; Kheradmand et al., 2021; Bronzini et al., 2023; Lee and Kim, 2023; Polignano et al., 2022; Kraus et al., 2023)
Refinitiv (Sautner et al., 2023; Bingler et al., 2023; Stammbach et al., 2023; Friederich et al., 2021; Brié et al., 2022; Lee and Raschke, 2023; Webersinke et al., 2022)
Reuters (Li et al., 2020; Clarkson et al., 2020)
ESG score private providers (Li et al., 2020; Sautner et al., 2023; Mehra et al., 2022; Sharma et al., 2024; Clarkson et al., 2020; Schimanski et al., 2023b)
CDP (Spokoyny et al., 2023)
Open for Good Data (Vinella et al., 2023)
ClimateWatch (Kraus et al., 2023)
Twitter/Reddit (Woloszyn et al., 2021; Luo et al., 2020; Vaid et al., 2022; Coan et al., 2021; Oppong-Tawiah and Webster, 2023; Sharma et al., 2024)
LobbyMap (Morio and Manning, 2023)
Net-zero tracker (Schimanski et al., 2023a)
Fact-checking websites (Jin et al., 2022; Leippold and Varini, 2020; Coan et al., 2021)
Table 10. Data Sources
Publicly available Reports
Many studies rely on 10-K reports (Doran and Quinn, 2009; Kölbel et al., 2021; Chou et al., 2023; Huang et al., 2020; Kraus et al., 2023) as they are easily accessible (EDGAR), uniformly structured, and mandatory for companies listed on US financial markets. Other reports, such as sustainable reports, are not centralized and mandatory, therefore they need to be collected from other sources. They are often collected manually from the company’s website (Kheradmand et al., 2021; Rouen et al., 2023; Lai and Chen, 2023; Kang and Kim, 2022; Auzepy et al., 2023; Brié et al., 2022; Webersinke et al., 2022), or from other websites (Luccioni et al., 2020; Kheradmand et al., 2021; Bronzini et al., 2023; Lee and Kim, 2023; Polignano et al., 2022; Kraus et al., 2023) such as from the GRI, SASB, TCFD, ResponsibilityReports, AnnualReports, Korean Standards Association. Polignano et al. (2022) relied on Sustainability Disclosure Database from GRI. To the best of our knowledge, it is not available any longer.
The reports are openly available, however they are most of the time in a PDF format, and must be parsed to produce machine-readable versions. While Khan (2024) proposed a machine-readable dataset for the 10-K reports, it is often not the case for other types of reports. (Bronzini et al., 2023; Polignano et al., 2022; Kang and Kim, 2022) provide a sample of processed reports.
Private Providers
Machine-readable reports, as well as earning call transcripts are often provided through a private data provider such as Refinitiv (Sautner et al., 2023; Bingler et al., 2023; Stammbach et al., 2023; Friederich et al., 2021; Brié et al., 2022; Lee and Raschke, 2023; Webersinke et al., 2022) or Reuters (Li et al., 2020; Clarkson et al., 2020). Similarly, financial data (e.g. ESG score) is often collected through data providers such as RavenPack or S&P Global Trucost or Wharton’s research platform WRDS, Reuters, Bloomberg, Refinitiv, RobecoSAM and Compustat (Li et al., 2020; Sautner et al., 2023; Mehra et al., 2022; Sharma et al., 2024; Clarkson et al., 2020; Schimanski et al., 2023b)
Carbon emission data
Carbon emissions of companies can be collected through the private data providers, or through CDP (Spokoyny et al., 2023), Open for Good Data (Vinella et al., 2023) or ClimateWatch website (Kraus et al., 2023).
Media and Social Media
As companies communicate with the public mainly through marketing and social media, other data sources include Twitter and Reddit (Woloszyn et al., 2021; Luo et al., 2020; Vaid et al., 2022; Coan et al., 2021; Oppong-Tawiah and Webster, 2023; Sharma et al., 2024).
Climate expert websites
Finally, projects such as LobbyMap or Net Zero Tracker, that track corporate policy and engagements, can help build annotated datasets (Morio and Manning, 2023; Schimanski et al., 2023a). Similarly, relying on a fact-checking website such as https://climatefeedback.org/ or www.skepticalscience.com is an efficient way of building a dataset (Jin et al., 2022; Leippold and Varini, 2020; Coan et al., 2021)
Appendix B Detailed model list
Table 11. Solutions proposed in each paper
Paper Dataset Baselines BERT Sent. Sim. Other
Pretraining Models on Climate-Related Text
Webersinke et al. (2022) Domain-specific corpus - Pre-Training ClimateBERT
Vaghefi et al. (2022) Deep Climate Change - Pre-Training ClimateGPT-2
Mullappilly et al. (2023) Deep Climate Change - Pre-Training stsb-xlm-r-multilingual Arabic Mini-ClimateGPT, Vicuna, Alpaca, Dolly V2, Alpaca ar
Schimanski et al. (2023b) Domain-specific corpus - Pre-Training EnvironmentalBERT
Yu et al. (2024) climateBUG-data - Pre-Training climateBUG-LM
Thulke et al. (2024) Domain-specific corpus - Pre-Training ClimateGPT-7/13/70B
Thulke et al. (2024) ClimaBENCH - Evaluation Stability-3B, Pythia-6.9B, Falcon-7B, Mistral-7B, Llama-2-7B, Jais-13B, Jais-13B-Chat, Llama-2-7/13/70B, Llama-2-Chat-7/13/70B, ClimateGPT-7/13/70B
Climate-related topic detection
Doran and Quinn (2009) 10-K filings - climate Keywords
Engle et al. (2020) WSJ - climate TF-IDF cosine similarity
Leippold and Varini (2020) climatext NB, Keywords BERT
Garrido-Merchán (Garrido-Merchán et al., 2023) climatext ClimateBERT, BERT
Spokoyny et al. (2023) climatext Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
Webersinke et al. (2022) climate detection distilRoBERTa, climateBERT
Bingler et al. (2023) climate detection NB, SVM + BoW, SVM + ELMo climateBERT
Brié et al. (2022) climate detection Random Forest climateBERT, FinBERT
Yu et al. (2024) climateBUG data BERT, FinBERT, ClimateBERT, climateBUG-LM
Lin et al. (2023b) sustainable signals review
Climate-related subtopic detection
Kang and Kim (2022) Sustainable Reports - topics sentence-BERT
Mishra and Mittal (2021) sciDCC
Spokoyny et al. (2023) sciDCC Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
Vaid et al. (2022) climateEng BERT, RoBERTa, DistilBERT FastText
Spokoyny et al. (2023) climateEng Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
Jain et al. (2023) ledger - topics TF-IDF BERT, ClimateBERT and RoBERTa. sentence-BERT Word2Vec-based Random Forest
Schimanski et al. (2024a) ESGBERT-nature Keywords ClimateBERT, RoBERTa, DistilRoBERTa, EnvironmentalBERT
Spokoyny et al. (2023) ClimaTOPIC Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
Detecting climate-related financial disclosure
Ding et al. (2023) Reports - Disclosure TF-IDF cosine similarity
Auzepy et al. (2023) Reports - TCFD BART+MNLI
Brié et al. (2022) Sustainable Reports structural topic model (STM)
Bingler et al. (2022) climate tcfd rec. TF-IDF + random forest RoBERTa, ClimateBERT (RoBERTa+LogReg) Sentence Enc.
Sampson et al. (2022) Reports - TCFD TF-IDF ClimateBERT, distilRoBERTa kNN (sentence-transformers embeddings), stacked models
In-depth Disclosure: Climate Risk Classification
Li et al. (2020) Reports - Risk dictionaries (risk)
Chou et al. (2023) 10-K filings - Risk 3 dictionaries (risk: Climate, Impact, Litigious)
Kheradmand et al. (2021) Annual/Sustainability Reports - Risk dictionaries (risk)
Sautner et al. (2023) earning calls - Risk dictionaries (risk/opportunity)
Polignano et al. (2022) Sustainability Reports - Sentiment TextBlob, Sentita
Kang and Kim (2022) Sustainability Reports - Sentiment DistilBERT (sentiment analysis)
Kölbel et al. (2021) 10-K filings - Risk BoW, TF-IDF BERT
Friederich et al. (2021) Reports - Risk SVM distilBERT, RoBERTa
Bingler et al. (2023) climate sentiment NB, SVM + BoW, SVM + ELMo distilRoBERTa, climateBERT
Webersinke et al. (2022) climate sentiment distilRoBERTa, climateBERT
Xiang and Fujii (2023) Web of Science, Scopus - Sentiment BERT, Domain-specific BERT, TinyBERT, DistilBERT, ClimateBERT CCLA+Max-Pooling, Bi-LSTM-Attention, POS-Bi-LSTM-Attention, DARE
Detecting environmental, social, and governance disclosure
Rouen et al. (2023) Sustainability Reports - ESG industry-topic dictionaries
Mehra et al. (2022) Sustainability Reports - ESG sentence BERT, Universal Sentence Encoder, and Deep Averaging Network
Huang et al. (2020) Sustainability Report, 10-K filings - ESG NB, SVM, RF FinBERT, BERT CNN, LSTM
Schimanski et al. (2023b) ESGBERT E,S,G BERT, RoBERTa, DistilRoBERTa, E/S/G-RoBERTa, E/S/G-DistilRoBERTa
Lee and Kim (2023) Korean Reports - ESG Multilingual BERT, KoBERT, DistilKoBERT, KB-ALBERT, KLUE-RoBERTa
Bronzini et al. (2023) Sustainability Reports - ESG INSTRUCTOR-xl
Green claim detection
Woloszyn et al. (2021) green claims RoBERTa, BERTweet Flair
Stammbach et al. (2023) environmental claims Random, Majority, TF-IDF + SVM, Character n-gram SVM DistilBERT, ClimateBERT, RoBERTa ClaimBuster RoBERTa, Pledge Detection RoBERTa
Lin et al. (2023b) Product Description - Claim environmental claim detection (finetuned distilBERT)
Green claim characteristics
Bingler et al. (2023) climate-commitments-actions NB, SVM + BoW, SVM + ELMo distilRoBERTa, climateBERT
Schimanski et al. (2023a) netzero/reduction ClimateBERT, DistilRoBERTa, RoBERTa GPT-3.5-turbo
Woloszyn et al. (2021) green claims 3 RoBERTa, BERTweet Flair
Bingler et al. (2023) climate specificity NB, SVM + BoW, SVM + ELMo distilRoBERTa, climateBERT
Green stance detection
Morio (Morio and Manning, 2023) LobbyMap Majority, TF-IDF BERT, RoBERTa, ClimateBERT, Longformer
Luo et al. (2020) DeSMOG
Vaid et al. (2022) ClimateStance BERT, RoBERTa, DistilBERT FastText
Lai and Chen (2023) Reports - Stance BERT, RoBERTa, ClimateBERT
Spokoyny et al. (2023) ClimateStance Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
Spokoyny et al. (2023) DeSMOG Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
Leippold (Leippold and Diggelmann, 2020) ClimateFEVER ALBERT(FEVER)
Wang et al. (2021) ClimateFEVER RoBERTa (ClimateFEVER) RoBERTa (FEVER), RoBERTa (ClimateFEVER) + UDA
Spokoyny et al. (2023) ClimateFEVER Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
Vaghefi et al. (2022) ClimateFEVER GPT-2, climateGPT-2
Xiang and Fujii (2023) ClimateFEVER BERT, RoBERTa, DistilRoBERTa, ClimateBERT, POS-Bi-LSTM-Attention
Webersinke et al. (2022) ClimateFEVER distilRoBERTa, climateBERT
Vaid et al. (2022) ClimateFEVER BERT, RoBERTa, DistilBERT FastText
Question Answering
Luccioni et al. (2020) Reports - QA RoBERTa
Spokoyny et al. (2023) ClimaQA BM25 MiniLM
Spokoyny et al. (2023) ClimaINS Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
Rony et al. (2022) CCMRC - QA pairs ALBERT (trained on SQuAD then finetuned) Sentence-BERT
Schimanski et al. (2023a) Net-Zero Tracker - Target RoBERTa Quad2
Ni et al. (2023) Reports - RAG ChatGPT
Vaghefi et al. (2023) Reports - RAG - Human Eval GPT-4
Mullappilly et al. (2023) Clima500-Instruct - QA instructions Vicuna (Mini-ClimateGPT)
Thulke et al. (2024) Instruction following corpus - Human Eval ClimateGPT-70B, ClimateGPT-7B, ClimateGPT-FSC-7B
Thulke et al. (2024) ClimRetrieve ColBERTv2, DRAGON, GTE-base, text-embedding-ada-002, text-embedding-3-small, text-embedding-3-large BM25
Deceptive Techniques
Oppong (Oppong-Tawiah and Webster, 2023) Tweets - Fake News linguistic indicators
Coan et al. (2021) contrarian claims Logistic, SVM RoBERTa ULMFiT (over sample, weighted, focal Loss, logisitic, SVM), RoBERTa + Logistic
Jin et al. (2022) logicClimate ELECTRA, BERT lectra-StructureAware
Bhatia et al. (2021) climate change scepticism - Neutralization SVM BERT MTEXT
Environmental performance prediction
Jain et al. (2023) Ledger - Scope 3 Emissions Alpaca, WizardLM, Flan-T5, ChatGPT
Bronzini et al. (2023) Sustainable Reports - ESG OLS
Mehra et al. (2022) Reports 10-Qs - ESG Majority BERT, ESGBERT
Clarkson et al. (2020) CSR Reports - CSR performer linguistic features + random forest, XGBoost, SVM
Lin et al. (2023b) Product Description / Review - Finch Score Lasso, Gradient Boosting NOCATE + transformer, SUSTAINABLESIGNALS + transformer
Sharma et al. (2024) Multimodel Tweet - ESG CLIP+MLP, Sentence-BERT+MLP
Appendix C Reported Performances
We have reported all the performances computed in each study when available. We did not include the performance table if the original study did not perform an evaluation (for example, in zero-short/unsupervised settings) or if the evaluation is a heavily modified metric (for example, when using a quantitative measure alongside other signals to compute a financial factor). We only reported traditional NLP metrics. When available, we reported the F1-score, however, if the F1-score is not computed in the study, we reported the available metrics. The computation methods of the F1-score (macro, micro, weighted) are specified when enough information is given to state it confidently. We reported one table per dataset, which might contain multiple tasks, but multiple studies can tackle the same dataset and are then grouped to be able to compare the results.
C.1. Climate-related Topic Classification
Source Model F1
Webersinke et al. (2022) distilRoBERTa 98.6% ±$\pm 1$
Webersinke et al. (2022) ClimateBERT 99.1% ±$\pm 0.8$
Bingler et al. (2023) Naive Bayes 87%
Bingler et al. (2023) SVM + BoW 87%
Bingler et al. (2023) SVM + ELMo 89%
Bingler et al. (2023) ClimateBERT 97%
Brié et al. (2022) ClimateBERT 98.59%
Brié et al. (2022) FinBERT 96.67%
Brié et al. (2022) Random Forest 91.65%
(a) F1-scores (weighted average) on ClimateBERT’s climate detection dataset (Bingler et al., 2023). For Webersinke et al. (2022) scores are displayed as mean ±$\pm std$. Bingler et al. (2023) also reported an uncertainty measure (p-value) which is not reported in this table for simplicity.
Source Model F1
Yu et al. (2024) BERT 90.81%
Yu et al. (2024) FinBERT 90.82%
Yu et al. (2024) ClimateBERT 91.07%
Yu et al. (2024) climateBUG-LM 91.36%
(b) F1-scores (macro average) on ClimateBUG-data (Yu et al., 2024).
Source Model Wikipedia 10-K Claims
F1 (bin.) F1 (bin.) F1 (bin.)
Leippold and Varini (2020) best BERT 80% ±$\pm 6$ 95% $\pm$ 2 83% $\pm$ 1
Leippold and Varini (2020) best NB 60% ±$\pm 7$ 90% ±$\pm 3$ 72% ±$\pm 2$
Leippold and Varini (2020) best Keywords 67% ±$\pm 7$ 88% ±$\pm 3$ 70% ±$\pm 1$
Leippold and Varini (2020) BERT-AL-wiki 69% ±$\pm 6$ 71% ±$\pm 4$ 81% ±$\pm 1$
Garrido-Merchán (Garrido-Merchán et al., 2023) BERT-AL-wiki - 91% ±$\pm 0.6$
Garrido-Merchán (Garrido-Merchán et al., 2023) climateBERT-AL-wiki - 93% ±$\pm 0.7$
Combined - F1 (macro)
Spokoyny et al. (2023) Majority 42.08%
Spokoyny et al. (2023) Random 46.86%
Spokoyny et al. (2023) SVM 83.39%
Spokoyny et al. (2023) BERT 87.04%
Spokoyny et al. (2023) RoBERTa 85.97%
Spokoyny et al. (2023) DistilRoBERTa 86.06%
Spokoyny et al. (2023) Longformer 87.80%
Spokoyny et al. (2023) SciBERT 83.29%
Spokoyny et al. (2023) ClimateBERT 85.14%
(c) F1-scores on ClimaText (Leippold and Varini, 2020). For Leippold and Varini (2020) scores are displayed as mean ±$\pm std$.
Table 12. Reported performances for the datasets on climate-related topic detection (which include climate, sustainability and environment).
C.2. Thematic Analysis
Source Model Prec.
Bingler et al. (2022) TF-IDF+random forest 24%
Bingler et al. (2022) Sentence Enc. 23%
Bingler et al. (2022) RoBERTa Para. 22%
Bingler et al. (2022) RoBERTa Sent. 75%
Bingler et al. (2022) ClimateBERT* 84%
Bingler et al. (2022) ClimateBERT* blahblF1: 84%
(a) Average Precisions and F1-score (macro average) on TCFD recommendations classification (Bingler et al., 2022). (*They refer to a finetuned RoBERTa with a Logistic Regression as ClimateBERT)
Source Model ROC-AUC
Sampson et al. (2022) TF-IDF (1) 0.867
Sampson et al. (2022) kNN (2) 0.847
Sampson et al. (2022) ClimateBERT (3) 0.852
Sampson et al. (2022) distilRoBERTa (4) 0.819
Sampson et al. (2022) (1) + (2) 0.884
Sampson et al. (2022) (1) + (3) 0.865
Sampson et al. (2022) (1) + (4) 0.857
(b) ROC-AUC scores on TCFD classification task (Sampson et al., 2022)
Source Model F1
Auzepy et al. (2023) BART+MNLI 56.68%
(c) F1-scores (macro average) on TCFD classification (Auzepy et al., 2023)
Table 13. Reported performances for the datasets on TFCD classification.
Source Model F1
Huang et al. (2020) FinBERT 89.6%
Huang et al. (2020) BERT 86.9%
Huang et al. (2020) NB 75.0%
Huang et al. (2020) SVM 75.0%
Huang et al. (2020) RF 78.3%
Huang et al. (2020) CNN 81.9%
Huang et al. (2020) LSTM 86.0%
(a) F1-scores (micro average) on ESG classification (Huang et al., 2020)
Source Model Accuracy
Lee and Kim (2023) Multiling. BERT 84.09% ±$\pm 0.092$
Lee and Kim (2023) KoBERT 85.98% ±$\pm 0.163$
Lee and Kim (2023) DistilKoBERT 84.89% ±$\pm 0.081$
Lee and Kim (2023) KB-ALBERT 85.62% ±$\pm 0.026$
Lee and Kim (2023) KLUE (base) 86.05% ±$\pm 0.125$
Lee and Kim (2023) KLUE (large) 86.66% ±$\pm 0.099$
(b) Accuracy on ESG classification (Lee and Kim, 2023). Accuracy reported as mean ±$\pm std$
Environment F1
Schimanski et al. (2023b) BERT 91.85% ±$\pm 1.25$
Schimanski et al. (2023b) RoBERTa 92.35% ±$\pm 2.29$
Schimanski et al. (2023b) DistilRoBERTa 90.97% ±$\pm 2.00$
Schimanski et al. (2023b) EnvRoBERTa 93.19% ±$\pm 1.40$
Schimanski et al. (2023b) EnvDistilRoBERTa 92.35% ±$\pm 1.65$
Social
Schimanski et al. (2023b) BERT 89.84% ±$\pm 0.79$
Schimanski et al. (2023b) RoBERTa 89.87% ±$\pm 1.35$
Schimanski et al. (2023b) DistilRoBERTa 90.59% ±$\pm 1.03$
Schimanski et al. (2023b) SocRoBERTa 91.90% ±$\pm 1.79$
Schimanski et al. (2023b) SocDistilRoBERTa 91.24% ±$\pm 1.86$
Governance
Schimanski et al. (2023b) BERT 76.86% ±$\pm 1.77$
Schimanski et al. (2023b) RoBERTa 77.03% ±$\pm 1.82$
Schimanski et al. (2023b) DistilRoBERTa 76.65% ±$\pm 2.39$
Schimanski et al. (2023b) GovRoBERTa 78.48% ±$\pm 2.62$
Schimanski et al. (2023b) GovDistilRoBERTa 78.86% ±$\pm 1.59$
(c) F1-scores (unspecified) on environment, social, and governance datasets (Schimanski et al., 2024a). F1-scores are reported as mean ±$\pm std$
Table 14. Reported performances for the datasets on ESG classification
Source Model Water Forest Biodiv. Nature
Schimanski et al. (2024a) EnvironmentalBERT 94.47% ±$\pm 1.37$ 95.37% ±$\pm 0.92$ 92.76% ±$\pm 1.01$ 94.19% ±$\pm 0.81$
Schimanski et al. (2024a) ClimateBERT 95.10% ±$\pm 1.13$ 95.34% ±$\pm 0.94$ 92.49% ±$\pm 1.03$ 93.50% ±$\pm 0.64$
Schimanski et al. (2024a) RoBERTa 94.55% ±$\pm 0.86$ 94.78% ±$\pm 0.48$ 92.46% ±$\pm 1.54$ 93.97% ±$\pm 0.26$
Schimanski et al. (2024a) DistilRoBERTa 94.98% ±$\pm 1.16$ 95.29% ±$\pm 0.65$ 92.29% ±$\pm 1.23$ 93.55% ±$\pm 0.72$
Schimanski et al. (2024a) Keywords - - 63.03% 61.00%
(a) F1-scores (unspecified) on Nature-related topic classification (Schimanski et al., 2024a). F1-scores are reported as mean ±$\pm std$.
Source Model F1
Spokoyny et al. (2023) Majority 0.79%
Spokoyny et al. (2023) Random 5.05%
Spokoyny et al. (2023) SVM 48.02%
Spokoyny et al. (2023) BERT 54.74%
Spokoyny et al. (2023) RoBERTa 52.90%
Spokoyny et al. (2023) DistilRoBERTa 51.13%
Spokoyny et al. (2023) Longformer 54.79%
Spokoyny et al. (2023) SciBERT 51.83%
Spokoyny et al. (2023) ClimateBERT 52.97%
(b) F1-scores (macro average) on SciDCC (Mishra and Mittal, 2021).
Source Models F1
Spokoyny et al. (2023) Majority 3.65%
Spokoyny et al. (2023) Random 6.45%
Spokoyny et al. (2023) SVM 58.34%
Spokoyny et al. (2023) BERT 64.64%
Spokoyny et al. (2023) RoBERTa 65.22%
Spokoyny et al. (2023) DistilRoBERTa 63.61%
Spokoyny et al. (2023) Longformer 64.03%
Spokoyny et al. (2023) SciBERT 63.62%
Spokoyny et al. (2023) ClimateBERT 64.24%
(c) F1-scores (macro average) on the climaTOPIC (Spokoyny et al., 2023).
Source Model F1
Jain et al. (2023) all-mpnet-base-v2 43.7%
Jain et al. (2023) TF-IDF 69%
Jain et al. (2023) Word2Vec 72%
Jain et al. (2023) RoBERTa 87.19%
Jain et al. (2023) BERT 87.14%
Jain et al. (2023) ClimateBERT 85.20%
(d) F1-scores (weighted average) on ledger transaction classification (Jain et al., 2023).
Source Model F1
Vaid et al. (2022) FastText 63.8%
Vaid et al. (2022) BERT-Base 69.6%
Vaid et al. (2022) BERT-Large 69.5%
Vaid et al. (2022) RoBERTa-Base 73.4%
Vaid et al. (2022) RoBERTa-Large 73.5%
Vaid et al. (2022) DistilBERT 69.4%
Spokoyny et al. (2023) Majority 13.83%
Spokoyny et al. (2023) Random 16.71%
Spokoyny et al. (2023) SVM 51.81%
Spokoyny et al. (2023) BERT 71.78%
Spokoyny et al. (2023) RoBERTa 74.58%
Spokoyny et al. (2023) DistilRoBERTa 72.33%
Spokoyny et al. (2023) Longformer 72.28%
Spokoyny et al. (2023) SciBERT 70.50%
Spokoyny et al. (2023) ClimateBERT 71.83%
(e) F1-scores (macro average) on ClimateEng (Vaid et al., 2022).
Table 15. Reported performances for the datasets on climate-related topic detection (other topics.)
C.3. In-depth Disclosure: Climate Risk Classification
Binary 2 Classes 5 Classes
Source Model REALISTIC REALISTIC REALISTIC HARD NEG. DISCRIMINATORY
Friederich et al. (2021) SVM 29.0% 35.1% 20.4% 45.7% 59.9%
Friederich et al. (2021) distilBERT 44.4% 49.7% 24.1% 43.1% 55.8%
Friederich et al. (2021) RoBERTa 49.6% 44.6% 35.6% 52.8% 59.6%
(a) F1-scores (macro average) for climate risk disclosure Friederich et al. (2021).
Source Model F1
Xiang and Fujii (2023) BERT-Base 93.1%
Xiang and Fujii (2023) BERT-Base (Domain-spe.) 95.5%
Xiang and Fujii (2023) TinyBERT 87.0%
Xiang and Fujii (2023) DistilBERT 89.9%
Xiang and Fujii (2023) ClimateBERT 87.5%
Xiang and Fujii (2023) CCLA+Max-Pooling 82.9%
Xiang and Fujii (2023) Bi-LSTM-Attention 71.9%
Xiang and Fujii (2023) POS-Bi-LSTM-Attention 88.2%
Xiang and Fujii (2023) DARE 89.4%
(b) F1-scores (macro average) on Risk/Opportunity classification (Xiang and Fujii, 2023).
Source Model F1
Webersinke et al. (2022) distilRoBERTa 82.5% $\pm$ 4.6
Webersinke et al. (2022) climateBERT 83.8% $\pm$ 3.6
Bingler et al. (2023) Naive Bayes 72%
Bingler et al. (2023) SVM + BoW 72%
Bingler et al. (2023) SVM + ELMo 75%
Bingler et al. (2023) climateBERT 80%
(c) F1-scores (weighted average) on risk/opportunity detection Webersinke et al. (2022); Bingler et al. (2023). Webersinke et al. (2022)’s F1-scores are reported as mean ±$\pm std$. Bingler et al. (2023) also reported an uncertainty measure (p-value) which is not shown in this table for simplicity.
Source Model F1
Kölbel et al. (2021) Bag-of-Words 83.63%
Kölbel et al. (2021) TF-IDF 80.05%
Kölbel et al. (2021) BERT 94.66%
(d) F1-scores (unspecified) for risk classification (General/Transition risk/Physical Risk) (Kölbel et al., 2021).
Table 16. Reported performances for the datasets on risk detection and classification.
C.4. Green Claim Detection
Source Model F1
Stammbach et al. (2023) Majority baseline 00.0%
Stammbach et al. (2023) Random baseline 33.5%
Stammbach et al. (2023) ClaimBuster RoBERTa 33.9%
Stammbach et al. (2023) Pledge Detection RoBERTa 26.4%
Stammbach et al. (2023) TF-IDF SVM 69.1%
Stammbach et al. (2023) Character n-gram SVM 70.9%
Stammbach et al. (2023) DistilBERT 83.7%
Stammbach et al. (2023) ClimateBERT 83.8%
Stammbach et al. (2023) RoBERTa_base 82.4%
Stammbach et al. (2023) RoBERTa_large 84.9%
(a) F1-scores (binary) on environmental claim detection (Stammbach et al., 2023).
Source Model F1
Woloszyn et al. (2021) Bin RoBERTa Bal 92.08%
Woloszyn et al. (2021) Bin BERTweet Bal 91.48%
Woloszyn et al. (2021) Bin Flair Bal 88.52%
Woloszyn et al. (2021) Bin RoBERTa Unbal 88.88%
Woloszyn et al. (2021) Bin BERTweet Unbal 84.30%
Woloszyn et al. (2021) Bin Flair Unbal 81.55%
(b) F1-scores (binary) on green claim detection (Woloszyn et al., 2021).
Table 17. Reported performances for the datasets on green claim detection.
C.5. Green Claim Characteristics
Source Model F1
Woloszyn et al. (2021) Mult RoBERTa Bal. 81.45%
Woloszyn et al. (2021) Mult BERTweet Bal. 75.14%
Woloszyn et al. (2021) Mult Flair Bal. 69.96%
Woloszyn et al. (2021) Mult RoBERTa Unbal. 76.82%
Woloszyn et al. (2021) Mult BERTweet Unbal. 59.71%
Woloszyn et al. (2021) Mult Flair Unbal. 38.15%
(a) F1-scores (unspecified) on Implicit/Explicit green claim detection (Woloszyn et al., 2021).
Source Model Accuracy
Schimanski et al. (2023a) ClimateBERT 96.2% ±$\pm 0.4$
Schimanski et al. (2023a) DistilRoBERTa 94.4% ±$\pm 0.7$
Schimanski et al. (2023a) RoBERTa-base 95.8% ±$\pm 0.6$
Schimanski et al. (2023a) GPT-3.5-turbo 92.0%
(b) Accuracy on net-zero/reduction classification (Schimanski et al., 2023a). Accuracy is reported as mean ±$\pm std$.
Source Model Commit. & Act. Specificity
Bingler et al. (2023) Naive Bayes 75% 75%
Bingler et al. (2023) SVM + BoW 76% 75%
Bingler et al. (2023) SVM + ELMo 79% 76%
Bingler et al. (2023) climateBERT 81% 77%
(c) F1-scores (weighted average) on Commitment&Action and Specificity classification (Bingler et al., 2023). Bingler et al. (2023) also reported an uncertainty measure (p-value) which is not reported in this table for simplicity.
Table 18. Reported performances for the datasets on claim characteristics classification.
C.6. Green Stance Detection
Source Model F1
Vaghefi et al. (2022) GPT-2 67%
Vaghefi et al. (2022) climateGPT-2 72%
Claim
Leippold and Diggelmann (2020) ALBERT(FEVER) 32.85%
Xiang and Fujii (2023) BERT 80.7%
Xiang and Fujii (2023) RoBERTa 72.3%
Xiang and Fujii (2023) DistilRoBERTa 72.0%
Xiang and Fujii (2023) ClimateBERT 76.8%
Xiang and Fujii (2023) POS-Bi-LSTM-Att. 79.3%
Webersinke et al. (2022) DistilRoBERTa 74.8% ±$\pm 3.6$
Webersinke et al. (2022) ClimateBERT 75.7% ±$\pm 4.4$
Subset 95 examples
Wang et al. (2021) RoBERTa (FEVER) 49.2%
Wang et al. (2021) + ClimateFEVER 64.2%
Wang et al. (2021) + UDA ratio 18 71.8%
Evidence
Spokoyny et al. (2023) Majority 26.08%
Spokoyny et al. (2023) Random 30.62%
Spokoyny et al. (2023) BERT 62.47%
Spokoyny et al. (2023) RoBERTa 60.74%
Spokoyny et al. (2023) DistilRoBERTa 61.54%
Spokoyny et al. (2023) Longformer 60.82%
Spokoyny et al. (2023) SciBERT 62.68%
Spokoyny et al. (2023) ClimateBERT 61.54%
(a) F1-score (macro average) on ClimateFEVER Leippold and Diggelmann (2020). The prediction can be made at Claim level or for each Claim-Evidence pair (Evidence). Webersinke et al. (2022)’s F1-scores are reported as mean ±$\pm std$.
Source Model Acc. F1
Luo et al. (2020) Majority class 0.43 17%
Luo et al. (2020) Linear 0.62 60%
Luo et al. (2020) BERT 0.75 73%
Luo et al. (2020) Human 0.71 -
(b) Accuracy and F1-scores (macro average) on GWSD (Luo et al., 2020).
Source Model F1
Vaid et al. (2022) FastText 34.3%
Vaid et al. (2022) BERT-Base 46.4%
Vaid et al. (2022) BERT-Large 48.9%
Vaid et al. (2022) RoBERTa-Base 51.0%
Vaid et al. (2022) RoBERTa-Large 48.9%
Vaid et al. (2022) DistilBERT 44.8%
Spokoyny et al. (2023) Majority 29.68%
Spokoyny et al. (2023) Random 25.52%
Spokoyny et al. (2023) SVM 42.92%
Spokoyny et al. (2023) BERT 55.37%
Spokoyny et al. (2023) RoBERTa 59.69%
Spokoyny et al. (2023) DistilRoBERTa 52.51%
Spokoyny et al. (2023) Longformer 34.68%
Spokoyny et al. (2023) SciBERT 48.67%
Spokoyny et al. (2023) ClimateBERT 52.84%
(c) F1-scores (macro average) on ClimateStance (Vaid et al., 2022).
Source Model macro bin.
Baselines
Lai and Chen (2023) BERT - 58%
Lai and Chen (2023) RoBERTa - 89%
Lai and Chen (2023) ClimateBERT 90% 90%
SeqGAN
Lai and Chen (2023) BERT - 83%
Lai and Chen (2023) RoBERTa - 87%
Lai and Chen (2023) ClimateBERT 90% 90%
MaliGAN
Lai and Chen (2023) BERT - 74%
Lai and Chen (2023) RoBERTa - 88%
Lai and Chen (2023) ClimateBERT 91% 91%
RankGAN
Lai and Chen (2023) BERT - 82%
Lai and Chen (2023) RoBERTa - 87%
Lai and Chen (2023) ClimateBERT 91% 90%
(d) F1-scores (binary and macro average) on stance on climate change remediation efforts) (Lai and Chen, 2023).
Table 19. Reported performance for the datasets on stance detection (Part 1)
Document Page overlap Strict
Source Model P Q S P Q S P Q S
Morio and Manning (2023) Most-frequent 46.7% 52.6% 36.8% 51.8% 25.6% 19.8% 41.2% 19.6% 17.5%
Morio and Manning (2023) Linear 66.0% 61.9% 50.3% 71.4% 44.5% 36.1% 52.0% 31.2% 27.0%
Morio and Manning (2023) BERT-base 71.0% 63.5% 51.6% 73.6% 48.1% 37.2% 50.2% 31.9% 25.8%
Morio and Manning (2023) ClimateBERT 71.8% 64.0% 52.8% 74.4% 48.9% 39.0% 50.2% 32.2% 26.8%
Morio and Manning (2023) RoBERTa-base 71.6% 64.5% 53.1% 73.8% 49.6% 38.3% 50.4% 33.4% 26.6%
Morio and Manning (2023) Longformer-base 73.7% 66.9% 54.6% 75.9% 53.0% 40.8% 52.5% 36.1% 28.6%
Morio and Manning (2023) Longformer-large 73.9% 68.8% 57.3% 76.5% 55.0% 44.1% 53.6% 38.7% 31.5%
(a) Custom metrics on LobbyMap (Morio and Manning, 2023).
Table 20. Reported performance for the datasets on stance detection (Part 2)
C.7. Question Answering
Source embeddings Quest. Def. Concepts Gen. inf_3 inf_all
Schimanski et al. (2024b) random 0.037 0.037 0.037 0.037 0.037 0.037
Schimanski et al. (2024b) BM25 0.113 0.114 0.126 0.139 0.172 0.174
Schimanski et al. (2024b) ColBERTv2 0.109 0.094 0.112 0.137 0.130 0.137
Schimanski et al. (2024b) DRAGON+ 0.139 0.121 0.161 0.160 0.161 0.160
Schimanski et al. (2024b) GTE-base 0.161 0.153 0.171 0.154 0.171 0.174
Schimanski et al. (2024b) text-embedding-ada-002 0.163 0.143 0.152 0.163 0.176 0.176
Schimanski et al. (2024b) text-embedding-3-small 0.163 0.143 0.152 0.163 0.179 0.179
Schimanski et al. (2024b) text-embedding-3-large 0.167 0.143 0.152 0.163 0.179 0.179
(a) F1-score (custom) of the different embedding models and information retrieval approaches (question, definition, concepts, generic, inf_3, inf_all) (Schimanski et al., 2024b)
Test In-house annotation
Source Model F1 BLEU METEOR F1 BLEU METEOR
Rony et al. (2022) ALBERT 81.6% 0.678 0.808 66.1% 0.416 0.694
(b) F1-score, BLEU and METEOR on CCMRC (Rony et al., 2022)
Source Model F1
Luccioni et al. (2020) large 85.5%
Luccioni et al. (2020) base 82%
(c) F1-scores (macro average) of RoBERTa on question answering (Luccioni et al., 2020).
Source Model F1
Spokoyny et al. (2023) Majority 4.11%
Spokoyny et al. (2023) Random 12.14%
Spokoyny et al. (2023) SVM 86.00%
Spokoyny et al. (2023) BERT 84.57%
Spokoyny et al. (2023) RoBERTa 85.61%
Spokoyny et al. (2023) DistilRoBERTa 84.38%
Spokoyny et al. (2023) Longformer 84.35%
Spokoyny et al. (2023) SciBERT 84.43%
Spokoyny et al. (2023) ClimateBERT 84.80%
(d) F1-scores (macro average) on ClimaINS Spokoyny et al. (2023)
Source Model Question Answer
R@1 R@5 R@1 R@5
Thulke et al. (2024) bge-base-en-v1.5 54.8 71.5 81.8 92.1
Thulke et al. (2024) bge-large-en-v1.5 55.8 73.6 83.3 93.1
Thulke et al. (2024) gtr-t5-large 48.8 67.4 79.6 90.1
Thulke et al. (2024) gtr-t5-xxl 47.6 66.3 79.2 89.7
Thulke et al. (2024) gte-large 50.7 68.2 80.9 91.4
Thulke et al. (2024) ember-v1 49.5 68.6 79.7 91.1
Thulke et al. (2024) instructor-large 50.0 68.2 81.7 91.8
Thulke et al. (2024) instructor-xl 53.3 69.7 83.3 92.1
(e) Recall (R@1 and R@5) of retrieving the correct document given the question or the answer from the IFT dataset (Thulke et al., 2024).
Model CITIES STATES CORP
No Finetuning on CDP
BM25 0.055 0.084 0.153
MiniLM 0.099 0.120 0.320
Finetuned on CDP (In-Domain)
ClimateBERT 0.331 0.422 0.753
MiniLM 0.366 0.482 0.755
Best Model Finetuned on all
MiniLM 0.352 0.489 0.745
(f) MRR@10 scores on the three subsets of CLIMA-QA (Spokoyny et al., 2023). Models finetuned and evaluated on the same subset fall under In-Domain.
Task Raw Optimal Confidence
Net-zero target year 0.95 0.95 0.97
Reduction target year 0.84 0.90 0.90
Reduction base year 0.68 0.80 0.82
Reduction percentage 0.88 0.91 0.93
(g) Accuracy of the Q&A approach in assessing the quantitative properties of the net zero and reduction targets (Schimanski et al., 2023a)
Table 21. Reported performances for the datasets on Question/Answering and Retrieval tasks.
Source Model ClimaBench Pira 2.0 MCQ Exeter Misinf. Weight. Avg.
Acc. Acc. F1 (macro)
Thulke et al. (2024) Stability-3B 71.4% 48.7% 52.6% 62.8%
Thulke et al. (2024) Pythia-6.9B 63.6% 22.9% 29.9% 48.9%
Thulke et al. (2024) Falcon-7B 62.9% 19.8% 39.9% 48.3%
Thulke et al. (2024) Mistral-7B 73.1% 19.0% 63.7% 53.7%
Thulke et al. (2024) Llama-2-7B 68.5% 51.1% 59.4% 62.6%
Thulke et al. (2024) Jais-13B 66.9% 26.4% 54.2% 54.4%
Thulke et al. (2024) Jais-13B-Chat 65.8% 66.4% 61.3% 64.6%
Thulke et al. (2024) Llama-2-Chat-7B 67.8% 72.0% 64.3% 68.5%
Thulke et al. (2024) Llama-2-Chat-13B 68.6% 79.3% 68.6% 71.4%
Thulke et al. (2024) Llama-2-Chat-70B 72.7% 88.2% 72.5% 77.8%
Thulke et al. (2024) ClimateGPT-7B 75.3% 86.6% 65.9% 77.1%
Thulke et al. (2024) ClimateGPT-13B 75.0% 89.0% 70.0% 78.0%
Thulke et al. (2024) ClimateGPT-70B 72.4% 89.9% 72.5% 77.7%
Thulke et al. (2024) ClimateGPT-FSC-7B 59.3% 17.2% 45.1% 46.2%
Thulke et al. (2024) ClimateGPT-FSG-7B 53.1% 17.4% 41.5% 42.1%
(a) F1 scores (macro average) and Accuracy on multiple climate benchmarks (Thulke et al., 2024).
Source Model F1
Spokoyny et al. (2023) Majority 20.10%
Spokoyny et al. (2023) Random 24.09%
Spokoyny et al. (2023) SVM -%
Spokoyny et al. (2023) BERT 70.57%
Spokoyny et al. (2023) RoBERTa 71.14%
Spokoyny et al. (2023) DistilRoBERTa 69.27%
Spokoyny et al. (2023) Longformer 67.72%
Spokoyny et al. (2023) SciBERT 68.45%
Spokoyny et al. (2023) ClimateBERT 69.44%
(b) F1-scores (macro average) on ClimaBench (Spokoyny et al., 2023)
Source Model Acc.
Vaghefi et al. (2023) hybrid ChatClimate 4.38
Vaghefi et al. (2023) ChatClimate 4.15
Vaghefi et al. (2023) GPT-4 2.62
(c) Accuracy on a set of 10 questions, human annotations (Vaghefi et al., 2023).
Table 22. Reported performance for the datasets on Question/Answering (few-shot and RAG) Tasks.
C.8. Deceptive Technique
Source Model F1
Trained on Logic applied on LogicCLimate
Jin et al. (2022) Electra-SA 22.72%
Jin et al. (2022) Electra 27.23%
Trained on LogicClimate
Jin et al. (2022) Electra-SA 23.71%
Jin et al. (2022) Electra 29.37%
Baseline: Trained on Logic applied on Logic
Jin et al. (2022) Electra-SA 58.77%
Jin et al. (2022) Electra 53.31%
Jin et al. (2022) BERT 45.80%
(a) F1-scores (micro average) on fallacy detection (Jin et al., 2022). Electra-StructureAware is abbreviated Electra-SA
Source Model F1
Bhatia et al. (2021) BERT 59%
Bhatia et al. (2021) BERT* 62%
Bhatia et al. (2021) MTEXT 66%
Bhatia et al. (2021) MTEXT* 67%
Bhatia et al. (2021) MTEXTmulti 68%
Bhatia et al. (2021) MTEXT*multi 67%
Bhatia et al. (2021) SVM 49%
Bhatia et al. (2021) Human 70%
(b) F1-score (micro average) on neutralization techniques classification (Bhatia et al., 2021)
Source Model F1 (macro)
Coan et al. (2021) Logistic (unweighted) 68%
Coan et al. (2021) Logistic (weighted) 72%
Coan et al. (2021) SVM (unweighted) 66%
Coan et al. (2021) SVM (weighted) 72%
Coan et al. (2021) ULMFiT 72%
Coan et al. (2021) ULMFiT (weighted) 65%
Coan et al. (2021) ULMFiT (over sample) 55%
Coan et al. (2021) ULMFiT (focal Loss) 61%
Coan et al. (2021) ULMFiT-logistic 75%
Coan et al. (2021) ULMFiT-SVM 71%
Coan et al. (2021) RoBERTa 77%
Coan et al. (2021) RoBERTa-logistic 79%
Coan et al. (2021) Coder (Human) blAcc (macro): 87%
Thulke et al. (2024) Llama-2-Chat-70Bbla F1 (Bin.): 72.5%
Thulke et al. (2024) ClimateGPT-70Bblah F1 (Bin.): 72.5%
(c) F1-score (macro average and Binary) and Accuracy (macro average) on contrarian claim classification (Coan et al., 2021).
Table 23. Deceptive techniques
C.9. Environmental Performance Prediction
Source Model MSE
Lin et al. (2023b) Baseline 1.173
Lin et al. (2023b) Lasso 0.848
Lin et al. (2023b) Gradient Boosting 0.818
Lin et al. (2023b) NOCATE(RB) 0.776
Lin et al. (2023b) NOCATE(CB) 0.763
Lin et al. (2023b) NOCATE(DB) 0.756
Lin et al. (2023b) SUSTAINABLESIGNALS(RB) 0.762
Lin et al. (2023b) SUSTAINABLESIGNALS(CB) 0.753
Lin et al. (2023b) SUSTAINABLESIGNALS(DB) 0.736
(a) MSE on Finch Score prediction (Finch Score range between 0 and 10) (Lin et al., 2023b)
Source Model F1
Clarkson et al. (2020) Random Forest 89.48%
Clarkson et al. (2020) XGBoost 95.73%
(b) F1-score (macro average) on good/bad CSR performer classification (Clarkson et al., 2020).
Source Model RMSE wMAPE
Bronzini et al. (2023) OLS 7.76 7.9%
Uniform Random 40 66%
Normal Average 20 31%
(c) RMSE and wMAPE on ESG score prediction (ESG score range between 0 and 100) (Bronzini et al., 2023). We added the average RMSE and wMAPE for a uniform distribution of score with a random classifier; and for a normal distribution (mean=50, scale=20) with the average value as prediction.
Source Model MAE RMSE
Sharma et al. (2024) CLIP+MLP 0.54 0.88
Sharma et al. (2024) Sent.-BERT+MLP 0.57 0.91
Uniform Random 3.97 4.88
Gaussian Average 1.59 1.99
(d) MAE and RMSE on ESG unmanaged risk score prediction (ESG risk between 0 and 13.5) (Sharma et al., 2024). We added the average RMSE and MAE for a uniform distribution of score with a random classifier; and for a normal distribution (mean=7, scale=2) with the average value a prediction.
Source Model Accuracy (Change) Accuracy (Direction of change)
Mehra et al. (2022) majority class 0.5791 0.5682
Mehra et al. (2022) BERT 0.5985 0.4317
Mehra et al. (2022) ESGBERT 0.6709 0.793
(e) Accuracy on ESG-risk change classification (Mehra et al., 2022)
Table 24. Environmental Performance Prediction

### v2

@@
-natural language processing; greenwashing detection; corporate climate-related greenwashing; climate pretrained models; climate topic detection; green claim detection; climate QA; climate-related deceptive techniques; enviromental performance prediction copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language processing
-Abstract.
-Greenwashing is an effort to mislead the public about the carbon transition of an entity, such as a state or company.
-We provide a comprehensive survey of the scientific literature addressing natural language processing methods to identify potentially misleading climate-related corporate communications, indicative of greenwashing.
-We break the detection of greenwashing into intermediate tasks, and review the state-of-the-art approaches for each of them.
-We discuss datasets, methods, and results, as well as limitations and open challenges.
-We also provide an overview of how far the field has come as a whole, and point out future research directions.
-natural language processing; greenwashing detection; corporate climate-related greenwashing; climate pretrained models; climate topic detection; green claim detection; climate QA; climate-related deceptive techniques; enviromental performance prediction
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language processing
-1. Introduction
-Since the first report of the Intergovernmental Panel on Climate Change in 1990  (Change, 1990), world leaders and citizens have increasingly recognized the significant adverse impacts of human activities on the environment, particularly on climate change (Peter Jackson, 2007).
-This increased awareness has translated into guidelines, laws, and investments, such as the European Green Deal (Publications Office of the European Union, 2019) or the Inflation Reduction Act in the US (The White House, 2023).
-Many companies have used the financial incentives offered by states, and the guidelines and legislation to make significant steps towards sustainability (Time, 2024).
+Detecting Greenwashing: A Natural Language Processing Literature Survey
+Abstract
+Greenwashing refers to practices by corporations or governments that intentionally mislead the public about their environmental impact.
+This paper provides a comprehensive and methodologically grounded survey of natural language processing (NLP) approaches for detecting greenwashing in textual data, with a focus on corporate climate communication.
+Rather than treating greenwashing as a single, monolithic task, we examine the set of NLP problems, also known as climate NLP tasks, that researchers have used to approximate it, ranging from climate topic detection to the identification of deceptive communication patterns.
+Our focus is on the methodological foundations of these approaches: how tasks are formulated, how datasets are constructed, and how model evaluation influences reliability.
+Our review reveals a fragmented landscape: several subtasks now exhibit near-perfect performance under controlled settings, yet tasks involving ambiguity, subjectivity, or reasoning remain challenging.
+Crucially, no dataset of verified greenwashing cases currently exist.
+We argue that advancing automated greenwashing detection requires principled NLP methodologies that combine reliable data annotations with interpretable model design.
+Future work should leverage third-party judgments, such as verified media reports or regulatory records, to mitigate annotation subjectivity and legal risk, and adopt decomposed pipelines that support human oversight, traceable reasoning, and efficient model design.
+keywords:
+Greenwashing detection, Natural Language Processing, Climate NLP, Climate communication, Corporate sustainability reporting
+1 Introduction
+Since the first report of the Intergovernmental Panel on Climate Change in 1990 [change1990ipcc], world leaders and citizens have increasingly recognized the significant adverse impacts of human activities on the environment, particularly on climate change [climatebriefhistory].
+This increased awareness has translated into laws and investments, such as the European Green Deal [europeangreendeal] or the Inflation Reduction Act in the US [inflationact].
+Many companies have used the financial incentives offered by states, and the guidelines and legislation, to make significant steps towards sustainability [greencompanies].
 At the same time, this growing attention also generated an advertising opportunity for companies that aim to promote themselves as environmentally aware and responsible.
 Indeed, some companies have been found to deliberately manipulate their data and statistics to appear more environment-friendly.
-The Diesel Scandal around the Volkswagen car company is a prominent example (US Environmental Protection Agency, 2024).
+The Diesel Scandal around the Volkswagen car company is a prominent example [volkswagen].
 However, such cases are not the norm.
-More commonly, companies avoid outright data manipulation but present themselves in a misleadingly positive light regarding their environmental impact – a practice called greenwashing.
-In addition to their broader communication strategies, most companies disseminate information about their environmental initiatives through dedicated sustainability 10-K and annual reports.
-In addition, corporations also leverage various media channels and traditional communication platforms to convey their environmental credentials.
-As a result, the corpus of documents used in textual analysis is diverse and heterogeneous.
-In such documents, greenwashing can appear in multiple forms. The Green Claims Directive of the European Union (US Environmental Protection Agency, 2023) identifies four types of greenwashing:
-  1. (1)
-    using unreliable sustainability labels (e.g., H&M and Decathlon have used labels such as eco design and conscious and were prompted by the Netherlands Authority for Consumers and Markets to stop the practice in 2022 as the retailers did not clearly describe why the products received those labels (Netherlands Authority for Consumers and Markets (2022), ACM))
-  2. (2)
-    presenting legal requirements for the product as its distinctive features (e.g., McDonald’s advertises its reduction of plastic waste by using reusable cutlery (McDonald’s France, 2024), while this is a legal obligation since January 2023 in France (LegiFrance, 2020))
-  3. (3)
-    making green claims about the entire product when the claim applies only to a part/aspect of the product (e.g., Kohl’s and Walmart were fined by the U.S.
-Federal Trade Commission for the misleading claim that their products were made of bamboo, a sustainable material, when they were made out of rayon, which is derived from bamboo via a highly toxic process that removes any environmental benefits of bamboo (Lesley Fair, Federal Trade Commission, 2022)), and finally
-  4. (4)
-    making environmental claims for which the company cannot provide evidence (e.g., Keurig was fined by Canada’s Competition Bureau for misleading claims about the recycling process of its single-use coffee pods (Competition Bureau Canada, 2022)).
-Greenwashing is a widely observed practice rather than a marginal occurrence: the Australian Competition & Consumer Commission performed a sweep analysis on 247 businesses in October 2022 and found that $57\%$ of them made misleading claims about their environmental performance (Australian Competition and Consumer Commission, 2023).
-The International Consumer Protection Enforcement Network, too, found in November 2020 that $40\%$ of around 500 websites of businesses covering different sectors contain misleading green claims (UK Competition and Markets Authority, 2021).
-Addressing greenwashing is not only in the broader public interest but also in the company’s best interest, as the practice is often subject to legal sanctions prohibiting misleading or deceptive conduct, with businesses facing fines or lawsuits as a consequence (Mary Foley, Forbes, 2024).
-In addition, political awareness and public awareness of the issue via traditional and social media (Thomas Wagner, Bonpote, 2021) can harm the reputation of a company involved in the practice.
+More commonly, companies avoid outright data manipulation but present themselves in a misleadingly positive light regarding their environmental impact, a practice called greenwashing.
+Greenwashing is a widely observed practice rather than a marginal occurrence: the Australian Competition & Consumer Commission performed a sweep analysis on 247 businesses in October 2022 and found that $57\%$ of them made misleading claims about their environmental performance [sweepaustralia].
+The International Consumer Protection Enforcement Network, too, found in November 2020 that $40\%$ of around 500 websites of businesses covering different sectors contain misleading green claims [globalsweep].
+Addressing greenwashing is not only in the broader public interest but also in the company’s best interest, as the practice is often subject to legal sanctions prohibiting misleading or deceptive conduct, with businesses facing fines or lawsuits as a consequence [litigatorsgreenwashing].
+In addition, political awareness and public awareness of the issue via traditional and social media [bonpote] can harm the reputation of a company involved in the practice.
 These factors make greenwashing a risk factor that can affect investments and that should be monitored at the company and portfolio levels.
-The question thus arises to what extent greenwashing can be detected automatically.
-Such automated detection could help investors, journalists, prosecutors, and activists spot potentially illicit behavior, but it could also help companies avoid unintentionally misleading communication.
-In this paper, we provide a comprehensive survey of natural language processing (NLP) methods to identify potentially misleading climate-related corporate communication that can be indicative of greenwashing.
-We have identified 61 works that propose ways to check one or more aspects of greenwashing in text.
-We cover the intermediary steps involved in a natural language processing (NLP) analysis of the environmental disclosure of a company, ranging from detecting environmental or climate-related text to greenwashing detection.
-We structure these works in a unified categorization and discuss tasks, datasets, methods, and results for each of them.
-Unlike other surveys on the topic (Moodaley and Telukdarie, 2023b; Bernini et al., 2023), which compile and discuss findings reported in the scientific literature, our approach dives into the actual methodologies of these works.
-Our main contributions and findings are:
-  • We provide the first survey on NLP-based detection of greenwashing that describes and systemizes the methods used in the literature.
-Our survey focuses on tasks and datasets but also contextualizes the performance of the models and provides future research directions.
-  • We find no datasets of positive and negative examples of greenwashing. Hence, it is not possible to train or evaluate models on the task of identifying greenwashing.
-  • While there are no datasets explicitly labeled for greenwashing, many works have studied intermediary tasks, such as detecting misleading claims, vague language, or environmental reporting inconsistencies.
-Researchers have used these characteristics to develop theoretical definitions of greenwashing and proposed measurable indicators to identify potential cases.
-Although not a direct substitute for labeled greenwashing datasets, these indicators provide tools for analyzing corporate disclosures and assessing the likelihood of greenwashing.
-  • Open challenges remain: While many models have been proposed for greenwashing-related tasks, it is difficult to compare approaches and solutions as many studies do not report measures of uncertainty (e.g., confidence interval, multiple runs, hyper-parameters tuning), nor do they report simple baselines that help assess the difficulty of datasets such as random baseline or TF-IDF baselines.
-Even the intermediary tasks thus remain active areas of research.
-Our survey is structured as follows.
-We start with a discussion of the related work on greenwashing surveys (Section 2), followed by the legal context around climate actions and the definition of greenwashing (Section 3).
-We then survey intermediate tasks that have been studied in the literature (Section 4): identifying climate-related statements; identifying topics aligned with frameworks such as the Task Force on Climate-related Financial Disclosures (TCFD) or Environmental, Social, and Governance (ESG) factors; assessing the presence of green claims; and examining sentiment, argumentation quality, deceptive practices, and stance to discern the tone and authenticity of a message.
-We then review approaches that aim to identify greenwashing as a whole (Section 5).
-We conclude by discussing challenges and opportunities in greenwashing detection (Sections 6 and 7).
-We note that we have an Appendix section that will be part of the supplemental online-only material.
-2. Related Work
-Moodaley and Telukdarie (2023b) analyze the literature on greenwashing, sustainability reporting, and the intersection with research in Artificial Intelligence (AI) and Machine Learning (ML).
+The growing public and regulatory attention to corporate climate claims raises a central question: to what extent can greenwashing be detected automatically?
+Automated detection systems could support investors, journalists, prosecutors, and activists in identifying potentially misleading or illicit communication, while also helping companies avoid unintentionally deceptive disclosures.
+Yet despite the relevance of this question, the methodological foundations of NLP-based greenwashing detection remain fragmented.
+Methodological grounding
+In this paper, we provide the first comprehensive and methodologically grounded survey of NLP techniques for identifying potentially misleading climate-related corporate communication.
+We analyze 61 scientific works that examine one or more aspects of greenwashing, ranging from identifying climate-related statements and assessing environmental claims to evaluating specificity, stance, argumentation, and the presence of deceptive linguistic cues.
+Unlike prior surveys [moodaley_greenwashing_2023, measuring_greenwashing], which primarily summarize findings reported in the literature, our focus is on the methodological foundations of these approaches: how tasks are formulated, how datasets are constructed, and how model evaluation influences reliability.
+Summary of findings
+Recent advances in NLP have enabled increasingly sophisticated analyses of climate-related corporate communication, yet our survey shows that progress across the building blocks of greenwashing detection is highly uneven.
+Several tasks, such as climate topic identification, climate risk classification, and the characterization of environmental claim types, now achieve near perfect performance when tackled with fine-tuned models.
+Even simple keyword-based approaches often achieve surprisingly strong results, suggesting that many of these tasks are driven by distinctive lexical terms rather than deep domain knowledge.
+In contrast, tasks that require nuance, subjectivity, or reasoning, such as distinguishing specific from vague commitments or assessing rhetorical deception, remain challenging.
+Low inter-annotator agreement, ambiguous label definitions, and small datasets, undermine the reliability of the reported performance.
+This pattern reflects a broader issue: many real indicators of greenwashing require judgment on subtle differences, multi-document context, or implicit reasoning, where current models struggle and where evaluation is especially difficult.
+These challenges are rendered more difficult by methodological weaknesses in research papers that limit the interpretability and reproducibility of results.
+Evaluation practices are often inadequate: few studies report uncertainty estimates, many rely on accuracy without considering semantic proximity between labels, and baselines are frequently missing.
+Only a minority compare to random, majority, or keyword-based baselines, even though these can drastically change how model performance is interpreted.
+Furthermore, many experiments are conducted on curated, balanced, and noise-free datasets, conditions that that might not hold in real-world datasets.
+A more fundamental limitation is the absence of datasets containing verified instances of greenwashing.
+While the literature proposes theoretical indicators such as overly positive sentiment, rhetorical cues, and stance inconsistencies, none are validated against real cases.
+As a result, most studies evaluate proxy tasks rather than greenwashing, leaving the field without empirical foundations for end-to-end detection.
+We believe that creating a greenwashing dataset requires principled data collection grounded in third-party judgments (e.g., regulators, courts, investigative journalism), clearer and narrower definitions of misleading communication, and decomposed pipelines that combine automated analysis with human oversight.
+In conclusion, these findings show that while some subtasks appear solved, building a reliable and trustworthy NLP system for detecting greenwashing remains an open challenge.
+Structure of the survey
+The remainder of the paper proceeds as follows.
+We begin by situating our work within the broader literature on greenwashing and automated climate communication analysis (Section 2).
+We then outline the regulatory landscape that shapes how greenwashing is defined (Section 3), as these definitions constrain what can be annotated or modeled.
+Section 4 introduces the building-block tasks that underpin automated greenwashing detection: identifying climate-related statements; identifying topics aligned with frameworks such as the Task Force on Climate-related Financial Disclosures (TCFD) or Environmental, Social, and Governance (ESG) factors; assessing the presence of green claims; and examining sentiment, argumentation quality, deceptive practices, and stance to discern the tone and authenticity of a message.
+For each task, we give the task description, annotated or annotation-free datasets, models and an analysis of their performance, and finally insights, in which we highlight important limitations and potential future direction.
+Section 5 reviews approaches that aim to identify greenwashing as a whole.
+Finally, Sections 6 and 7 discuss open methodological challenges and outline research opportunities for building principled, reliable systems for greenwashing detecting.
+2 Related Work
+moodaley_greenwashing_2023 analyze the literature on greenwashing, sustainability reporting, and the intersection with research in Artificial Intelligence (AI) and Machine Learning (ML).
 The authors provide an overview of how the literature has changed over time by examining the number of publications and the different keywords covered by such publications.
 Unlike our work, this survey does not describe the actual approaches of the publications.
-Bernini et al.
-(2023) review the definitions of greenwashing in business, management, and accounting.
+measuring_greenwashing review the definitions of greenwashing in business, management, and accounting.
 They conclude that there is no agreed-upon definition of greenwashing, except that it is defined as a gap between the company’s actions and its disclosure.
 In contrast to our study, the review provides only a brief overview of the techniques for detecting greenwashing.
-While our review focuses on the semantic and linguistic aspects of greenwashing identification, other works focus on quantitative approaches that can serve the same objective (Dao et al., 2024).
+While our review focuses on the semantic and linguistic aspects of greenwashing identification, other works focus on quantitative approaches that can serve the same objective dao2024introduction.
+LUBLOY2025123399 studied the quantification of firm-level greenwashing; however, our work is focused on natural language processing approaches, while theirs is more generic.
 Other works are less directly related to our survey.
 For example, a growing community is focused on using AI to tackle climate change [^1] See for example https://www.climatechange.ai/.
-Rolnick et al.
-(2022) review how machine learning has been used and could be used to tackle climate change.
+Rolnick_tackling_climate review how machine learning has been used and could be used to tackle climate change.
 They explore different sectors of the economy, such as electricity, transportation, industry, and agriculture, among others.
 The survey discusses climate investments and the effects of climate change on finance, but not greenwashing.
-Alonso et al.
-(2023) provide a survey on ML approaches for climate finance, also known in the literature as green finance or carbon finance.
+alonsoMachineLearningMethods2023 provide a survey on ML approaches for climate finance, also known in the literature as green finance or carbon finance.
 Climate finance is a subfield in economics dedicated to providing tools from financial economics to tackle climate change.
-Finally, several works study the impact of AI on climate change (Hershcovich et al., 2022; Kaack et al., 2022; Rohde et al., 2023; Verdecchia et al., 2023).
+Finally, several works study the impact of AI on climate change [hershcovich_towards_2022, kaack_aligning_2022, rohde_broadening_2023, verdecchia_systematic_2023].
 Our work, in contrast, surveys approaches that use AI to detect greenwashing.
-3. Preliminaries
-3.1. Legal Context
+3 Preliminaries
+3.1 Legal Context
 In this section, we provide the legal context for the definition of greenwashing. We first present international and national laws, followed by how they are applied in the corporate context.
 International and Nation Level Legislation and Guidelines on Climate Change Mitigation
-The first international summit on the effect of humans on the environment was the United Nations Conference on the Human Environment, held in 1972 in Sweden (United Nations, European Commission, 2024).
-The Intergovernmental Panel on Climate Change (IPCC) was created in 1988 and released its first report in 1990  (Change, 1990).
-It was only in 1992 that the first treaty between countries was signed, the United Nations Framework Convention on Climate Change (UNFCCC) (United Nations, European Commission, 2024).
-The treaty only encouraged countries to reduce their emissions, and in 1997, the Kyoto Protocol set commitments for the countries to follow (United Nations Framework Convention on Climate Change, 2011).
+The first international summit on the effect of humans on the environment was the United Nations Conference on the Human Environment, held in 1972 in Sweden [timelineclimate].
+The Intergovernmental Panel on Climate Change (IPCC) was created in 1988 and released its first report in 1990 [change1990ipcc].
+It was only in 1992 that the first treaty between countries was signed, the United Nations Framework Convention on Climate Change (UNFCCC) [timelineclimate].
+The treaty only encouraged countries to reduce their emissions, and in 1997, the Kyoto Protocol set commitments for the countries to follow [kyotoprotocol].
 The 36 countries that participated in the Protocol reduced their emissions in 2008–2012 by a large margin in respect to the levels in 1990; however, the global emissions increased by $32\%$ in the same period.
@@
 The trading system limits the amount of CO2 emitted by European industries and covers $46\%$ of the EU’s CO2 emissions.
-It is estimated that the trading system reduced CO2 emissions in the EU by $3.8\%$ between 2008 and 2016 (Bayer and Aklin, 2020).
+It is estimated that the trading system reduced CO2 emissions in the EU by $3.8\%$ between 2008 and 2016 [bayer2020european].
 At COP21, in 2015, 194 nations plus the European Union signed the Paris Agreement, a treaty by which the nations commit to keep the rise in global surface temperature below 2 °C (3.6 °F) above pre-industrial levels.
@@
 The first such assessment took place at COP28, where it was established that an important direction to achieve the long-term goals of countries was to transition away from fossil fuels to renewable energy.
-In 2021, the European Climate Law was adopted: the EU commits to reducing its emissions by at least $55\%$ by 2030 with respect to the 1990 levels and becoming climate neutral by 2050 (Council of the EU Press release, 2021).
-Similarly, countries such as Canada (Government of Canada, 2024), Taiwan (The Republic of China Ministry of Environment(2023), Taiwan), South Korea (The Korean Law Information Center, 2024), and Australia (Government of New South Wales, 2023), among others, aim to achieve carbon neutrality by 2050 and have passed laws setting this goal.
+In 2021, the European Climate Law was adopted: the EU commits to reducing its emissions by at least $55\%$ by 2030 with respect to the 1990 levels and becoming climate neutral by 2050 [euclimatelaw].
+Similarly, countries such as Canada [canadaclimatelaw], Taiwan [taiwanclimatelaw], South Korea [koreaclimatelaw], and Australia [australiaclimatelaw], among others, aim to achieve carbon neutrality by 2050 and have passed laws setting this goal.
 Corporate Level Laws and Guidelines
@@
 Even without laws, investors can be concerned about environmental or social issues; hence, companies have used voluntary disclosures for a long time.
-One of the first organizations to provide standards for reporting on climate change or social aspects is the Global Reporting Initiative (GRI), with its first guidelines published in 2000 (GRI, 2023).
-In 2015, the Financial Stability Board and the Group of 20 created the Task Force for Climate-related Financial Disclosure (TCFD) guidelines on disclosure in response to shortcomings of COP21, in particular the lack of standards climate-related disclosure (Wikipedia contributors, 2024).
-In 2021, the International Sustainability Standards Board (ISSB) was created to establish standards for climate-related disclosure, and starting in 2024, the standard released by this board will be applied worldwide (IFRS Foundation, 2024).
-ISSB standards were aligned with GRI disclosure standards to make them complementary and interoperable (Global Reporting Initiative, 2023), while the ISSB standards are expected to take over TCFD (IFRS Foundation, 2024).
-Differently from the voluntary reporting standards of GRI, TCFD, and ISSB, the Corporate sustainability reporting law in the EU required the creation of the European Sustainability Reporting Standards (ESRS), which are mandatory for companies subject to EU law (Directorate-General for Financial Stability, Financial Services and Capital Markets Union, 2023).
+One of the first organizations to provide standards for reporting on climate change or social aspects is the Global Reporting Initiative (GRI), with its first guidelines published in 2000 [grihistory].
+In 2015, the Financial Stability Board and the Group of 20 created the Task Force for Climate-related Financial Disclosure (TCFD) guidelines on disclosure in response to shortcomings of COP21, in particular the lack of standards climate-related disclosure [enwiki:1257716178].
+In 2021, the International Sustainability Standards Board (ISSB) was created to establish standards for climate-related disclosure, and starting in 2024, the standard released by this board will be applied worldwide [ifrs].
+ISSB standards were aligned with GRI disclosure standards to make them complementary and interoperable [gri], while the ISSB standards are expected to take over TCFD [ifrs].
+Differently from the voluntary reporting standards of GRI, TCFD, and ISSB, the Corporate sustainability reporting law in the EU required the creation of the European Sustainability Reporting Standards (ESRS), which are mandatory for companies subject to EU law [esrs].
 ESRS has high interoperability with GRI and ISSB.
@@
 Some European countries anticipated this legislation with their own, for example, the 2001 New Economic Regulations Act in France.
-Switzerland, which is not part of the EU, has imposed a mandatory TCFD disclosure for large public companies, banks, and insurance companies starting from January 2024 (The Swiss Federal Council, 2022), and similar laws exist in New Zealand (New Zealand Ministry of Environment, 2023).
+Switzerland, which is not part of the EU, has imposed a mandatory TCFD disclosure for large public companies, banks, and insurance companies starting from January 2024 [disclosureswiss], and similar laws exist in New Zealand [disclosurezealand].
 While not compulsory, initiatives like the Carbon Disclosure Project (CDP) have played a pivotal role by standardizing responses related to the climate disclosure of a company via structured questionnaires, thereby facilitating more systematic and comparable reporting.
-3.2. Definition of Greenwashing
-A widely cited and comprehensive definition, synthesizing those commonly found in the literature, is provided by the Oxford English Dictionary (OED, 2023).
+3.2 Definition of Greenwashing
+A widely cited and comprehensive definition, synthesizing those commonly found in the literature, is provided by the Oxford English Dictionary [GreenwashMeaningsEtymology2023].
 Definition 1.
@@
 As climate-related disclosures face increasing regulation and greenwashing poses significant risks to a company’s reputation, some companies adopt a strategy of silence, avoiding discussions about their environmental impact.
-This deliberate lack of communication is known as greenhushing (Letzing, 2022).
+This deliberate lack of communication is known as greenhushing [Letzing].
 These definitions imply that greenwashing is a deliberate act; however, in many cases, it results from an error or miscommunication by companies genuinely trying to showcase their sustainability efforts, and that are trying to best follow disclosure standards.
-As shown by Definitions 3.1 and 3.2, greenwashing is not defined by easily identifiable properties but as a general concept.
+As shown by Definitions 1 and 2, greenwashing is not defined by easily identifiable properties but as a general concept.
 Since the concept is so unspecific, researchers focused on components indicative of potentially misleading communications but easier to define.
@@
 They should all be understood in the context of climate-related misleading communications, as components associated or indicative of potential greenwashing even if they are not synonymous.
-4. Intermediary tasks
-4.1. Pretraining Models on Climate-Related Text
+4 Foundational Tasks for Greenwashing Detection
+4.1 Pretraining Models on Climate-Related Text
 The first step in applying a language model for a given task is typically the pretraining process, which involves training the model on relevant corpora.
-Although general-purpose models such as BERT (Devlin et al., 2019) and LLaMA (Touvron et al., 2023) have demonstrated strong performance in various tasks, they are inherently limited by the knowledge and vocabulary present in their training corpora.
-This limitation can lead to suboptimal results when these models are applied to highly specialized domains with unique terminology and concepts.
-Hence, domain-specific models are usually trained on targeted datasets to improve their performance in niche areas.
+Although general-purpose models such as BERT [devlin-etal-2019-bert] and LLaMA [touvron2023llamaopenefficientfoundation] have demonstrated strong performance in various tasks, they are inherently limited by the knowledge and vocabulary present in their training corpora.
+This has motivated the development of climate-focused language models, which aim to encode domain knowledge directly into the pretrained representations.
 Task Description
-A language model is pretrained on climate-related text, usually on the classical tasks of masked token prediction and next sentence prediction.
-The goal is to produce a domain-specific pre-trained language model that will be subsequently fine-tuned to specific tasks.
-There are two approaches: (1) training a model with a domain-specific corpus from scratch, or (2) further-training [^2] Further-training refers to the process of training a pretrained model on its original task (e.g., next-token prediction or masked-token prediction) using additional domain-specific data to specialize its knowledge for that domain.
+A language model is pretrained on climate-related text, to produce a domain-specific language model that will be subsequently fine-tuned to specific tasks. There are two approaches:
+  1. (i)
+    Training a model with a domain-specific corpus from scratch, or
+  2. (ii)
+    Further-training [^2] Further-training refers to the process of training a pretrained model on its original task using additional domain-specific data to specialize its knowledge for that domain.
 a generalist pretrained model on a domain-specific corpus.
 Annotated datasets
-There have been several efforts to create a climate-related corpus.
-Webersinke et al.
-(2022) gathered climate-related news articles, research abstracts, and corporate reports.
-Vaghefi et al.
-(2022) introduced Deep Climate Change, a dataset composed of abstracts of articles from climate scientists, and built a corpus specific to climate research texts.
-Schimanski et al.
-(2023b) introduced a dataset that focuses on text related to Environment Social and Governance (ESG).
-Similarly, Mehra et al.
-(2022) built a dataset using text from the Knowledge Hub of Accounting for Sustainability for an ESG domain-specific corpus.
-More recently, Thulke et al.
-(2024) proposed a dataset comprising news, publications (abstracts and articles), books, patents, the English Wikipedia, policy and finance-related texts, Environmental Protection Agency documents, and ESG, and IPCC reports.
-Mullappilly et al.
-(2023) proposed a climate-specific multilingual dataset.
-Yu et al.
-(2024) published a pretraining dataset constructed with annual and sustainable reports from EU banks.
-Unfortunately, to the best of our knowledge, only Yu et al.
-(2024)’s ClimateBUG-data dataset is publicly available.
-Supervised Solutions
-Based on the above datasets, the models ClimateBERT, ClimateGPT-2, climateBUG-LM and EnvironmentalBERT, ESGBERT were proposed (Webersinke et al., 2022; Vaghefi et al., 2022; Yu et al., 2024; Schimanski et al., 2023b; Mehra et al., 2022).
-More recently, generative domain-specific models such as Llama-2 for ClimateGPT (Thulke et al., 2024) or Vicuna for Arabic Mini-ClimateGPT (Mullappilly et al., 2023) have also been proposed.
+A variety of climate-related corpora have been assembled for this purpose, ranging from scientific abstracts and policy documents to sustainability reports and ESG disclosures.
+These corpora differ in scope and quality, and only a few are publicly accessible.
+The lack of standardized, openly available pretraining datasets remains a barrier to consistent benchmarking and reproducibility.
+nicolas_webersinke_climatebert_2021 gathered climate-related news articles, research abstracts, and corporate reports.
+vaghefi2022deep introduced Deep Climate Change, a dataset composed of abstracts of articles from climate scientists, and built a corpus specific to climate research texts.
+schimanski_bridging_2023 introduced a dataset that focuses on text related to Environment Social and Governance (ESG).
+Similarly, Mehra_2022 built a dataset using text from the Knowledge Hub of Accounting for Sustainability for an ESG domain-specific corpus.
+More recently, thulke2024climategpt proposed a dataset comprising news, publications (abstracts and articles), books, patents, the English Wikipedia, policy and finance-related texts, Environmental Protection Agency documents, and ESG, and IPCC reports.
+mullappilly-etal-2023-arabic proposed a climate-specific multilingual dataset.
+yu_climatebug_2024 published a pretraining dataset constructed with annual and sustainable reports from EU banks.
+Unfortunately, to the best of our knowledge, only [yu_climatebug_2024]’s ClimateBUG-data dataset is publicly available.
+Models
+Based on the above datasets, the models ClimateBERT, ClimateGPT-2, climateBUG-LM and EnvironmentalBERT, ESGBERT were proposed [nicolas_webersinke_climatebert_2021, vaghefi2022deep, yu_climatebug_2024, schimanski_bridging_2023, Mehra_2022].
+More recently, generative domain-specific models such as Llama-2 for ClimateGPT [thulke2024climategpt] or Vicuna for Arabic Mini-ClimateGPT [mullappilly-etal-2023-arabic] have also been proposed.
 Performance of Models
-The domain-specific models drastically improve the performance on domain-specific masked-language modeling (Webersinke et al., 2022; Yu et al., 2024) and next token prediction (Vaghefi et al., 2022).
+The domain-specific models drastically improve the performance on domain-specific masked-language modeling [nicolas_webersinke_climatebert_2021, yu_climatebug_2024] and next token prediction [vaghefi2022deep].
 They were also evaluated on domain-specific downstream tasks.
-These tasks are either based on pre-existing datasets such as ClimateFEVER (Leippold and Diggelmann, 2020) or introduced by the authors, as in ClimateBERT’s climate detection (Webersinke et al., 2022).
+These tasks are either based on pre-existing datasets such as ClimateFEVER [diggelmann_climate-fever_2020] or introduced by the authors, as in ClimateBERT’s climate detection [nicolas_webersinke_climatebert_2021].
 At this stage, we evaluate whether fine-tuning enhances downstream performance, with a detailed analysis of the tasks presented in subsequent sections.
-Cl.BERT DisRoBERTa NB
-Climatext (Spokoyny et al., 2023; Leippold and Varini, 2020) 85.14 86.06 83.39
-Climate Detection (Webersinke et al., 2022) 99.1±$\pm 1$ 98.6±$\pm 0.8$
-Sentiment (Webersinke et al., 2022) 83.8±$\pm 3.6$ 82.5±$\pm 4.6$
-Net-zero/Reduction (Schimanski et al., 2023a) 96.2±$\pm 0.4$ 94.4±$\pm 0.6$
-Cl.BERT DisRoBERTa TF-IDF
-TCFD classification (Sampson et al., 2022) 0.852 0.819 0.867
-Cl.BERT DisRoBERTa SVM
-SciDCC (Spokoyny et al., 2023; Mishra and Mittal, 2021) 52.97 51.13 48.02
-ClimaTOPIC (Spokoyny et al., 2023) 64.24 63.61 58.34
-Cl.FEVER (claim) (Xiang and Fujii, 2023; Leippold and Diggelmann, 2020) 76.8 72
-Cl.FEVER (claim) (Webersinke et al., 2022; Leippold and Diggelmann, 2020) 75.7±$\pm 4.4$ 74.8±$\pm 3.6$
-Cl.FEVER (evid.) (Spokoyny et al., 2023; Leippold and Diggelmann, 2020) 61.54 61.54
-Cl.Stance (Spokoyny et al., 2023; Vaid et al., 2022) 52.84 52.51 42.92
-ClimateEng (Spokoyny et al., 2023; Vaid et al., 2022) 71.83 72.33 51.81
-ClimaINS (Spokoyny et al., 2023) 84.80 84.38 86.00
-ClimaBENCH (Spokoyny et al., 2023) 69.44 69.27
-Nature (Schimanski et al., 2024a) 94.11 94.03
-Cl.BERT SVM+ELMo SVM+BoW
-Commitment&Actions (Bingler et al., 2023) 81 79 76
-Specificity (Bingler et al., 2023) 77 76 75
-Cl.BERT DisBERT SVM
-Env. Claim (Stammbach et al., 2023) 83.8 83.7 70.9
-Cl.BERT DisBERT LSTM
-DARE sentiment (Xiang and Fujii, 2023) 87.4 89.9 88.2
-(a) F1-scores of climateBERT when compared to a similarly sized models. ROC-AUC for TCFD classification.
-RoBERTa EnvRoBERTa DisRoBERTa EnvDisRoB.
-Environment (Schimanski et al., 2023b) 92.35±$\pm 2.29$ 93.19±$\pm 1.65$ 90.97±$\pm 2$ $92.35\pm 1.65$
-Social (Schimanski et al., 2023b) 89.87±$\pm 1.35$ 91.90±$\pm 1.79$ 90.59±$\pm 1.03$ $91.24\pm 1.86$
-Governance (Schimanski et al., 2023b) 77.03±$\pm 1.82$ 78.48±$\pm 2.62$ 76.65±$\pm 2.39$ 78.86±$\pm 1.59$
-EnvDisRoB. ClimateBERT RoBERTa DisRoBERTa
-Water (Schimanski et al., 2024a) 94.47±$\pm 1.37$ 95.10±$\pm 1.13$ 94.55±$\pm 0.86$ $94.98\pm 1.16$
-Forest (Schimanski et al., 2024a) 95.37±$\pm 0.92$ 95.34±$\pm 0.94$ 94.78±$\pm 0.48$ 95.29±$\pm 0.65$
-Biodiversity (Schimanski et al., 2024a) 92.76±$\pm 1.01$ 92.49±$\pm 1.03$ 92.46±$\pm 1.54$ 92.29±$\pm 1.23$
-Nature (Schimanski et al., 2024a) 94.19±$\pm 0.81$ 93.50±$\pm 0.64$ 93.97±$\pm 0.26$ 93.55±$\pm 0.72$
-(b) F1-scores of multiple models compared to EnvironmentalBERT (based on RoBERTa and distilRoBERTa).
-GPT-2 climateGPT-2
-ClimateFEVER (Vaghefi et al., 2022) 62 72
-LLama-2-Chat (7B) ClimateGPT (7B)
-Multiple Benchmarks (Thulke et al., 2024) 71.4 77.1
-(c) F1-scores of ClimateGPT-2 and custom performances (average Accuracy and F1-score over multiple benchmarks) of ClimateGPT compared to similarly sized models.
-Table 1.
-Performance of domain-specific models on domain-specific tasks.
-The figure displayed in the tables are the values reported by authors in the corresponding studies; with the following abbreviations: CL for Climate, Dis for Distil, EnvDisRob for EnvDistilRoBERTa, Evid.
+Table 1: Performance of ClimateBERT on domain-specific tasks (using F1-score, except for TCFD classification which is using ROC-AUC).
+The figure displayed in the tables are the values reported by authors in the corresponding studies; with the following abbreviations: CL for Climate, Dis for Distil, Evid.
 for evidences.
@@
 Detailed performances are reported in Appendix C.
-As shown in Table 1(c), both ClimateGPT-2 and ClimateGPT significantly outperform the baseline model in their designated tasks.
+Cl. BERT Dis. RoBERTa NB
+Climatext [spokoyny2023answering, varini_climatext_2020] 85.14 86.06 83.39
+[…]
+[…]
+[…]
+Cl. BERT Dis. RoBERTa TF-IDF
+TCFD classification [sampson_tcfd-nlp_nodate] 0.852 0.819 0.867
+Cl. BERT Dis. RoBERTa SVM
+SciDCC [spokoyny2023answering, mishra2021neuralnere] 52.97 51.13 48.02
+ClimaTOPIC [spokoyny2023answering] 64.24 63.61 58.34
+Cl.FEVER (claim) [xiang_dare_2023, diggelmann_climate-fever_2020] 76.8 72
+[…]
+Cl.FEVER (evid.) [spokoyny2023answering, diggelmann_climate-fever_2020] 61.54 61.54
+Cl.Stance [spokoyny2023answering, vaid-etal-2022-towards] 52.84 52.51 42.92
+ClimateEng [spokoyny2023answering, vaid-etal-2022-towards] 71.83 72.33 51.81
+ClimaINS [spokoyny2023answering] 84.80 84.38 86.00
+ClimaBENCH [spokoyny2023answering] 69.44 69.27
+Nature [Schimanski2024nature] 94.11 94.03
+Cl. BERT SVM+ ELMo SVM+ BoW
+Commitment&Actions [bingler2023cheaptalkspecificitysentiment] 81 79 76
+Specificity [bingler2023cheaptalkspecificitysentiment] 77 76 75
+Cl. BERT Dis. BERT SVM
+Env. Claim [stammbach_environmental_2023] 83.8 83.7 70.9
+Cl. BERT Dis. BERT LSTM
+DARE sentiment [xiang_dare_2023] 87.4 89.9 88.2
+Table 2: F1-scores of multiple models compared to EnvironmentalBERT.
+The figure displayed in the tables are the values reported by authors in the corresponding studies; with the following abbreviations: Dis for Distil, EnvDisRob for EnvDistilRoBERTa.
+Each row reports the performances of models fine-tuned in the same experimental setting.
+Detailed performances are reported in Appendix C.
+RoBERTa EnvRoBERTa DisRoBERTa EnvDisRoB.
+[…]
+[…]
+[…]
+EnvDisRoB. ClimateBERT RoBERTa DisRoBERTa
+[…]
+[…]
+[…]
+[…]
+Table 3: F1-scores of ClimateGPT-2 and custom performances of ClimateGPT compared to similarly sized models on domain-specific benchmarks.
+The figure displayed in the tables are the values reported by authors in the corresponding studies.
+Detailed performances are reported in Appendix C.
+GPT-2 climateGPT-2
+ClimateFEVER [vaghefi2022deep] 62 72
+LLama-2-Chat (7B) ClimateGPT (7B)
+Multiple Benchmarks [thulke2024climategpt] 71.4 77.1
+As shown in Table 1 and Table 2, evaluations on climate-related downstream tasks show mixed but informative results: (i) domain-adapted models often outperform their base counterparts on specialized tasks, (ii) gains are frequently modest, and (iii) strong generalist baselines—when fine-tuned on the same domain-specific datasets—remain highly competitive.
+These findings indicate that domain adaptation helps encode environmental terminology and discourse structure, but its benefits are task-dependent and not uniformly substantial.
+In Table 3, both ClimateGPT-2 and ClimateGPT significantly outperform the baseline model.
 These models benefit from additional training on climate-related data, highlighting the advantages of domain adaptation.
-However, when trained from scratch, Thulke et al.
-(2024) report substantially lower performance, underscoring the importance of large-scale pretraining data.
-For smaller models (climateBERT, climateBUG-LM and EnvironmentalBERT), the domain-adaptation also enhanced the performance on domain-specific tasks, though the significance of this improvement is less pronounced.
-As detailed in Table 1(a), out of the 15 evaluations of fine-tuned domain-specific climateBERT (Webersinke et al., 2022) on climate-related tasks compared to the fine-tuned base model (distilRoBERTa), 4 reported an improvement compared to the base model (¿1%), 8 reported a marginal improvement (¡1%) and 3 indicate no measurable difference (¡=0%).
-When compared to DistilBERT in the last two tasks from Table 1(a), climateBERT exhibits a slight improvement in one case, while performing comparably in the other.
-Additionally, climateBERT achieves performance similar to the simple baselines in certain instances (Table 1(a)).
-For example, in Specificity Classification (Bingler et al., 2023), climateBERT outperforms an SVM with bag-of-word by 2%, and on the TCFD classification (Sampson et al., 2022) climateBERT performs slightly below the TF-IDF baseline model.
-Likewise, climateBUG-LM (Yu et al., 2024) demonstrates a slight improvement of less than 1% over BERT and FinBERT on a climate-related detection task.
-As shown in Table 1(c), the domain-adapted models shows a small improvement in the performance on the Environmental task, by 1.38% for distilRoBERTa and 0.84% for RoBERTa.
+However, when trained from scratch, thulke2024climategpt report substantially lower performance, underscoring the importance of large-scale pretraining data.
 Insights
-Domain-specific language models, particularly those developed for environmental and climate-related texts, offer an interesting avenue for improving model performance through enhanced knowledge within specialized domains.
-Theoretically, such models should outperform general models by leveraging their tailored vocabulary and contextual understanding.
-However, our findings suggest that the improvements provided by the domain-specific models on the proposed domain-specific downstream tasks are limited.
+In theory, domain-specific language models should outperform general models by leveraging their tailored vocabulary and contextual understanding.
+However, our findings suggest that the improvements provided by these models on the proposed domain-specific downstream tasks are limited.
 Indeed, we found that fine-tuning general models on domain-specific datasets consistently yields competitive performance.
-However, these findings do not undermine the significance of existing works that have been fundamental in demonstrating the potential of domain-specific models, particularly in generating domain-specific text (MLM and next token prediction).
-Moreover, ClimateGPT (Thulke et al., 2024) demonstrated significant improvement for the domain-adapted Llama-2, showing that domain-adaptation is highly relevant.
+These findings do not undermine the significance of existing works that have been fundamental in demonstrating the potential of domain-specific models, particularly in generating domain-specific text.
+Moreover, ClimateGPT [thulke2024climategpt] demonstrated significant improvement for the domain-adapted Llama-2, showing that domain-adaptation is highly relevant.
 Future research could focus on designing tasks where a lack of domain understanding—whether in terms of vocabulary or knowledge—significantly hinders performance.
-4.2. Climate-Related Topic Detection
+4.2 Climate-Related Topic Detection
 The first step toward detecting climate-related greenwashing is identifying text addressing climate-related topics.
@@
 Given an input sentence or a paragraph, output a binary label, “climate-related” or “not climate-related”.
+Table 4: Label definitions of the datasets related to climate change and sustainability topic detection.
 Dataset Input Labels Positive Label Details
-ClimateBug-data (Yu et al., 2024) sentences from Banks’ reports relevant/ irrelevant Climate change and sustainability (including ESG, SDGs related to the environment, recycling and more)
-ClimateBERT’s climate detection (Bingler et al., 2023) paragraphs from reports 1/0 Climate policy, climate change or an environmental topic
-Climatext (Wikipedia, 10-K, claims) (Leippold and Varini, 2020) sentences from Wikipedia, 10-Ks or web scraping 1/0 Directly related to climate-change
-Climatext (wiki-doc) (Leippold and Varini, 2020) sentences from a Wikipedia page 1/0 Extracted from a Wikipedia page related to climate-change
-Sustainable signals’s reviews (Lin et al., 2023b) online product reviews (user comments) relevant/ irrelevant Contains terms related to sustainability
-Table 2. Label definitions of the datasets related to climate change and sustainability topic detection.
-Unsupervised Solutions
-The earliest work on identifying corporate climate-related text was conducted by Doran and Quinn (2009), where the authors collected a large corpus of 10-K filings from 1995 to 2008, and filtered climate-related reports using hand-selected keywords.
-While Doran and Quinn (2009) focused on identifying climate-related text within company reports to assess how businesses addressed climate change, other early efforts took a different approach by analyzing the external impacts of media attention toward climate change on companies.
-For instance, Engle et al.
-(2020) developed the WSJ Climate Change News Index based on identifying climate-related news articles in media sources using climate-change vocabulary frequency.
+ClimateBug-data yu_climatebug_2024 Sentences from bank reports relevant/ irrelevant Climate change and sustainability (including ESG, SDGs related to the environment, recycling and more)
+ClimateBERT’s climate detection bingler2023cheaptalkspecificitysentiment Paragraphs from reports 1/0 Climate policy, climate change or an environmental topic
+Climatext (Wikipedia, 10-K, claims) varini_climatext_2020 Sentences from Wikipedia, 10-Ks or web scraping 1/0 Directly related to climate-change
+Climatext (wiki-doc) varini_climatext_2020 Sentences from a Wikipedia page 1/0 Extracted from a Wikipedia page related to climate-change
+Sustainable signals’s reviews linSUSTAINABLESIGNALSijcai2023 Online product reviews (user comments) relevant/ irrelevant Contains terms related to sustainability
+Annotation-free
+The earliest work on identifying corporate climate-related text was conducted by doran_risk_disclosure, where the authors collected a large corpus of 10-K filings from 1995 to 2008, and filtered climate-related reports using hand-selected keywords.
+While doran_risk_disclosure focused on identifying climate-related text within company reports to assess how businesses addressed climate change, other early efforts took a different approach by analyzing the external impacts of media attention toward climate change on companies.
+For instance, Engle_hedging_climatechange developed the WSJ Climate Change News Index based on identifying climate-related news articles in media sources using climate-change vocabulary frequency.
 This index served as a tool to measure the influence of climate-change attention in the media on companies.
-Gorovaia and Makrominas (2025) proposed measuring the environmental content of CSR (corporate social responsibility) reports using a lexicon-based approach, with the goal of comparing CSR reports of environmental violators to companies with clean records.
+csr_report_greenwashing proposed measuring the environmental content of CSR (corporate social responsibility) reports using a lexicon-based approach, with the goal of comparing CSR reports of environmental violators to companies with clean records.
 Annotated datasets
 More recent studies have introduced several annotated datasets aimed at climate-related corporate text classification.
-One of the most prominent is ClimaText (Leippold and Varini, 2020), a large dataset from diverse sources such as Wikipedia, U.S.
+One of the most prominent is ClimaText [varini_climatext_2020], a large dataset from diverse sources such as Wikipedia, U.S.
 Security and Exchange Commission 10-K filings, and web-scraped content.
@@
 Climatext (Wiki-doc) is annotated automatically, while Climatext (Wikipedia, 10-k and claims) is annotated by humans.
-In contrast, ClimateBERT’s authors (Webersinke et al., 2022) provide a smaller, more focused dataset consisting of paragraphs from companies’ annual and sustainability reports.
+In contrast, ClimateBERT’s authors [nicolas_webersinke_climatebert_2021] provide a smaller, more focused dataset consisting of paragraphs from companies’ annual and sustainability reports.
 This dataset was created as a downstream task for evaluating the ClimateBERT model, focusing on corporate disclosures.
-The dataset was later extended by Bingler et al.
-(2023), increasing its size and refining the annotation to capture both specificity and sentiment related to corporate climate discourse (see Sections 4.44.6).
-Another large dataset, ClimateBUG-data (Yu et al., 2024), also focuses on corporate communications but in EU banks.
-Beyond corporate disclosures, Sustainable Signals (Lin et al., 2023b) introduces a dataset of product reviews, each classified based on its relevance to sustainability.
+The dataset was later extended by bingler2023cheaptalkspecificitysentiment, increasing its size and refining the annotation to capture both specificity and sentiment related to corporate climate discourse (see Sections 4.44.6).
+Another large dataset, ClimateBUG-data [yu_climatebug_2024], also focuses on corporate communications but in EU banks.
+Beyond corporate disclosures, Sustainable Signals [linSUSTAINABLESIGNALSijcai2023] introduces a dataset of product reviews, each classified based on its relevance to sustainability.
 In addition to focusing on consumer perspectives through reviews, this dataset also examines sustainability aspects in product descriptions, providing a broader understanding of how sustainability is communicated in the context of consumer products.
@@
 Human annotations ensure that the labels align with the task’s goals, something weak labels alone cannot achieve.
-Supervised Solutions
-The solutions proposed to tackle that task range from keyword-based models (Leippold and Varini, 2020; Bingler et al., 2023) to fine-tuning BERT-like models (Leippold and Varini, 2020; Webersinke et al., 2022; Garrido-Merchán et al., 2023; Bingler et al., 2023; Yu et al., 2024) (see Table LABEL:tab:models in the appendix for details).
-The performance of the best models is above 90% (Garrido-Merchán et al., 2023; Bingler et al., 2023; Yu et al., 2024) (see Table 12 in the appendix for details).
+Models
+The solutions proposed to tackle that task range from keyword-based models [varini_climatext_2020, bingler2023cheaptalkspecificitysentiment] to fine-tuning BERT-like models [varini_climatext_2020, nicolas_webersinke_climatebert_2021, garridomerchán2023finetuning, bingler2023cheaptalkspecificitysentiment, yu_climatebug_2024].
+The performance of the best models is above 90% [garridomerchán2023finetuning, bingler2023cheaptalkspecificitysentiment, yu_climatebug_2024].
+[^3] See Table 13 in the appendix for details.
 Insights
@@
 Moreover, relying only on weak labels might result in tasks that fail to capture the complexities of the subject.
-4.3. Thematic Analysis
+4.3 Thematic Analysis
 Once a text is known to be related to climate, one aims to know the exact topic of the text.
@@
 Task Description
-Given an input sentence or a paragraph, output a subtopic related to climate change.
-This is a multiclass classification task that can be either supervised or unsupervised, depending on the availability of labeled data.
+Given an input sentence or paragraph, output a subtopic related to climate change.
+This is a multiclass classification task that can be supervised or unsupervised, depending on the availability of labeled data.
 Alternatively, it can be framed as a clustering task with the goal of discovering latent subtopic structures.
+Table 5: Summary of datasets, labels, scopes, and sources for climate-related text classification tasks.
 Dataset Input Labels Label details
-TCFD rec. (Bingler et al., 2022) paragraphs from corporate annual reports TCFD 4 main categories: Metrics and Targets, Risk Management, Strategy, Governance and General (see appendix A for details)
-TCFD (Sampson et al., 2022) paragraphs from regulatory and discretionary reports TCFD 11 recommendations: Metrics and Targets (board’s oversight, management’s role), Risk Management (identified risk, impact, resilience), … (see Appendix A for details)
-FineBERT’s ESG (Huang et al., 2020) Sentences from 10-K Environmental, Social, Governance, General Environmental - e.g., climate change, natural capital, pollution and waste, and environmental opportunities
-ESGBERT’s ESG (Schimanski et al., 2023b) Sentences from reports and corporate news Environment, Social, Governance and None Environmental criteria comprise a company’s energy use, waste management, pollution, […] as well as compliance with governmental regulations.
+TCFD rec. bingler_cheap_2021 Paragraphs from corporate annual reports TCFD 4 main categories: Metrics and Targets, Risk Management, Strategy, Governance and General (see appendix A for details)
+TCFD sampson_tcfd-nlp_nodate Paragraphs from regulatory and discretionary reports TCFD 11 recommendations: Metrics and Targets (board’s oversight, management’s role), Risk Management (identified risk, impact, resilience), … (see Appendix A for details)
+FineBERT’s ESG huangFinBERTLargeLanguage2020 Sentences from 10-K Environmental, Social, Governance, General Environmental - e.g., climate change, natural capital, pollution and waste, and environmental opportunities
+ESGBERT’s ESG schimanski_bridging_2023 Sentences from reports and corporate news Environment, Social, Governance and None Environmental criteria comprise a company’s energy use, waste management, pollution, […] as well as compliance with governmental regulations.
 Special areas of interest are climate change and environmental sustainability.
-Multilingual ESG classification (Lee and Kim, 2023) Sentences from Korean reports (English and Korean) Environment, Social, Governance, Neutral, and Irrelevant Environmental factors include the reduction of hazardous substances, eco-friendly management, climate change, carbon emission, natural resources, […]
-SciDCC (Mishra and Mittal, 2021) News articles (Title, Summary, Body) Environment, Geology, Animals, Ozone Layer, Climate, etc. Category in which the article was published (Automatic Label)
-Transaction Ledger (Jain et al., 2023) Transaction ledger entry (description of transaction) Accounting Adjustments, Administration, Advertising, Benefits & Insurance, etc.
-Standardized commodity classes
-ClimateEng (Vaid et al., 2022) Tweets posted during COP25 filtered by keywords (relevant to climate-change) Ocean/Water, Politics, Disaster, Agriculture/Forestry, General Sub-categories of climate-change
-ESGBERT’s Nature (Schimanski et al., 2024a) Paragraphs from reports General, Nature, Biodiversity, Forest, Water Multi-label Nature-related topics
-ClimaTOPIC (Spokoyny et al., 2023) CDP responses (short texts) Adaptation, Buildings, Climate Hazards, Emissions, Water, etc. Category of the question (Automatic Label)
-Table 3. Summary of datasets, labels, scopes, and sources for climate-related text classification tasks.
-4.3.1. Classifying Climate-Related Financial Disclosure (TCFD)
-The Task Force on Climate-Related Financial Disclosures (TCFD) proposes four main categories of climate-related disclosure (Governance, Strategy, Risk Management, and Metrics and Targets) and 11 recommendations for climate-related disclosures (see Appendix A for details).
+Multilingual ESG classification LEE2023119726 Sentences from Korean reports (English and Korean) Environment, Social, Governance, Neutral, and Irrelevant Environmental factors include the reduction of hazardous substances, eco-friendly management, climate change, carbon emission, natural resources, […]
+SciDCC mishra2021neuralnere News articles (Title, Summary, Body) Environment, Geology, Animals, Ozone Layer, Climate, etc. Category in which the article was published (Automatic Label)
+Transaction Ledger jain_supply_2023 Transaction ledger entry (description of transaction) Accounting Adjustments, Administration, Advertising, Benefits & Insurance, etc. Standardized commodity classes
+ClimateEng vaid-etal-2022-towards Tweets posted during COP25 filtered by keywords (relevant to climate-change) Ocean/Water, Politics, Disaster, Agriculture/Forestry, General Sub-categories of climate-change
+ESGBERT’s Nature Schimanski2024nature Paragraphs from reports General, Nature, Biodiversity, Forest, Water Multi-label Nature-related topics
+ClimaTOPIC spokoyny2023answering CDP responses (short texts) Adaptation, Buildings, Climate Hazards, Emissions, Water, etc. Category of the question (Automatic Label)
+4.3.1 Classifying Climate-Related Financial Disclosure (TCFD)
+The Task Force on Climate-Related Financial Disclosures (TCFD) proposes four main categories of climate-related disclosure (Governance, Strategy, Risk Management, and Metrics and Targets) and 11 recommendations for climate-related disclosures.444see Appendix A for details.
 Finding texts associated to these categories in company communication can help identify gaps and inconsistencies in reporting.
-Unsupervised Solutions
-Ding et al.
-(2023) introduced a score to quantify the extent of climate-related content in corporate disclosures, as well as 4 TCFD-category-based similarity scores.
+Annotation-free
+dingCarbonEmissionsTCFD2023 introduced a score to quantify the extent of climate-related content in corporate disclosures, as well as 4 TCFD-category-based similarity scores.
 The score is based on sentence similarity with sentences from TCFD referential documents.
 They identified discrepancies that could signal greenwashing by comparing these scores with actual carbon emission data.
-Auzepy et al.
-(2023) examined reports from banks endorsing the TCFD recommendations.
+auzepy_evaluating_2023 examined reports from banks endorsing the TCFD recommendations.
 They proposed a fine-grained analysis of TCFD categories using a zero-shot entailment classification method on sentences.
@@
 Although unsupervised methods can be employed for thematic analysis, recent studies have demonstrated the value of curated datasets for categorizing corporate disclosures.
-For instance, Bingler et al.
-(2022) leveraged report headings to annotate paragraphs into the four main TCFD categories.
-Similarly, Sampson et al.
-(2022) assessed the quality of climate-related disclosures using a dataset of 162k sentences, where each sentence was labeled according to the TCFD recommendations by experts through a question answering (QA) process.
-4.3.2. Classifying Environmental, Social, and Governance Disclosure (ESG)
+For instance, bingler_cheap_2021 leveraged report headings to annotate paragraphs into the four main TCFD categories.
+Similarly, sampson_tcfd-nlp_nodate assessed the quality of climate-related disclosures using a dataset of 162k sentences, where each sentence was labeled according to the TCFD recommendations by experts through a question answering (QA) process.
+4.3.2 Classifying Environmental, Social, and Governance Disclosure (ESG)
 Beyond climate-specific reporting, Environmental, Social, and Governance (ESG) disclosures represent a broader set of non-financial metrics that evaluate a company’s sustainability performance.
@@
 The standardized nature of ESG metrics makes them valuable for computational models that aim to detect inconsistencies in corporate claims, such as overstating environmental impacts through greenwashing.
-However, as noted in Berg et al.
-(2022)’s aggregated analysis of ESG ratings, discrepancies persist across different evaluation frameworks.
+However, as noted in berg2022aggregate’s aggregated analysis of ESG ratings, discrepancies persist across different evaluation frameworks.
 These inconsistencies raise the question of whether artificial intelligence can be harnessed to improve the consistency and reliability of ESG assessments.
-Unsupervised Solutions
-Rouen et al.
-(2023) proposed an ML algorithm to describe the content of ESG reports.
+Annotation-free
+rouenEvolutionESGReports2023 proposed an ML algorithm to describe the content of ESG reports.
 They used industry-topic dictionaries to compute the topic frequency using a TF-IDF-based algorithm.
 They used that classification to identify selective disclosure.
-Mehra et al.
-(2022) experimented with sentence similarity to extract sentences that were the most relevant to environmental factors, and (Bronzini et al., 2023) selected sentences related to ESG using INSTRUCTOR-xl (Su et al., 2023).
+Mehra_2022 experimented with sentence similarity to extract sentences that were the most relevant to environmental factors, and bronzini_glitter_2023 selected sentences related to ESG using INSTRUCTOR-xl [su2023embeddertaskinstructionfinetunedtext].
 They both subsequently used the extracted sentences to predict ESG scores.
 Annotated datasets
-Multiple works (Huang et al., 2020; Schimanski et al., 2023b; Lee and Kim, 2023) proposed datasets for topic classification with the labels: Environment, Social, Governance and General.
-Huang et al.
-(2020) focused on reports while Schimanski et al.
-(2023b) added corporate news.
-Schimanski et al.
-(2023b) introduced a larger dataset for pre-training domain-specific models and an evaluation dataset of 2,000 sentences.
+Multiple works huangFinBERTLargeLanguage2020, schimanski_bridging_2023, LEE2023119726 proposed datasets for topic classification with the labels: Environment, Social, Governance and General.
+huangFinBERTLargeLanguage2020 focused on reports while schimanski_bridging_2023 added corporate news.
+schimanski_bridging_2023 introduced a larger dataset for pre-training domain-specific models and an evaluation dataset of 2,000 sentences.
 While most ESG-related reports are typically available in English, some local companies provide them only in their native language.
-To partially address this, Lee and Kim (2023) proposed a multilingual dataset constructed from Korean corporate reports.
-4.3.3. Other Topics
+To partially address this, LEE2023119726 proposed a multilingual dataset constructed from Korean corporate reports.
+4.3.3 Other Topics
 Although TCFD and ESG are established frameworks for analyzing corporate communications, many other topics may be of interest.
 These topics can follow other frameworks such as CDP, but the thematic analysis can also be unsupervised, effectively discovering the topics mentioned.
-Unsupervised Solutions
-Kang and Kim (2022) analyzed the content of sustainable reports with thematic analysis.
+Annotation-free
+hyewon_kang_analyzing_2022 analyzed the content of sustainable reports with thematic analysis.
 It was conducted using sentence similarity between the reports and content from the SDGs website.
-Yu et al.
-(2024) examined the themes covered in the climate-related sections of bank reports by employing a dual approach: using hand-selected keywords to target anticipated topics and applying clustering techniques to group climate-related sentences and uncover underlying themes.
-Brié et al.
-(2022) leveraged ClimateBERT to extract climate-related paragraphs from corporate documents.
-They then applied a Structural Topic Model (STM) (Margaret E.
-Roberts and Airoldi, 2016) to identify thematic clusters within the data.
-This STM-based approach was also used by Fortes and Le Guenedal (2020) to identify influential topics for EUR/USD exchange rate.
+yu_climatebug_2024 examined the themes covered in the climate-related sections of bank reports by employing a dual approach: using hand-selected keywords to target anticipated topics and applying clustering techniques to group climate-related sentences and uncover underlying themes.
+bjarne_brie_mandatory_2022 leveraged ClimateBERT to extract climate-related paragraphs from corporate documents.
+They then applied a Structural Topic Model (STM) [STM] to identify thematic clusters within the data.
+This STM-based approach was also used by fortes2020tracking to identify influential topics for EUR/USD exchange rate.
 Their findings highlighted a significant discrepancy between the European Central Bank (ECB) and the Federal Reserve (FED), with the ECB attributing greater importance to climate change in the context of financial stability than the FED.
@@
 Annotated datasets
-Spokoyny et al.
-(2023) introduced ClimaTOPIC, a topic classification dataset derived from responses to the CDP questionnaires, organized into 12 categories that align with the thematic structure of the CDP questions.
+spokoyny2023answering introduced ClimaTOPIC, a topic classification dataset derived from responses to the CDP questionnaires, organized into 12 categories that align with the thematic structure of the CDP questions.
 This can be used to identify sections relevant to a particular CDP topic (e.g., Emissions, Adaptation, or Energy), enabling the extraction of structured information from unstructured documents.
-Schimanski et al.
-(2024a) presents a dataset focused on nature-related topics (such as Water, Forest, and Biodiversity) extracted from annual reports, sustainability reports, and earnings call transcripts (ESGBERT Nature).
+Schimanski2024nature presents a dataset focused on nature-related topics (such as Water, Forest, and Biodiversity) extracted from annual reports, sustainability reports, and earnings call transcripts (ESGBERT Nature).
 While less explicitly linked to climate change, these topics often reflect the consequences of environmental degradation associated with climate issues.
 While NLP is often used to analyze corporate report texts, it can also be used to analyze tables.
-For example, Jain et al.
-(2023) proposes estimating Scope 3 emissions  [^3] Scope 3 emissions refer to all indirect greenhouse gas (GHG) emissions that occur in a company’s value chain, excluding emissions from the company’s own operations (Scope 1) and its purchased electricity (Scope 2).
+For example, jain_supply_2023 proposes estimating Scope 3 emissions  [^5] Scope 3 emissions refer to all indirect greenhouse gas (GHG) emissions that occur in a company’s value chain, excluding emissions from the company’s own operations (Scope 1) and its purchased electricity (Scope 2).
 based on corporate transactions (see Section 4.10 for details).
@@
 As previously discussed, social media and news articles offer a broader perspective for cross-referencing corporate communications.
-For example, Vaid et al.
-(2022) introduced ClimateEng, a climate-related topic classification dataset composed of tweets grouped into five categories: Disaster, Ocean/Water, Agriculture/Forestry, Politics, and General.
+For example, vaid-etal-2022-towards introduced ClimateEng, a climate-related topic classification dataset composed of tweets grouped into five categories: Disaster, Ocean/Water, Agriculture/Forestry, Politics, and General.
 This dataset can be used to detect greenwashing by comparing the public’s concerns with the narrative presented in corporate sustainability reports.
-Similarly, Mishra and Mittal (2021) published the Science Daily Climate Change dataset (SciDCC), consisting of 11,000 news articles grouped into 20 climate-related categories.
+Similarly, mishra2021neuralnere published the Science Daily Climate Change dataset (SciDCC), consisting of 11,000 news articles grouped into 20 climate-related categories.
 This dataset offers another avenue for identifying greenwashing, as companies’ communications can be cross-referenced with external news coverage, potentially highlighting contradictions or omissions in corporate reports.
-4.3.4. Models and Conclusion
-Supervised Solutions
-The most common solution proposed for these tasks is fine-tuning a Transformer model (Huang et al., 2020; Schimanski et al., 2023b; Lee and Kim, 2023; Bingler et al., 2022; Sampson et al., 2022; Vaid et al., 2022; Spokoyny et al., 2023; Schimanski et al., 2024a; Jain et al., 2023).
-To contextualize the performance of Transformer models, some studies also report performances of classical baselines (Huang et al., 2020; Spokoyny et al., 2023; Schimanski et al., 2024a; Jain et al., 2023; Bingler et al., 2022; Sampson et al., 2022).
-Bingler et al.
-(2022) proposed a custom approach combining logistic regression with features from a fine-tuned language model, while Sampson et al.
-(2022) also evaluated clustering techniques and stacked models.
+4.3.4 Models and Conclusion
+Models
+The most common solution proposed for these tasks is fine-tuning a Transformer model [huangFinBERTLargeLanguage2020, schimanski_bridging_2023, LEE2023119726, bingler_cheap_2021, sampson_tcfd-nlp_nodate, vaid-etal-2022-towards, spokoyny2023answering, Schimanski2024nature, jain_supply_2023].
+To contextualize the performance of Transformer models, some studies also report performances of classical baselines [huangFinBERTLargeLanguage2020, spokoyny2023answering, Schimanski2024nature, jain_supply_2023, bingler_cheap_2021, sampson_tcfd-nlp_nodate].
+bingler_cheap_2021 proposed a custom approach combining logistic regression with features from a fine-tuned language model, while sampson_tcfd-nlp_nodate also evaluated clustering techniques and stacked models.
 The performance of the best-performing model is systematically high (above 80%), except on SciDCC and ClimaTOPIC.
-Classical approaches (SVM, NB) also reach good performances, under-performing by less than 20% on all datasets (except on Bingler et al.
-(2022)’s dataset).
-The TF-IDF baseline even outperformed fine-tuned Transformers on Sampson et al.
-(2022)’s dataset.
+Classical approaches (SVM, NB) also reach good performances, under-performing by less than 20% on all datasets (except on bingler_cheap_2021’s dataset).
+The TF-IDF baseline even outperformed fine-tuned Transformers on sampson_tcfd-nlp_nodate’s dataset.
 This shows that those topics have distinguishable vocabularies.
 Automatic Labels
-The performance gap observed in Bingler et al.
-(2022) may be attributed to using paragraph-level inputs rather than sentence-level inputs aggregated for classification, as fine-tuned models also achieve around 20% precision at the paragraph level.
-The lower performances on SciDCC(Mishra and Mittal, 2021) and ClimaTOPIC(Spokoyny et al., 2023) are likely caused by the automation of the labeling process.
+The performance gap observed in bingler_cheap_2021 may be attributed to using paragraph-level inputs rather than sentence-level inputs aggregated for classification, as fine-tuned models also achieve around 20% precision at the paragraph level.
+The lower performances on SciDCCmishra2021neuralnere and ClimaTOPICspokoyny2023answering are likely caused by the automation of the labeling process.
 The labels are therefore not designed to be labels.
@@
 For ClimaTOPIC, the labels are question categories, which are designed to group questions not to precisely identify them.
-Therefore a question about ”the emission of a building” might fit in either Emissions or Building, yet it is assigned only one label.
+Therefore, a question about ”the emission of a building” might fit in either Emissions or Building, yet it is assigned only one label.
 Insights
@@
 Future research could investigate topics that are lexically similar but distinct in their subject or focus.
-Topic detection serves as a valuable tool for structuring documents and has been utilized to analyze the extent to which companies address specific topics, enabling the detection of selective disclosure (Bingler et al., 2023, 2022).
+Topic detection serves as a valuable tool for structuring documents and has been utilized to analyze the extent to which companies address specific topics, enabling the detection of selective disclosure bingler2023cheaptalkspecificitysentiment, bingler_cheap_2021.
 However, it is critical to ensure that labels are well-defined and data is accurately annotated.
@@
 These works might be revisited to conduct human annotations.
-4.4. In-depth Disclosure: Climate Risk Classification
+4.4 In-depth Disclosure: Climate Risk Classification
 Climate change can bring both risks and opportunities for companies.
@@
 Some works focus only on risks, classifying them into types of risks (e.g., physical risk, reputational risk, regulatory risk, or transition risk)
+Table 6: Summary of datasets, labels, scopes, and sources for climate-related Risk classification.
 Dataset Input Labels
-Ask BERT’s Climate Risk (Kölbel et al., 2021) Sentences from TCFD’s example reports and non-climate-related sentences Risk type: Transition risk, physical risk, and general risk (no guidelines)
-Climate Risk (Friederich et al., 2021) Paragraphs from European companies annual reports Risk type:Acute, Chronic, Policy & legal, Tech & Market, Reputational, and Negative (no guidelines)
-ClimateBERT’s Sentiment (Bingler et al., 2023) Paragraphs from companies’ annual reports Risk or threat that negatively impacts an entity of interest (negative sentiment); or Opportunity arising due to climate change (positive sentiment); Neutral otherwise.
-Sentiment Analysis (Xiang and Fujii, 2023) paragraphs from academic texts on climate change and health published between 2013 and 2020 Risk (negative) if it discusses climate change causing public health issues, serious consequences, or worsening trends, greenwashing.
+Ask BERT’s Climate Risk kolbel_ask_2021 Sentences from TCFD’s example reports and non-climate-related sentences Risk type: Transition risk, physical risk, and general risk (no guidelines)
+Climate Risk Friederich_climate_risk_disclosure Paragraphs from European companies annual reports Risk type:Acute, Chronic, Policy & legal, Tech & Market, Reputational, and Negative (no guidelines)
+ClimateBERT’s Sentiment bingler2023cheaptalkspecificitysentiment Paragraphs from companies’ annual reports Risk or threat that negatively impacts an entity of interest (negative sentiment); or Opportunity arising due to climate change (positive sentiment); Neutral otherwise.
+Sentiment Analysis xiang_dare_2023 Paragraphs from academic texts on climate change and health published between 2013 and 2020 Risk (negative) if it discusses climate change causing public health issues, serious consequences, or worsening trends, greenwashing.
 Opportunity (positive): highlights potential benefits, positive actions, or research addressing gaps.
 Neutral otherwise.
-Table 4. Summary of datasets, labels, scopes, and sources for climate-related Risk classification.
-Unsupervised Solutions
+Annotation-free
 Similarly to climate-related detection, climate risk classification has been tackled with keyword-based approaches.
-Li et al.
-(2020), Kheradmand et al.
-(2021) and Chou et al.
-(2023) proposed using dictionaries of words related to climate risk to identify paragraphs dealing with climate risk.
-Li et al.
-(2020) constructed risk measures based on the frequency of the terms in the risk dictionaries.
-Chou et al.
-(2023) analyzed the topics mentioned in conjunction with physical and transition risks.
-Sautner et al.
-(2023) proposed using one dictionary for risk and one for opportunity classification.
+liCorporateClimateRisk2020, kheradmand2021a and chou_ESG proposed using dictionaries of words related to climate risk to identify paragraphs dealing with climate risk.
+liCorporateClimateRisk2020 constructed risk measures based on the frequency of the terms in the risk dictionaries.
+chou_ESG analyzed the topics mentioned in conjunction with physical and transition risks.
+SAUTNER_cliamte_change_exp proposed using one dictionary for risk and one for opportunity classification.
 Annotated datasets
 Several studies have introduced specialized annotated datasets for risk/opportunity classification.
-For instance, Kölbel et al.
-(2021) created a dataset for climate-related risk classification with three categories: physical risk, transition risk, and general risk, using active learning for annotation.
-Similarly, Friederich et al.
-(2021) developed an annotated dataset for risk classification with five labels, covering acute physical risk, chronic physical risk, policy and legal risks, technology and market risks, and reputational transition risks.
-Bingler et al.
-(2023) released a dataset as part of the ClimateBERT downstream tasks, focusing on classifying paragraphs from corporate reports into three categories: opportunity, neutral, or risk.
-Extending beyond corporate disclosures, Xiang and Fujii (2023) compiled a climate-related risk/opportunity classification dataset of academic texts from the Web of Science and Scopus.
-Supervised Solutions
-The solutions proposed are similar to the other tasks: fine-tuned Transformer models (Kang and Kim, 2022; Kölbel et al., 2021; Friederich et al., 2021; Bingler et al., 2023; Webersinke et al., 2022; Xiang and Fujii, 2023) and keyword-based-features with simple models (Kölbel et al., 2021; Friederich et al., 2021; Bingler et al., 2023) (see Table LABEL:tab:models in the appendix for details).
-The notable exception is Xiang and Fujii (2023), who evaluated LSTM-based solutions alongside the Transformers models.
-Most studies (Kang and Kim, 2022; Kölbel et al., 2021; Bingler et al., 2023; Webersinke et al., 2022; Xiang and Fujii, 2023) reported high performances, above 80%.
-Kölbel et al.
-(2021) and Bingler et al.
-(2023) also reported good performances for the keyword-based baselines (72% and 84%).
-The performance reported by Friederich et al.
-(2021) on climate risk tasks highlights distinct challenges across subtasks.
+For instance, kolbel_ask_2021 created a dataset for climate-related risk classification with three categories: physical risk, transition risk, and general risk, using active learning for annotation.
+Similarly, Friederich_climate_risk_disclosure developed an annotated dataset for risk classification with five labels, covering acute physical risk, chronic physical risk, policy and legal risks, technology and market risks, and reputational transition risks.
+bingler2023cheaptalkspecificitysentiment released a dataset as part of the ClimateBERT downstream tasks, focusing on classifying paragraphs from corporate reports into three categories: opportunity, neutral, or risk.
+Extending beyond corporate disclosures, xiang_dare_2023 compiled a climate-related risk/opportunity classification dataset of academic texts from the Web of Science and Scopus.
+Models
+The solutions proposed are similar to the other tasks: fine-tuned Transformer models hyewon_kang_analyzing_2022, kolbel_ask_2021, Friederich_climate_risk_disclosure, bingler2023cheaptalkspecificitysentiment, nicolas_webersinke_climatebert_2021, xiang_dare_2023 and keyword-based-features with simple models kolbel_ask_2021, Friederich_climate_risk_disclosure, bingler2023cheaptalkspecificitysentiment.
+The notable exception is xiang_dare_2023, who evaluated LSTM-based solutions alongside the Transformers models.
+Most studies [hyewon_kang_analyzing_2022, kolbel_ask_2021, bingler2023cheaptalkspecificitysentiment, nicolas_webersinke_climatebert_2021, xiang_dare_2023] reported high performances, above 80%.
+kolbel_ask_2021 and bingler2023cheaptalkspecificitysentiment also reported good performances for the keyword-based baselines (72% and 84%).
+The performance reported by Friederich_climate_risk_disclosure on climate risk tasks highlights distinct challenges across subtasks.
 When tasked with identifying the specific type of risk in sentences already known to be about climate risks, word-based models outperformed fine-tuned Transformers.
@@
 It is important to distinguish risk/opportunity classification from sentiment analysis.
-While risk/opportunity classification is inspired by sentiment analysis (Bingler et al., 2023), they play a complementary role.
-A statement such as “With proactive climate risk management, we are ready to tackle extreme weather disruptions, ensuring resilience” is classified as positive using a sentiment analysis model (Pérez et al., 2021), and as mentioning a risk using a risk/opportunity classification model (Bingler et al., 2023).
+While risk/opportunity classification is inspired by sentiment analysis bingler2023cheaptalkspecificitysentiment, they play a complementary role.
+A statement such as “With proactive climate risk management, we are ready to tackle extreme weather disruptions, ensuring resilience” is classified as positive using a sentiment analysis model perez2021pysentimiento, and as mentioning a risk using a risk/opportunity classification model bingler2023cheaptalkspecificitysentiment.
 Sentiment analysis focuses on the form of the statement, while risk is about the content.
-As the literature on sentiment analysis (Wankhade et al., 2022) is broad, the analysis of the tone of the statement can be done using existing models (Polignano et al., 2022; Kang and Kim, 2022) complementing approaches on risk classification.
+As the literature on sentiment analysis Wankhade2022 is broad, the analysis of the tone of the statement can be done using existing models marco_polignano_nlp_2022, hyewon_kang_analyzing_2022 complementing approaches on risk classification.
 Insights
 The performance of fine-tuned models might indicate that the task of identifying texts talking about risks is solved.
-However, the work by Friederich et al.
-(2021) shows limitations when experimenting with heavily imbalanced datasets, which actually correspond to real-world settings.
+However, the work by Friederich_climate_risk_disclosure shows limitations when experimenting with heavily imbalanced datasets, which actually correspond to real-world settings.
 Future research could focus on such skewed settings.
-4.5. Green Claim Detection
+4.5 Green Claim Detection
 Identifying greenwashing involves more than just finding text related to climate change, as a company might simply state factual information about the climate without making any commitments on it.
-Hence, we now focus on detecting green claims, i.e., claims that a product, service, or corporate practice either contributes positively to environmental sustainability or is less harmful to the environment than alternatives (Stammbach et al., 2023; Woloszyn et al., 2021).
+Hence, we now focus on detecting green claims, i.e., claims that a product, service, or corporate practice either contributes positively to environmental sustainability or is less harmful to the environment than alternatives stammbach_environmental_2023, vinicius_woloszyn_towards_2021.
 The European Commission calls these claims “environmental claims”:
+Table 7: Label definitions of the datasets on green claims detection.
 Dataset Input Positive Label description
-Green Claims (Woloszyn et al., 2021) Marketing Tweets Environmental (or green) advertisements refer to all appeals that include ecological, environmental sustainability, or nature-friendly messages that target the needs and desires of environmentally concerned stakeholders.
-Environmental Claims (Stammbach et al., 2023) Paragraph from reports Environmental claims refer to the practice of suggesting or otherwise creating the impression […] that a product or a service is environmentally friendly (i.e., it has a positive impact on the environment) or is less damaging to the environment than competing goods or services […] In our case, claims relate to products, services, or specific corporate environmental performance.
-Table 5. Label definitions of the datasets on green claims detection.
+Green Claims vinicius_woloszyn_towards_2021 Marketing Tweets Environmental (or green) advertisements refer to all appeals that include ecological, environmental sustainability, or nature-friendly messages that target the needs and desires of environmentally concerned stakeholders.
+Environmental Claims stammbach_environmental_2023 Paragraph from reports Environmental claims refer to the practice of suggesting or otherwise creating the impression […] that a product or a service is environmentally friendly (i.e., it has a positive impact on the environment) or is less damaging to the environment than competing goods or services […] In our case, claims relate to products, services, or specific corporate environmental performance.
 Task Description
@@
 Annotated datasets
-The most notable approaches aiming at identifying climate-related claims introduced annotated datasets (Woloszyn et al., 2021; Stammbach et al., 2023).
-Woloszyn et al.
-(2021) proposed to focus on social media marketing through the detection of green claims in tweets.
-Stammbach et al.
-(2023) proposed a dataset for environmental claim detection in paragraphs from corporate reports.
-Supervised Solutions
-Both Stammbach et al.
-(2023) and Woloszyn et al.
-(2021) evaluated fine-tuned Transformer models on the datasets.
-Stammbach et al.
-(2023) also evaluated classical approaches and experimented with a model fine-tuned on general claim detection (Arslan et al., 2020) applied to their domain-specific task (see Table LABEL:tab:models in the Appendix for details).
-Woloszyn et al.
-(2021) and Stammbach et al.
-(2023) demonstrated good performances with fine-tuned Transformer (above 84%), yet they have very different Inter-Annotator Agreement (IAA), with Krippendorff’s $\alpha=0.8223$ for Woloszyn et al.
-(2021) and $\alpha=0.47$ for Stammbach et al.
-(2023).
+The most notable approaches aiming at identifying climate-related claims introduced annotated datasets vinicius_woloszyn_towards_2021, stammbach_environmental_2023.
+vinicius_woloszyn_towards_2021 proposed to focus on social media marketing through the detection of green claims in tweets.
+stammbach_environmental_2023 proposed a dataset for environmental claim detection in paragraphs from corporate reports.
+Models
+Both stammbach_environmental_2023 and vinicius_woloszyn_towards_2021 evaluated fine-tuned Transformer models on the datasets.
+stammbach_environmental_2023 also evaluated classical approaches and experimented with a model fine-tuned on general claim detection Arslan_Hassan_Li_Tremayne_2020 applied to their domain-specific task.
+vinicius_woloszyn_towards_2021 and stammbach_environmental_2023 demonstrated good performances with fine-tuned Transformer (above 84%), yet they have very different Inter-Annotator Agreement (IAA), with Krippendorff’s $\alpha=0.8223$ for vinicius_woloszyn_towards_2021 and $\alpha=0.47$ for stammbach_environmental_2023.
 This might be explained by the difference in context: marketing tweets might be easier to understand for humans and more self-sufficient compared to paragraphs extracted from reports.
-Additionally, Woloszyn et al.
-(2021) experimented with adversarial attack, showing the sensibility to character-swap and word-swap.
+Additionally, vinicius_woloszyn_towards_2021 experimented with adversarial attack, showing the sensibility to character-swap and word-swap.
 This shows that the models tend to rely heavily on particular words, showing only a superficial understanding.
@@
 Insights
-Green claims have been analyzed in company reporting (Stammbach et al., 2023) and social media communication (Woloszyn et al., 2021).
+Green claims have been analyzed in company reporting [stammbach_environmental_2023] and social media communication [vinicius_woloszyn_towards_2021].
 Fine-tuned models can solve the task rather well, even if challenges remain:
-  • Inter annotator agreement: The annotator agreement remains low for green claim detection in reports (Stammbach et al., 2023).
+  • Inter annotator agreement: The annotator agreement remains low for green claim detection in reports [stammbach_environmental_2023].
   • Integration with existing claim detection literature: While claim detection is a well-established field with extensive literature, studies on environmental claims do not fully connect with existing research on claim detection.
-In particular, they do not distinguish between Claims, Verifiable Claims (claims that can be checked), and Check-worthy Claims (claims that are interesting to verify) (Panchendrarajan and Zubiaga, 2024).
+In particular, they do not distinguish between Claims, Verifiable Claims (claims that can be checked), and Check-worthy Claims (claims that are interesting to verify) [panchendrarajan2024claim].
   • Evaluation Sensitivity and Real-World Robustness: The literature shows that fine-tuned models are sensitive to adversarial attacks, meaning small perturbations of the text influence greatly the performance of the classifier.
@@
 Poor data quality might induce perturbations in the texts, reducing the performance of models.
-4.6. Green Claim Characteristics
+4.6 Green Claim Characteristics
 Once we have established that a sentence is climate-related (Section 4.2) and that it is a claim about the company (Section 4.5), we can endeavor to further classify the claim into fine-grained categories.
 Table 8 shows various characteristics of claims that have been studied.
+Table 8: Label definitions the datasets related to characterization of green claims.
 Dataset Input Labels
-Implicit/Explicit Green Claims (Woloszyn et al., 2021) Marketing Tweets Implicit green claims raise the same ecological and environmental concerns as explicit green claims (see definition in Section 4.5), but without showing any commitment from the company.
+Implicit/Explicit Green Claims vinicius_woloszyn_towards_2021 Marketing Tweets Implicit green claims raise the same ecological and environmental concerns as explicit green claims (see definition in Section 4.5), but without showing any commitment from the company.
 If the tweet does not contain a green claim then No Claim.
-Specificity (Bingler et al., 2023) Paragraph from reports A paragraph is Specific if it includes clear, tangible, and firm-specific details about events, goals, actions, or explanations that directly impact or clarify the firm’s operations, strategy, or objectives.
+Specificity bingler2023cheaptalkspecificitysentiment Paragraph from reports A paragraph is Specific if it includes clear, tangible, and firm-specific details about events, goals, actions, or explanations that directly impact or clarify the firm’s operations, strategy, or objectives.
 Non-specific otherwise.
-Commitments & Actions(Bingler et al., 2023) Paragraph from reports A paragraph is a commitment or an action if it contains targets for the future or actions already taken in the past.
-Net Zero/Reduction (Schimanski et al., 2023a) Paragraph from Net Zero Tracker (Lang et al., 2023) The paragraph contains either a Net-Zero target, a Reduction target or no target (None)
-Table 6. Label definitions the datasets related to characterization of green claims.
+Commitments & Actionsbingler2023cheaptalkspecificitysentiment Paragraph from reports A paragraph is a commitment or an action if it contains targets for the future or actions already taken in the past.
+Net Zero/Reduction tobias_schimanski_climatebert-netzero_2023 Paragraph from Net Zero Tracker [netzerotracker2023] The paragraph contains either a Net-Zero target, a Reduction target or no target (None)
 Task Description
@@
 Existing works characterized both the content and the form of claims.
-On the content dimension, there are datasets for identifying if the claim is about an action or a commitment (Bingler et al., 2023; Schimanski et al., 2023a).
-Bingler et al.
-(2023) introduce a dataset for the identification of commitments and actions (Yes/No), and Schimanski et al.
-(2023a) released one for the identification of reduction targets (net zero, reduction, general).
-On the form dimension, Woloszyn et al.
-(2021) characterized claims as implicit or explicit, and Bingler et al.
-(2023) annotated their dataset on the specificity of claims (Specific/Not specific).
-Bingler et al.
-(2023) ultimately used the specificity and commitment/action characteristics to identify cheap talks related to climate disclosure.
-Supervised Solutions
-Woloszyn et al.
-(2021); Schimanski et al.
-(2023a) and Bingler et al.
-(2023) experimented with fine-tuned Transformers.
-Bingler et al.
-(2023) also evaluated classical baselines (SVM, NB), and Schimanski et al.
-(2023a) experimented with GPT-3.5-turbo (see Table LABEL:tab:models in the Appendix for details).
+On the content dimension, there are datasets for identifying if the claim is about an action or a commitment [bingler2023cheaptalkspecificitysentiment, tobias_schimanski_climatebert-netzero_2023].
+bingler2023cheaptalkspecificitysentiment introduce a dataset for the identification of commitments and actions (Yes/No), and tobias_schimanski_climatebert-netzero_2023 released one for the identification of reduction targets (net zero, reduction, general).
+On the form dimension, vinicius_woloszyn_towards_2021 characterized claims as implicit or explicit, and bingler2023cheaptalkspecificitysentiment annotated their dataset on the specificity of claims (Specific/Not specific).
+bingler2023cheaptalkspecificitysentiment ultimately used the specificity and commitment/action characteristics to identify cheap talks related to climate disclosure.
+Models
+vinicius_woloszyn_towards_2021, tobias_schimanski_climatebert-netzero_2023 and bingler2023cheaptalkspecificitysentiment experimented with fine-tuned Transformers.
+bingler2023cheaptalkspecificitysentiment also evaluated classical baselines (SVM, NB), and tobias_schimanski_climatebert-netzero_2023 experimented with GPT-3.5-turbo.
 All approaches reached good performances, in particular fine-tuned Transformers reaching performances above 80%.
@@
 This low performance might be intrinsic to the task.
-Indeed, humans have disagreement when distinguishing specific and unspecific claims: Bingler et al.
-(2023) measured a low IAA on Specificity (Krippendorff’s $\alpha=0.1703$).
+Indeed, humans have disagreement when distinguishing specific and unspecific claims: bingler2023cheaptalkspecificitysentiment measured a low IAA on Specificity (Krippendorff’s $\alpha=0.1703$).
 This might indicate that the task is not well defined in the first place.
@@
 The high performances of fine-tuned models indicate that this type of task is solved.
-Furthermore, the demonstrated ability of GPT-3.5 as a zero-shot classifier (Schimanski et al., 2023a) highlights the potential of large language models to classify these characteristics without requiring extensive annotated datasets.
+Furthermore, the demonstrated ability of GPT-3.5 as a zero-shot classifier tobias_schimanski_climatebert-netzero_2023 highlights the potential of large language models to classify these characteristics without requiring extensive annotated datasets.
 These characteristics are particularly useful for identifying statements that may indicate greenwashing.
-For example, Bingler et al.
-(2023) utilized the proportion of non-specific commitments as a Cheap Talk Index demonstrating their practical applications.
+For example, bingler2023cheaptalkspecificitysentiment utilized the proportion of non-specific commitments as a Cheap Talk Index demonstrating their practical applications.
 However, characteristics such as specificity are inherently ambiguous and subjective, as evidenced by the low inter-annotator agreement (IAA).
 Therefore, future research could focus on disambiguating what constitutes a specific statement from a non-specific one more objectively.
-4.7. Green Stance Detection
+4.7 Green Stance Detection
 Beyond making claims about the environmental impact of their products and processes, companies contribute to environmental discussions and can have an impact on regulatory frameworks through their communications and industry presence.
 It is thus helpful to understand the stance of an organization on the existence and gravity of climate change, on climate mitigation and adaptation efforts, and on climate-related regulations.
+Table 9: Summary of datasets, labels, scopes, and sources for climate-related stance detection.
 Dataset Input Labels Label details
-ClimateFEVER (evidence)  (Leippold and Diggelmann, 2020) A claim and an evidence sentence from Wikipedia Support, Refutes, Not Enough Information Determines the relation between a claim and a single evidence sentence
-LobbyMap (Stance) (Morio and Manning, 2023) Page from a company communications (report, press release, …) Strongly supporting, Supporting, No or mixed position, Not supporting, Opposing Given the policy and the page, classifies the stance
-Global Warming Stance Detection (GWSD) (Luo et al., 2020) Sentences from news about global warming Stance of the evidence (Agree, Disagree, Neutral) toward the claim: Climate-Change is a serious concern.
-ClimateStance (Vaid et al., 2022) Tweets posted during COP25 filtered by keywords (relevant to climate-change) Stance towards climate change prevention: Favor, Against, Ambiguous.
+ClimateFEVER (evidence) [diggelmann_climate-fever_2020] A claim and an evidence sentence from Wikipedia Support, Refutes, Not Enough Information Determines the relation between a claim and a single evidence sentence
+LobbyMap (Stance) [morio2023an] Page from a company communications (report, press release, …) Strongly supporting, Supporting, No or mixed position, Not supporting, Opposing Given the policy and the page, classifies the stance
+Global Warming Stance Detection (GWSD) [luo_detecting_2020] Sentences from news about global warming Stance of the evidence (Agree, Disagree, Neutral) toward the claim: Climate-Change is a serious concern.
+ClimateStance [vaid-etal-2022-towards] Tweets posted during COP25 filtered by keywords (relevant to climate-change) Stance towards climate change prevention: Favor, Against, Ambiguous.
 (Stance used as a broad notion including sentiment, evaluation, appraisal, …)
-Stance on Remediation Effort (Lai and Chen, 2023) Texts extracted from the TCFD sections of the financial reports The text indicates support for climate change remediation efforts or refutation of such efforts.
+Stance on Remediation Effort [lai_using_2023] Texts extracted from the TCFD sections of the financial reports The text indicates support for climate change remediation efforts or refutation of such efforts.
 (No guidelines)
 Related subtask
-ClimateFEVER (claim)  (Leippold and Diggelmann, 2020) A claim and multiple evidence sentences from Wikipedia Support, Refutes, Debated, Not Enough Information Determines if a claim is supported by a set of retrieved evidence sentences
-LobbyMap (Page) (Morio and Manning, 2023) Page from a company communications (report, press release, …) 1/0 Contains a stance on a remediation policy
-LobbyMap (Query) (Morio and Manning, 2023) GHG emission regulation, Renewable energy, Carbon tax, … Classifies the remediation policy
-Table 7. Summary of datasets, labels, scopes, and sources for climate-related stance detection.
+ClimateFEVER (claim) [diggelmann_climate-fever_2020] A claim and multiple evidence sentences from Wikipedia Support, Refutes, Debated, Not Enough Information Determines if a claim is supported by a set of retrieved evidence sentences
+LobbyMap (Page) [morio2023an] Page from a company communication (report, press release, etc) Binary Contains a stance on a remediation policy
+LobbyMap (Query) [morio2023an] GHG emission regulation, Renewable energy, Carbon tax, etc) Classifies the remediation policy
 Task Description
@@
 if they were collected manually).
-In order to build these datasets, research can rely on simple heuristics such as downloading all tweets published during the COP25 filtered with keywords (Vaid et al., 2022).
+In order to build these datasets, research can rely on simple heuristics such as downloading all tweets published during the COP25 filtered with keywords vaid-etal-2022-towards.
 However, other researchers performed a more elaborate procedure.
-Leippold and Diggelmann (2020) proposed a pipeline for collecting evidence.
+diggelmann_climate-fever_2020 proposed a pipeline for collecting evidence.
 To retrieve relevant evidence from Wikipedia for a given claim, the pipeline involves three steps: document-level retrieval using entity-linking and BM25 to identify top articles, sentence-level retrieval using sentence embeddings trained on the FEVER dataset to extract relevant sentences, and sentence re-ranking using a pretrained ALBERT model to classify and rank evidence based on relevance.
 The second one encompasses the broader process: identifying claims and finding evidence, before predicting the relation.
-Wang et al.
-(2021) used a generic claim detection model trained on ClaimBuster (Arslan et al., 2020) for the claims and Google Search API to collect evidence.
+Wang2021EvidenceBA used a generic claim detection model trained on ClaimBuster Arslan_Hassan_Li_Tremayne_2020 for the claims and Google Search API to collect evidence.
 Finally, the last subtask is to train multiple models on each step: identifying evidences, identifying the claim and classifying the stance.
-This is the approach used by (Morio and Manning, 2023).
+This is the approach used by morio2023an.
 Annotated datasets
-As described previously, researchers built datasets of evidence related to one claim while focusing stance in tweets on seriousness of climate change such as Global Warming Stance Detection (GWSD) (Luo et al., 2020), ClimateStance (Vaid et al., 2022) stance on climate change of tweets posted during the COP25, or Lai and Chen (2023)’s dataset which focuses on the stance toward climate change remediation efforts.
-Another research initiative built a dataset of claim-evidence pairs, ClimateFEVER, and then trained a model on the stance classification (Leippold and Diggelmann, 2020; Wang et al., 2021).
-Finally, Morio and Manning (2023) proposed a dataset to assess corporate policy engagement built upon LobbyMap, which tackles the 3 steps: finding pages with evidence, identifying the claims targeted by that evidence, and classifying the stance.
-Supervised Solutions
-The solution proposed for to tackle stance classification are fine-tuned Transformer models (Vaid et al., 2022; Vaghefi et al., 2022; Xiang and Fujii, 2023; Morio and Manning, 2023; Webersinke et al., 2022; Wang et al., 2021; Spokoyny et al., 2023; Lai and Chen, 2023; Luo et al., 2020) and classical approaches (Spokoyny et al., 2023; Morio and Manning, 2023; Luo et al., 2020).
-Vaid et al.
-(2022) also experimented with FastText and Xiang and Fujii (2023) with LSTM-based models (See Table LABEL:tab:models in the Appendix for details).
-The performances are quite heterogeneous, Lai and Chen (2023) reaching F1-scores around 90% on classification of stance on remediation efforts, Luo et al.
-(2020) around 72% on GWSD, while performance on ClimateStance could not exceed 60%.
-Performances on ClimateFEVER are also quite low, however, when selecting only non-ambiguous examples, Xiang and Fujii (2023) could reach performances around 80%.
-Finally, performances on LobbyMap (Morio and Manning, 2023) are quite low (between 31% and 57.3% depending on the strictness of the metric).
-Overall, the performances show that the datasets are challenging (see Table 19 in the Appendix for details).
+As described previously, researchers built datasets of evidence related to one claim while focusing stance in tweets on seriousness of climate change such as Global Warming Stance Detection (GWSD) luo_detecting_2020, ClimateStance vaid-etal-2022-towards stance on climate change of tweets posted during the COP25, or lai_using_2023’s dataset which focuses on the stance toward climate change remediation efforts.
+Another research initiative built a dataset of claim-evidence pairs, ClimateFEVER, and then trained a model on the stance classification diggelmann_climate-fever_2020, Wang2021EvidenceBA.
+Finally, morio2023an proposed a dataset to assess corporate policy engagement built upon LobbyMap, which tackles the 3 steps: finding pages with evidence, identifying the claims targeted by that evidence, and classifying the stance.
+Models
+The solution proposed for to tackle stance classification are fine-tuned Transformer models vaid-etal-2022-towards, vaghefi2022deep, xiang_dare_2023, morio2023an, nicolas_webersinke_climatebert_2021, Wang2021EvidenceBA, spokoyny2023answering, lai_using_2023, luo_detecting_2020 and classical approaches spokoyny2023answering, morio2023an, luo_detecting_2020.
+vaid-etal-2022-towards also experimented with FastText and xiang_dare_2023 with LSTM-based models .
+The performances are quite heterogeneous, lai_using_2023 reaching F1-scores around 90% on classification of stance on remediation efforts, luo_detecting_2020 around 72% on GWSD, while performance on ClimateStance could not exceed 60%.
+Performances on ClimateFEVER are also quite low, however, when selecting only non-ambiguous examples, xiang_dare_2023 could reach performances around 80%.
+Finally, performances on LobbyMap [morio2023an] are quite low (between 31% and 57.3% depending on the strictness of the metric).
+Overall, the performances show that the datasets are challenging.666see Table 20 in the Appendix for details.
 Exhaustivity
-Datasets such as ClimateFEVER (Leippold and Diggelmann, 2020) and LobbyMap (Morio and Manning, 2023) rely on automated construction methods that may not ensure exhaustivity of evidence coverage.
+Datasets such as ClimateFEVER [diggelmann_climate-fever_2020] and LobbyMap [morio2023an] rely on automated construction methods that may not ensure exhaustivity of evidence coverage.
 ClimateFEVER utilizes BM25 and Wikipedia for evidence retrieval, which could result in missing relevant information, particularly when compared to more advanced retrieval methods.
@@
 While existing datasets provide strong foundations, they may suffer from a lack of exhaustivity, particularly in evidence retrieval and coverage, which requires further investigation.
-4.8. Question Answering
-Question answering (QA) is a known task:
+4.8 Question Answering
 Task Description
 Given an input question and a set of resources (paragraphs or documents), produce an answer to the question.
+Table 10: Summary of datasets, labels, scopes, and sources for climate-related QA datasets.
 Dataset Input Labels
-ClimaQA  (Spokoyny et al., 2023) The text from a response to one of the CDP questions; and one of the questions from the CDP questionnaire 1: the response answers this question 0: The response does not answer this question, but another one
-ClimateQA (Luccioni et al., 2020) Sentence from reports and a question based on the TCFD recommendations 1: the sentence answers the question 0: The sentences does not answer the question
-ClimaINS  (Spokoyny et al., 2023) The text from a response to one of the questions from the NAIC questionnaire MANAGE, RISK PLAN, MITIGATE, ENGAGE, ASSESS, RISKS The labels correspond the 8 questions asked in the NAIC questionnaires
-Table 8. Summary of datasets, labels, scopes, and sources for climate-related QA datasets.
-QA can be used for climate-specific applications such as structuring the information related to climate change from documents (Luccioni et al., 2020; Schimanski et al., 2023a), building chatbots with climate-related knowledge to make it more accessible to non-experts (Vaghefi et al., 2023; Rony et al., 2022), or helping identify potentially misleading information (Ni et al., 2023).
+ClimaQA  [spokoyny2023answering] The text from a response to one of the CDP questions; and one of the questions from the CDP questionnaire 1: the response answers this question; 0: The response does not answer this question, but another one
+ClimateQA [luccioni_analyzing_2020] Sentence from reports and a question based on the TCFD recommendations 1: the sentence answers the question; 0: The sentences does not answer the question
+ClimaINS  [spokoyny2023answering] The text from a response to one of the questions from the NAIC questionnaire MANAGE, RISK PLAN, MITIGATE, ENGAGE, ASSESS, RISKS The labels correspond the 8 questions asked in the NAIC questionnaires
+QA can be used for climate-specific applications such as structuring the information related to climate change from documents [luccioni_analyzing_2020, tobias_schimanski_climatebert-netzero_2023], building chatbots with climate-related knowledge to make it more accessible to non-experts [s_vaghefi_chatclimate_2023, cliamtebot_2022], or helping identify potentially misleading information [jingwei_ni_paradigm_2023].
 Solution
-The QA process can be divided into two main steps (Kraus et al., 2023): the retrieval step and the answer generation step.
+The QA process can be divided into two main steps [krausEnhancingLargeLanguage2023]: the retrieval step and the answer generation step.
 The retrieval step involves locating the relevant information or answer to the user’s query in external documents.
@@
 Retriever systems can be generic, using techniques such as sentence similarity, BM25, or document tags to filter documents and retrieve specific passages.
-Schimanski et al.
-(2024b) published ClimRetrieve, a benchmark for climate-related information retrieval in corporate reports.
+schimanski-etal-2024-climretrieve published ClimRetrieve, a benchmark for climate-related information retrieval in corporate reports.
 They found that simple embedding approaches are limited.
-Therefore, several works have specifically focused on improving the identification of answers within climate-related documents (Luccioni et al., 2020; Spokoyny et al., 2023).
-Luccioni et al.
-(2020) reformulated TCFD recommendations as questions and annotated reports to identify answers to the questions.
+Therefore, several works have specifically focused on improving the identification of answers within climate-related documents [luccioni_analyzing_2020, spokoyny2023answering].
+luccioni_analyzing_2020 reformulated TCFD recommendations as questions and annotated reports to identify answers to the questions.
 The dataset includes questions paired with potential answers, each labeled to indicate whether the answer addresses the question or not.
 Based on the QA model trained on this dataset, they also proposed a tool called ClimateQA.
-Spokoyny et al.
-(2023) focused on questions from the CDP questionnaire and the NAIC Climate Risk Disclosure survey.
+spokoyny2023answering focused on questions from the CDP questionnaire and the NAIC Climate Risk Disclosure survey.
 As the responses to those questionnaires are publicly available, they constructed two datasets, ClimaINS and ClimaQA.
@@
 As ClimaQA is framed as a retrieval task, the authors used BM25, sentence-BERT, and ClimateBERT to rank possible answers.
-Using the methodology evaluated on ClimaQA, they proposed a system to extract responses to CDP questions directly from unstructured reports and automatically fill the questionnaire (Spokoyny et al., 2023).
-Luccioni et al.
-(2020) trained and evaluated a RoBERTa-based model called ClimateQA on finding answers to TCFD-based questions.
-They all reached good performance but with room for improvement (see Table 21 in the Appendix for details), demonstrating the feasibility of the task but also the need for further research.
+Using the methodology evaluated on ClimaQA, they proposed a system to extract responses to CDP questions directly from unstructured reports and automatically fill the questionnaire [spokoyny2023answering].
+luccioni_analyzing_2020 trained and evaluated a RoBERTa-based model called ClimateQA on finding answers to TCFD-based questions.
+They all reached good performance but with room for improvement,  [^7] See Table 21 in the Appendix for details.
+demonstrating the feasibility of the task but also the need for further research.
 Answer generation
@@
 The generation can be simple such as finding a particular information in a paragraph.
-For example, Schimanski et al.
-(2023a) used a QA model (RoBERTa SQuaD v2 (Rajpurkar et al., 2016)) to extract the target year, the percentage of reduction, and the baseline year from the reduction target of a company from an input text that contains an emission reduction target.
-While it is possible to rely on generalist models, as they are proficient few-shot learners (Brown et al., 2020; Thulke et al., 2024), multiple studies proposed using domain-specific models.
-Climate Bot (Rony et al., 2022) built a dataset of climate-specific QA.
+For example, tobias_schimanski_climatebert-netzero_2023 used a QA model (RoBERTa SQuaD v2 [rajpurkar-etal-2016-squad]) to extract the target year, the percentage of reduction, and the baseline year from the reduction target of a company from an input text that contains an emission reduction target.
+While it is possible to rely on generalist models, as they are proficient few-shot learners [lm_few_shot_learner, thulke2024climategpt], multiple studies proposed using domain-specific models.
+Climate Bot [cliamtebot_2022] built a dataset of climate-specific QA.
 Given a question and a document (scientific/news), the model should find the span answering the question.
-Mullappilly et al.
-(2023) proposed a dataset of question-answer pairs based on ClimaBench (Spokoyny et al., 2023) and CCMRC (Rony et al., 2022) to train a model to generate answers.
-Thulke et al.
-(2024) introduced a climate-specific corpus of prompt/completion pairs for Instruction Fine-Tuning created by experts and non-experts.
-(Rony et al., 2022) finetuned an ALBERT model on their climate-specific QA dataset.
-Mullappilly et al.
-(2023) trained a Vicuna-7B on their climate-specific answer generation dataset.
-And Thulke et al.
-(2024) trained LLama-2-based models (ClimateGPT) on their domain-specific prompt/completion dataset.
+mullappilly-etal-2023-arabic proposed a dataset of question-answer pairs based on ClimaBench [spokoyny2023answering] and CCMRC [cliamtebot_2022] to train a model to generate answers.
+thulke2024climategpt introduced a climate-specific corpus of prompt/completion pairs for Instruction Fine-Tuning created by experts and non-experts.
+cliamtebot_2022 finetuned an ALBERT model on their climate-specific QA dataset.
+mullappilly-etal-2023-arabic trained a Vicuna-7B on their climate-specific answer generation dataset.
+And thulke2024climategpt trained LLama-2-based models (ClimateGPT) on their domain-specific prompt/completion dataset.
 They also evaluated their ClimateGPT model alongside multiple LLMs on climate-related benchmarks.
-They found that the domain-specific models outperformed generalist ones (see Table 21 in the Appendix for details).
+[^8] They found that the domain-specific models outperformed generalist ones (see Table 21 in the Appendix for details).
 Retrieval Augmented Generation
 By combining the last two components (a retriever and a QA system), one can develop Retrieval-Augmented Generation (RAG) systems.
-Rony et al.
-(2022) provide the first example of a RAG system in climate-related tasks based on sentence-BERT to retrieve documents and AlBERT to identify the answer.
-More recently, Ni et al.
-(2023) proposed ChatReport, a methodology based on ChatGPT for analyzing corporate reports through the lens of TCFD questions using RAG.
+cliamtebot_2022 provide the first example of a RAG system in climate-related tasks based on sentence-BERT to retrieve documents and AlBERT to identify the answer.
+More recently, jingwei_ni_paradigm_2023 proposed ChatReport, a methodology based on ChatGPT for analyzing corporate reports through the lens of TCFD questions using RAG.
 The reports, the answers to the TCFD questions, and a TCFD conformity assessment are chunked and stored in a vector store for easy retrieval.
@@
 They added in the prompt the notions of greenwashing and cheap talk to invite the model to provide a critical analysis of the retrieved answers.
-Similarly, Vaghefi et al.
-(2023) introduced ChatClimate, a RAG-based pipeline to augment GPT-4 with knowledge about IPCC reports.
-While the previous LLM-based approaches rely on closed-source models, Mullappilly et al.
-(2023) and Thulke et al.
-(2024) proposed RAG-based pipelines relying on open-source models.
-Rony et al.
-(2022) is the only study that evaluated their RAG pipeline using classical metrics (F1-score, BLEU, METEOR).
-Unfortunately, those metrics are not well aligned with human judgment for text generation (Chen et al., 2019).
-Therefore Vaghefi et al.
-(2023) and Mullappilly et al.
-(2023) relied on human or ChatGPT evaluations.
+Similarly, s_vaghefi_chatclimate_2023 introduced ChatClimate, a RAG-based pipeline to augment GPT-4 with knowledge about IPCC reports.
+While the previous LLM-based approaches rely on closed-source models, mullappilly-etal-2023-arabic and thulke2024climategpt proposed RAG-based pipelines relying on open-source models.
+cliamtebot_2022 is the only study that evaluated their RAG pipeline using classical metrics (F1-score, BLEU, METEOR).
+Unfortunately, those metrics are not well aligned with human judgment for text generation chen-etal-2019-evaluating.
+Therefore s_vaghefi_chatclimate_2023 and mullappilly-etal-2023-arabic relied on human or ChatGPT evaluations.
 They conclude that their approaches improve on non-domain-specific RAG systems.
@@
 Modern LLMs achieve strong results in QA due to instruction-following fine-tuning, which improves adherence to specific instructions, and Retrieval-Augmented Generation (RAG), which bases its answers on retrieved documents.
-However, retrieval systems remain a performance bottleneck (Maekawa et al., 2024).
-Furthermore, answer generation is challenging to evaluate (Sai et al., 2022), often requiring human feedback or advanced model-based assessments.
-4.9. Classification of Deceptive Techniques
+However, retrieval systems remain a performance bottleneck [maekawa-etal-2024-retrieval].
+Furthermore, answer generation is challenging to evaluate [survey_nlg_eval], often requiring human feedback or advanced model-based assessments.
+4.9 Classification of Deceptive Techniques
 In analyzing climate-related discourse, it can be useful to identify rhetorical strategies that might obscure or misrepresent an entity’s stance.
-For example, expressing support for climate remediation policies on one side (Morio and Manning, 2023), and promoting arguments downplaying the urgency of climate change on the other (Coan et al., 2021) could signal a lack of authenticity.
-Another example could be promoting a product through misleading rhetoric, by, for example, claiming that a product is better because it is natural, which would be an “appeal to nature” – a fallacious argument (Vaid et al., 2022; Jain et al., 2023).
+For example, expressing support for climate remediation policies on one side morio2023an, and promoting arguments downplaying the urgency of climate change on the other coanComputerassistedClassificationContrarian2021 could signal a lack of authenticity.
+Another example could be promoting a product through misleading rhetoric, by, for example, claiming that a product is better because it is natural, which would be an “appeal to nature” – a fallacious argument vaid-etal-2022-towards, jain_supply_2023.
 Detecting these deceptive techniques could help identify misleading communications.
+Table 11: Summary of datasets, labels, scopes, and sources for tasks related to deceptive techniques in climate-related context.
 Dataset Input Labels (Labels details)
-LogicClimate (Jin et al., 2022) texts from climatefeedback.org Faulty Generalization, Ad Hominem, Ad Populum, False Causality, … Classifies fallacies (Multi-label)
-Neutralization Techniques (Bhatia et al., 2021) paragraphs from other previous works on climate-change Denial of Responsibility, Denial of Injury, Denial of Victim, Condemnation of the Condemner, … Classifies neutralization techniques
-Contrarian Claims (Coan et al., 2021) paragraphs from conservative think tank No Claim, Global Warming is not happening, Climate Solutions won’t work, Climate impacts are not bad, … Classifies arguments into super/sub-categories of climate science denier’s arguments
-Table 9. Summary of datasets, labels, scopes, and sources for tasks related to deceptive techniques in climate-related context.
+LogicClimate [jin-etal-2022-logical] texts from climatefeedback.org Faulty Generalization, Ad Hominem, Ad Populum, False Causality, … Classifies fallacies (Multi-label)
+Neutralization Techniques [bhatia_automatic_2021-1] paragraphs from other previous works on climate-change Denial of Responsibility, Denial of Injury, Denial of Victim, Condemnation of the Condemner, … Classifies neutralization techniques
+Contrarian Claims [coanComputerassistedClassificationContrarian2021] paragraphs from conservative think tank No Claim, Global Warming is not happening, Climate Solutions won’t work, Climate impacts are not bad, … Classifies arguments into super/ sub-categories of climate science denier arguments
 Task Description
 The goal is to classify statements into argumentative categories: fallacies, types of arguments, or rhetorical techniques.
-Unsupervised Solutions
-Oppong-Tawiah and Webster (2023) frame greenwashing as fake news.
+Annotation-free
+divinus_oppong-tawiah_corporate_2023 frame greenwashing as fake news.
 They propose to tackle this identification through the form perspective.
@@
 “Solutions will not work”, or “Human influence is not demonstrated”).
-Therefore, Coan et al.
-(2021) constructed a taxonomy of such contrarian claims and published a large dataset annotated with such claims.
-Similarly, Bhatia et al.
-(2021) proposed a dataset to classify neutralization techniques, i.e., rationalizations that individuals use to justify deviant or unethical behavior (e.g.
+Therefore, coanComputerassistedClassificationContrarian2021 constructed a taxonomy of such contrarian claims and published a large dataset annotated with such claims.
+Similarly, bhatia_automatic_2021-1 proposed a dataset to classify neutralization techniques, i.e., rationalizations that individuals use to justify deviant or unethical behavior (e.g.
 Denial of the victim).
-Finally, Jin et al.
-(2022) studied fallacious arguments and how they apply to climate change more broadly.
+Finally, jin-etal-2022-logical studied fallacious arguments and how they apply to climate change more broadly.
 They constructed LogicClimate, a climate-specific dataset of sentences annotated with fallacies, using articles from the Climate Feedback website.
-Supervised Solutions
-Fine-tuned Transformer architectures and classical approaches have been evaluated on each dataset (Jin et al., 2022; Bhatia et al., 2021; Coan et al., 2021).
-Jin et al.
-(2022) experimented with an ELECTRA model, fine-tuned to use the structure of the argument.
+Models
+Fine-tuned Transformer architectures and classical approaches have been evaluated on each dataset [jin-etal-2022-logical, bhatia_automatic_2021-1, coanComputerassistedClassificationContrarian2021].
+jin-etal-2022-logical experimented with an ELECTRA model, fine-tuned to use the structure of the argument.
 The performance of fine-tuned approaches ranges from 58.77% to 79%.
 The models struggle with fallacy detection in LogicClimate, but perform well with generalist fallacy detection, contrarian claims, and neutralization techniques.
-Thulke et al.
-(2024) experimented with multiple zero-shot approaches using multiple LLM on the binary classification of contrarian claims using Coan et al.
-(2021)’s dataset.
-Their climateGPT-70B and LLama-2-Chat-70B models both reached an F1-score of 72.5% (see Table 23 in the Appendix for details).
+thulke2024climategpt experimented with multiple zero-shot approaches using multiple LLM on the binary classification of contrarian claims using coanComputerassistedClassificationContrarian2021’s dataset [^9] Their climateGPT-70B and LLama-2-Chat-70B models both reached an F1-score of 72.5% (see Table in the Appendix for details)..
 Insights
 The low performance of models on tasks such as fallacy detection and neutralization classification indicates the inherent complexity of these tasks.
-Fallacy detection, for one, is known to be inherently subjective (Helwe et al., 2024).
+Fallacy detection, for one, is known to be inherently subjective [helwe-etal-2024-mafalda].
 The detection of neutralization techniques, too, appears to be subjective, as even human annotators achieve only a moderate performance level of 70% F1-score.
-Furthermore, both LogicClimate (Jin et al., 2022) and the neutralization dataset (Bhatia et al., 2021) are very small in size.
-To address these limitations, future research could explore several directions, including increasing the size of the datasets, defining the labels more precisely, or permitting multiple correct annotations (Helwe et al., 2024) to acknowledge viable disagreement.
+Furthermore, both LogicClimate [jin-etal-2022-logical] and the neutralization dataset [bhatia_automatic_2021-1] are very small in size.
+To address these limitations, future research could explore several directions, including increasing the size of the datasets, defining the labels more precisely, or permitting multiple correct annotations [helwe-etal-2024-mafalda] to acknowledge viable disagreement.
 Additionally, the analyses on rhetorical techniques could be combined with other analyses (such as stance detection) to search for communication patterns.
-4.10. Environmental Performance Prediction
+4.10 Environmental Performance Prediction
 Greenwashing can be interpreted as a misrepresentation of the company or product’s environmental performance.
@@
 In variants of this task, the goal is to predict not a value, but a range for that value, out of a set of possible ranges.
-Unsupervised Solutions
-Jain et al.
-(2023) proposed to predict the Scope 3 emissions from the list of financial transactions of the company.
+Annotation-free
+jain_supply_2023 proposed to predict the Scope 3 emissions from the list of financial transactions of the company.
 They first performed classification of all transaction descriptions to map them to their emission factors (emission per \$ spent), and then compute the emissions of each transaction.
-For ESG score prediction, Bronzini et al.
-(2023) proposes to use LLMs as few-shot learners to extract triples of ESG category, action, and company from sustainable reports, and to construct a graph representation with few-shot learning.
+For ESG score prediction, bronzini_glitter_2023 proposes to use LLMs as few-shot learners to extract triples of ESG category, action, and company from sustainable reports, and to construct a graph representation with few-shot learning.
 They demonstrated the triplet generation using Alpaca, WizardLM, Flan-T5, and ChatGPT on a few examples.
@@
 Annotated datasets
-Mehra et al.
-(2022) proposed to predict the changes in the ESG score instead of the actual value.
+Mehra_2022 proposed to predict the changes in the ESG score instead of the actual value.
 They constructed their dataset by extracting the three sentences most relevant to the environment from financial reports and associated them with the E score [^11] ESG scores are usually aggregated scores along multiple dimensions.
 In this study, they are focusing on the ”Environment” part of the ESG score.’s change and direction of change.
-Instead of relying on scores, Clarkson et al.
-(2020) proposed to focus on good/bad CSR (corporate social responsibility) performers.
+Instead of relying on scores, clarkson_nlp_us_csr proposed to focus on good/bad CSR (corporate social responsibility) performers.
 Their approach evaluates CSR performance based on linguistic style rather than content, aiming to identify whether language patterns alone influence the perception of CSR quality.
-Sharma et al.
-(2024) introduced a multi-modal dataset of the tweets of companies to predict a company’s ESG unmanaged risk.
-Finally, focusing on products and not companies, Lin et al.
-(2023b) introduced a dataset of online product descriptions and reviews used to predict the Finch Score.
+Greenscreen introduced a multi-modal dataset of the tweets of companies to predict a company’s ESG unmanaged risk.
+Finally, focusing on products and not companies, linSUSTAINABLESIGNALSijcai2023 introduced a dataset of online product descriptions and reviews used to predict the Finch Score.
 All these scores can be used to identify companies and products that are actually environmentally friendly, helping users distinguish actual sustainability from greenwashing.
@@
 ESG score or quantity of emissions), which might indicate cases of greenwashing.
-Supervised Solutions
-Mehra et al.
-(2022) evaluated fine-tuned Transformers on ESG change prediction.
-Clarkson et al.
-(2020) experimented with hand-chosen features plugged into random forest and SVM on CSR performers prediction.
-Lin et al.
-(2023b) also experimented with traditional approaches (e.g.
+Models
+Mehra_2022 evaluated fine-tuned Transformers on ESG change prediction.
+clarkson_nlp_us_csr experimented with hand-chosen features plugged into random forest and SVM on CSR performers prediction.
+linSUSTAINABLESIGNALSijcai2023 also experimented with traditional approaches (e.g.
 gradient boosting) and custom architectures based on Transformer models on Finch score prediction.
-Sharma et al.
-(2024) introduced baseline models, which are simple image and text embedding models (e.g., CLIP or sentence-BERT) with a classification head on ESG score prediction.
-A detailed list of models is reported in Table LABEL:tab:models in the Appendix.
-Bronzini et al.
-(2023); Clarkson et al.
-(2020); Sharma et al.
-(2024) did not report baselines, which makes the study difficult to analyze.
-(Mehra et al., 2022) reported good accuracy, demonstrating that a few sentences hold a significant amount of information to predict ESG score changes.
+Greenscreen introduced baseline models, which are simple image and text embedding models (e.g., CLIP or sentence-BERT) with a classification head on ESG score prediction.
+bronzini_glitter_2023, clarkson_nlp_us_csr, Greenscreen did not report baselines, which makes the study difficult to analyze.
+Mehra_2022 reported good accuracy, demonstrating that a few sentences hold a significant amount of information to predict ESG score changes.
 However, there is definitely room for improvement, indicating that the information selected is not sufficient.
-Lin et al.
-(2023a) reported the performance of the baseline (average score) reaching already mean squared error (MSE) of 11.7% already quite low.
-Their Transformer-based approach reached MSE of 7.4% improving slightly on classical approaches (Gradient Boosting reaching MSE of 8.2%).
-Detailed performances can be found in Table 24 in the Appendix.
+lin-etal-2023-linear reported the performance of the baseline (average score) reaching already mean squared error (MSE) of 11.7% already quite low.
+Their Transformer-based approach reached MSE of 7.4% improving slightly on classical approaches (Gradient Boosting reaching MSE of 8.2%) [^12] Detailed performances can be found in Table 24 in the Appendix..
 Insights
 We expected the prediction of ESG and CSR metrics to be difficult and require an extensive understanding of a company.
-However, Bronzini et al.
-(2023); Lin et al.
-(2023b); Mehra et al.
-(2022) and Clarkson et al.
-(2020) showed that it is possible to build strong predictors relying only on textual elements.
-This can be explained because analysts reward transparency (Bronzini et al., 2023), so the quantity and complexity of disclosure have strong predictive power (Clarkson et al., 2020).
-Although existing studies have examined specific dimensions, future research should investigate the interaction between form (Clarkson et al., 2020) and content (Bronzini et al., 2023), as this could help uncover inconsistencies that may indicate greenwashing practices.
-5. Greenwashing Detection
+However, bronzini_glitter_2023, linSUSTAINABLESIGNALSijcai2023, Mehra_2022 and clarkson_nlp_us_csr showed that it is possible to build strong predictors relying only on textual elements.
+This can be explained because analysts reward transparency [bronzini_glitter_2023], so the quantity and complexity of disclosure have strong predictive power [clarkson_nlp_us_csr].
+Although existing studies have examined specific dimensions, future research should investigate the interaction between form [clarkson_nlp_us_csr] and content [bronzini_glitter_2023], as this could help uncover inconsistencies that may indicate greenwashing practices.
+5 Greenwashing Detection
 After having discussed different subtasks of greenwashing detection, we now come to the final, all-englobing task:
@@
 Quantifying climate-related disclosure can highlight greenhushing and selective disclosure.
-Bingler et al.
-(2022) identify a lack of disclosure in the Strategy and Metrics&Targets categories, which they describe as cherry-picking.
+bingler_cheap_2021 identify a lack of disclosure in the Strategy and Metrics&Targets categories, which they describe as cherry-picking.
 They also highlight that merely announcing support for the TCFD does not lead to an increase in disclosure – on the contrary, it is a practice called cheap talk.
-Similarly, Auzepy et al.
-(2023) conducted a more fine-grained analysis of TCFD-related disclosure in the banking industry.
+Similarly, auzepy_evaluating_2023 conducted a more fine-grained analysis of TCFD-related disclosure in the banking industry.
 They also highlighted large differences across TCFD categories.
@@
 Climate Communication as an Image-Building Strategy
-Ding et al.
-(2023) and Chou et al.
-(2023) found a correlation between climate-related disclosure and carbon emission: companies that emit more tend to disclose more climate-related information.
-Bingler et al.
-(2023) proposed a Cheap talk index based on claims specificity.
+dingCarbonEmissionsTCFD2023 and chou_ESG found a correlation between climate-related disclosure and carbon emission: companies that emit more tend to disclose more climate-related information.
+bingler2023cheaptalkspecificitysentiment proposed a Cheap talk index based on claims specificity.
 They reached the same conclusion that larger emitters tend to disclose more, but also that negative news coverage is correlated with cheap talk.
-Polignano et al.
-(2022) showed that while disclosing more, reports mostly focus on positive disclosure.
-Kang and Kim (2022) proposed a sentiment ratio metric that highlights overly positive corporate reports.
+marco_polignano_nlp_2022 showed that while disclosing more, reports mostly focus on positive disclosure.
+hyewon_kang_analyzing_2022 proposed a sentiment ratio metric that highlights overly positive corporate reports.
 They identified periods of overly positive communications that followed negative environmental controversies; in other words, companies that tried to rebuild their image after a controversy.
-Gorovaia and Makrominas (2025) analyzed CSR reports of environmental violators and companies with clean records.
+csr_report_greenwashing analyzed CSR reports of environmental violators and companies with clean records.
 They studied an ensemble of variables: quantification of environmental content, readability score, and sentiment analysis.
@@
 This suggests that companies use climate disclosure as a tool to mitigate controversies and/or to improve their image.
-Zhao et al.
-(2023) propose to use the discrepancies between internal disclosure and social media perception of a company to identify potential greenwashing.
+kdir23 propose to use the discrepancies between internal disclosure and social media perception of a company to identify potential greenwashing.
 The Role of Disclosure Style in Perceived Commitment
-Clarkson et al.
-(2020) found that companies that use a more complex language in their CSR disclosure are associated with a higher CSR rating.
-Schimanski et al.
-(2023b) concluded that a higher quantity of ESG communication is associated with higher ESG rating.
-Rouen et al.
-(2023) also highlighted the relationship between disclosure quantity and complexity with the ESG score.
+clarkson_nlp_us_csr found that companies that use a more complex language in their CSR disclosure are associated with a higher CSR rating.
+schimanski_bridging_2023 concluded that a higher quantity of ESG communication is associated with higher ESG rating.
+rouenEvolutionESGReports2023 also highlighted the relationship between disclosure quantity and complexity with the ESG score.
 This might indicate that the linguistic style is a good predictor of the substance of the discourse, or that analysts are rewarding companies that communicate on ESG-related issues.
 ESG Score in Greenwashing Detection
-Lee and Raschke (2023) proposed a greenwashing index based on the difference between the ESG score and the ESG score, weighted by the quantity of communication on each topic (E, S and G).
-Sharma et al.
-(2024) define two types of risks: managed risk (the company is addressing it), and unmanaged risk (the company is not currently addressing it but it could).
+LEE_greenwashing proposed a greenwashing index based on the difference between the ESG score and the ESG score, weighted by the quantity of communication on each topic (E, S and G).
+Greenscreen define two types of risks: managed risk (the company is addressing it), and unmanaged risk (the company is not currently addressing it but it could).
 Based on these, they propose an ESG unmanaged risk score as a measure for environmental performance, which can then be used as a signal for identifying greenwashing.
 Contrasting Stances: A Method to Identify Greenwashing in Climate Communications
-Morio and Manning (2023) proposed to evaluate a company’s stance towards climate change mitigation policies.
+morio2023an proposed to evaluate a company’s stance towards climate change mitigation policies.
 They hypothesize that companies with a mixed stance might be engaging in greenwashing.
 The hypothesis was not tested and remains unconfirmed; still, such methodology helps understand the narratives of companies around climate change.
-Coan et al.
-(2021) were able to analyze the type of contrarian claims used by conservative think-tank (CTTs) websites and contrarian blogs.
+coanComputerassistedClassificationContrarian2021 were able to analyze the type of contrarian claims used by conservative think-tank (CTTs) websites and contrarian blogs.
 They showed a shift over time from climate change denying arguments to opposing climate change mitigation policies.
 Despite this shift, they found that the primary donors of these think tanks continue to support organizations promoting a climate-denier narrative aimed at discrediting scientific evidence and scientists.
-Discrepancies between the stance extracted from official communications (Morio and Manning, 2023) on the one hand, and the stance of third-party media organizations such as think tanks founded by the company (Coan et al., 2021) on the other, might indicate greenwashing.
+Discrepancies between the stance extracted from official communications [morio2023an] on the one hand, and the stance of third-party media organizations such as think tanks founded by the company [coanComputerassistedClassificationContrarian2021] on the other, might indicate greenwashing.
 Defining the linguistic features of greenwashing
-Oppong-Tawiah and Webster (2023) we propose to identify greenwashing using the linguistic profile of tweets.
+divinus_oppong-tawiah_corporate_2023 we propose to identify greenwashing using the linguistic profile of tweets.
 They estimate the deceptiveness of the text using a keyword-based approach and linguistic indicators (e.g.
@@
 Insights
-Several indicators for greenwashing have been proposed: overly positive sentiment (Kang and Kim, 2022), the use of specific linguistic cues (Oppong-Tawiah and Webster, 2023), discrepancies between ESG scores and corporate disclosures (Lee and Raschke, 2023; Sharma et al., 2024), ambiguous stances (Morio and Manning, 2023), and inconsistencies between social media perception and official disclosures (Zhao et al., 2023).
-While these approaches laid the theoretical groundwork for understanding indicators of greenwashing, they have a significant limitation: they are not empirically evaluated (Bernini et al., 2023).
-Only a few studies, such as that by Kang and Kim (2022), have made attempts to validate their signals against actual cases, but even these efforts have not resulted in a comprehensive dataset of greenwashing examples.
+Several indicators for greenwashing have been proposed: overly positive sentiment [hyewon_kang_analyzing_2022], the use of specific linguistic cues [divinus_oppong-tawiah_corporate_2023], discrepancies between ESG scores and corporate disclosures [LEE_greenwashing, Greenscreen], ambiguous stances [morio2023an], and inconsistencies between social media perception and official disclosures [kdir23].
+While these approaches laid the theoretical groundwork for understanding indicators of greenwashing, they have a significant limitation: they are not empirically evaluated [measuring_greenwashing].
+Only a few studies, such as that by hyewon_kang_analyzing_2022, have made attempts to validate their signals against actual cases, but even these efforts have not resulted in a comprehensive dataset of greenwashing examples.
 This gap between theory and practice highlights the necessity of developing datasets containing real-world instances of greenwashing.
@@
 Finally, the publication of any such dataset exposes the authors to charges of defamation by the companies they accuse of greenwashing.
-6. Open Challenges in Greenwashing Detection
-We discuss open challenges in the evaluation (Section 6.1), model robustness (Section 6.2), and data sources (Section 6.3).
-6.1. Evaluation Methodology
+6 Open Challenges in Greenwashing Detection
+While the lack of real-world greenwashing datasets constitutes a major obstacle, it is only one of several unresolved challenges.
+Building on the reviewed literature, this section examines additional limitations that affect greenwashing detection and, more broadly, NLP tasks characterized by ambiguity, subjectivity, and real-world constraints.
+Although discussed through the lens of greenwashing-related applications, these challenges are not specific to this domain and reflect broader methodological issues in applied NLP.
+We focus on three interconnected dimensions: (6.1) evaluation methodology, where ambiguity and task complexity complicate performance assessment; (6.2) model robustness, including sensitivity to noise and data imbalance; and (6.3) data sources, encompassing reproducibility, dataset availability, and definitional issues.
+6.1 Evaluation Methodology
+Firstly, as models improve, they are able to tackle more complex tasks, moving from classification to text generation, and from extraction to reasoning.
+This progression in task complexity creates an increase in evaluation complexity.
+Classification Evaluation
+Many tasks in greenwashing detection are framed as classification tasks.
+Such a framing may be too simplistic, in that it is blind to the semantic similarity between the classes.
+For example, in the Contrarian Claims dataset  [coanComputerassistedClassificationContrarian2021], models may fail on the fine-grained classification, but still get the super-category right.
+For such cases, morio2023an proposed strict and relaxed metrics, where the relaxed metric grouped similar labels together (e.g.
+”strongly supporting” and ”supporting” grouped in ”supporting” label).
 Ambiguous Labels
-Many tasks in greenwashing detection are framed as classification tasks.
-However, in some cases, the labels are not sufficiently distinguishable.
-For example, specificity classification (Bingler et al., 2023) is the task of distinguishing specific claims from unspecific claims.
-It turns out that this task is itself too unspecific, as the low inter-annotator agreement shows.
-However, even for tasks that seem easy to distinguish, some edge cases remain ambiguous, e.g., the distinction between “climate-related” and “not climate-related”.
-The climate detection model climateBERT (Bingler et al., 2023) classifies the sentence “This year, we are producing more than 10,000 [bicycle/electric bicycle], demonstrating our continued growth and commitment to meeting market demand.” as “climate-related” for bicycle but not for electric bicycle.
-While electric bicycles are often associated with solutions to reduce emissions, the statement is not explicitly making the link.
-If we consider indirect links, then non-electric bicycles are also solutions.
-There lies the question of which product is inherently linked to climate-change, and which one requires contextualization.
-When labels in classification tasks are not clearly distinguishable, the task is not well-defined.
-Consequently, it does not mean much if models perform well on such a task.
-It merely means that the model has learned whatever label distribution the dataset had.
-Work in this area should thus always first check human inter-annotator agreement on the task to make sure it is meaningful.
+One particular issue is that the main hypothesis of the classification task is that categories are distinguishable from each other.
+This is not always the case.
+For example, specificity classification [bingler2023cheaptalkspecificitysentiment] is the task of distinguishing specific claims from unspecific claims.
+It turns out that this task is itself too unspecific, as the low IAA shows.
+While this issue appears obvious for this particularly ambiguous task, other less ambiguous classes also contain an ambiguous frontier.
+For example, climate/non-climate labels can be debated in some instance: ”The company produced 10 million computers this year” or ”The company produced 10 million solar panels this year” are respectively predicted ”non-climate” and ”climate” by climateBERT’s climate detection model [bingler2023cheaptalkspecificitysentiment].
+In this example, the sentences are fundamentally similar.
+If the company’s business model is to produce solar panels, then this statement is not necessarily about climate, at least not more than the other one.
 Generated Text Evaluation
 Some tasks are not classification tasks, but text generation tasks.
-Evaluating generated text is a challenging task (Sai et al., 2022), and different techniques are being used: BLEU, METEOR, human-based evaluation, and LLM-based evaluation (Mullappilly et al., 2023; Rony et al., 2022; Vaghefi et al., 2023; Thulke et al., 2024).
-BLEU and METEOR are not well-suited for evaluating tasks such as question answering  (Chen et al., 2019), as they rely only on n-gram similarity.
-Human evaluation also has its shortcomings.
-First, it is limited to smaller sizes and variations in settings.
-Second, human evaluation of generated texts should be based on explicit guidelines (Thulke et al., 2024; Bulian et al., 2024), but sometimes resorts to comparisons (Mullappilly et al., 2023; Thulke et al., 2024) or simple user preferences (Mullappilly et al., 2023).
-Third, human evaluation can be influenced by factors other than the quality of the answer – for example, the observer-expectancy effect (Rosenthal, 1976).
-For instance, in ChatClimate (Vaghefi et al., 2023), humans evaluate the accuracy of text generated by a generic ChatGPT vs.
+Evaluating generated text is a challenging task  [survey_nlg_eval].
+We argue that evaluating generated text is still a challenge deserving future research.
+In the papers we considered for this review, to evaluate generative models, we came across solutions such as using BLEU, METEOR, human-based evaluation, and LLM-based evaluation  [mullappilly-etal-2023-arabic, cliamtebot_2022, s_vaghefi_chatclimate_2023, thulke2024climategpt].
+BLEU and METEOR are not great to evaluate tasks such as question answering [chen-etal-2019-evaluating].
+Human evaluation also has many shortcomings.
+Firstly, it is not possible to perform an automatic evaluation; the evaluations are therefore limited to smaller sizes and variations in settings.
+Secondly, human evaluation of generated texts often relies on comparisons [mullappilly-etal-2023-arabic, thulke2024climategpt].
+While thulke2024climategpt refers to specific comparison guidelines [bulian2024assessinglargelanguagemodels], [mullappilly-etal-2023-arabic] relies only on the evaluator’s preferences.
+Thirdly, human evaluation can be influenced by factors other than the quality of the answer – for example, the observer-expectancy effect [rosenthal1976].
+For example, in ChatClimate  [s_vaghefi_chatclimate_2023], humans evaluate the accuracy of text generated by a generic ChatGPT vs.
 the proposed ChatClimate chatbot.
 The text produced by the ChatClimate bot always gives specific references (down to page numbers), while the generic ChatGPT does not.
-This allows evaluators to easily identify the more precise answer based solely on the type of references used, raising the concern that this criterion may have disproportionately influenced assessments, potentially overlooking broader aspects such as overall accuracy or quality.
-LLM-based evaluation has its own challenges: Language models are known to reproduce human biases (Gallegos et al., 2024).
-They are sensitive to the way questions are formulated, to the order of answers in multiple-choice questions (Pezeshkpour and Hruschka, 2024), and to cues that can influence human evaluation such as using “Our Model”/“Their model” (which is the wording used in Mullappilly et al.
-(2023)).
+It is thus very easy for evaluators to identify the more accurate answer simply by looking at the type of references, so that there is the danger that this was in fact the only facet of the texts that was evaluated (and not, say, the general accuracy or quality of the answer).
+A third shortcoming of evaluations appears with LLM-based evaluation.
+Language models are known to reproduce human biases  [gallegos2024biasfairnesslargelanguage]: they are sensitive to the way questions are formulated, to the order of answers in multiple-choice questions  [pezeshkpour-hruschka-2024-large], and to cues that can influence human evaluation such as using “Our Model”/“Their model” (which is the wording used in mullappilly-etal-2023-arabic).
 LLM-based evaluation should thus always adhere to the same rigorous standards that are applied to human evaluation.
-Figure 1.
-Upset plot (Lex et al., 2014) of the methods used in the studies mentioned in this literature review.
-Each bar counts the number of studies reporting a given set of methods.
-The number of a papers using a given method is reported on the horizontal bars.
+Proper control of the evaluation setting is crucial.
+Subtle factors, such as the specific wording used or the unintentional leakage of task-related information, can bias the results.
+Therefore, strict protocols and a meticulously designed evaluation environment are essential to mitigate these risks.
+Figure 1: Upset plot [upset_plot] of the methods used in the studies mentioned in this literature review.
+Each bar counts the number of study reporting a given set of methods.
+The sum of a paper using a given method is reported on the horizontal bars.
 We have highlighted in red the intersection sets that include both fine-tuned Transformers and keyword-based methods.
 Baselines
-Figure 1 shows the approaches that are commonly used in greenwashing detection tasks (more details in Table LABEL:tab:models).
+Figure 1 shows the approaches that are commonly used in greenwashing detection tasks.
 We see that the most common solution is fine-tuning transformer models; more recent approaches include using LLMs as few-shot learners.
-The figure also shows which baselines have been used:
+However, when evaluating, the state-of-the-art models are compared to baselines.
+We argue that some basic baselines are often forgotten, even though they provide important insights:
   • Random/Majority Baseline: Any proposed approach should beat a random baseline.
 However, we find that only 9% of papers we studied compare their results to the random or majority baseline.
-For example, Friederich et al.
-(2021) reported performances under 50% for a binary classification task, which is worse than a theoretical random classifier on a balanced dataset.
+For example, Friederich_climate_risk_disclosure reported performances under 50% for a binary classification task, which is worse than a theoretical random classifier on a balanced dataset.
 Their dataset is not balanced, but it is difficult to draw conclusions without the performance of a random classifier.
-Random classifiers cannot just suggest that a model underperforms, but also that a model has a comparatively good performance.
-For example, on SciDCC or ClimaTOPIC, the performance of models is around 50-60%.
-While this looks mediocre, it is in fact an excellent performance as the random classifier performed around 5-7% (Spokoyny et al., 2023).
-In case of a heavily imbalanced dataset, the majority baseline is more meaningful.
-If such a baseline is not provided, but the dataset is available, it can even be computed a posteriori.
-Unfortunately, neither is the case for the work of Friederich et al.
-(2021).
-Similarly, a constant value baseline can help understand the performance of a regression model.
-Sharma et al.
-(2024) and Bronzini et al.
-(2023) evaluated the prediction of the ESG score, but do not provide such a baseline.
-Thus, the reader is left to wonder whether a model that always predicts the mean value might not actually perform better.
+Another interesting aspect of reporting the random classifier performance is to have a point of reference, or a lower baseline.
+For example, on SciDCC or ClimaTOPIC, the performance of models is around 50-60%, yet this is excellent performances as the random classifier performed around 5-7% [spokoyny2023answering].
+In case of an heavily imbalanced dataset, providing the performance of the majority baseline might be more meaningful.
+If the dataset is not avaiable, it is not possible to estimate that value, which is the case for [Friederich_climate_risk_disclosure].
+Similarly a constant baseline help understand the performance of a regression model.
+Greenscreen and bronzini_glitter_2023 evaluated the prediction of ESG score, yet they do not provide baseline.
+If the values are really concentrated, a model which always predict the mean value might have a good performance.
+Unfortunately, without the distribution of values, it is not possible to estimate it.
   • Simple Baselines: If a keyword-based baseline performs well, or even better than a proposed method, then the proposed method is likely an overkill.
-For example, Rouen et al.
-(2023); Clarkson et al.
-(2020) showed that even for the prediction of an ESG score (which is a complex target), linguistic features such as the quantity and complexity of text are good predictors.
-This suggests that simple predictors could perform well even on complex tasks.
-Indeed, simple baselines can outperform finetuned approaches, for example on TCFD classification (Sampson et al., 2022) or on the ClimaINS dataset (Spokoyny et al., 2023).
+And yet, less than half of the works we studied report the performance of a keyword-based baseline.
+Studies such as rouenEvolutionESGReports2023, clarkson_nlp_us_csr, which showed that linguistic features such as the quantity and complexity of text are good predictors for the ESG score, therefore seemingly simple predictors (e.g.
+a keyword-based classifier) might perform surprisingly well, even on complex tasks.
+Moreover, we observed that the simple baseline can even outperform finetuned approaches, for example on TCFD classification [sampson_tcfd-nlp_nodate] or on ClimaINS dataset [spokoyny2023answering].
 Beyond a validation (or invalidation) of the proposed approaches, a keyword-based baseline would also allow us to understand if models rely on an “understanding” of the text or merely on word cues.
-And yet, less than half of the works we studied report the performance of a keyword-based baseline.
   • Human Baselines: Using a human as a baseline can help determine whether a task is well-defined.
-If even a qualified human cannot achieve high accuracy, this most likely means that the task is subjective, ill-defined, or in disagreement with the dataset.
+If even a qualified human cannot achieve high accuracy, this most likely means that the task is subjective, ill-defined, or in disagreement with the gold standard.
 Vice versa, if a human achieves high accuracy, but the model achieves even higher accuracy, this could mean that the model has acquired expert knowledge.
-However, with the exception of 3 studies, all studies discussed here not come with a human evaluation.
-6.2. Model Robustness
+However, most of the studies discussed here do not incorporate an evaluation of human performance (Only 3 studies reported Human performances).
+6.2 Model Robustness
 When evaluating a model, it is important to know whether the performance on a dataset is due to an inherent capacity of the model or due to random factors that may be entirely different on another dataset.
@@
 data sampling, hyper-parameters, random initialization, model architecture).
-Thus, a better performance in an experiment could be due also to these factors rather than due to the choice of the model.
-This is even more true when studies report differences in performance of less than 1%.
-For this reason, the uncertainty of the evaluation should always be quantified – by confidence intervals, p-values or at least the standard deviation.
-However, we found that only seven papers, out of the 61 we studied, quantify the uncertainty of the model performance (Webersinke et al., 2022; Leippold and Varini, 2020; Bingler et al., 2023; Schimanski et al., 2024a, 2023b; Lee and Kim, 2023; Schimanski et al., 2023a).
-Noisy Text and Adversarial Attacks
-Most studies mentioned in this survey trained and evaluated models on datasets that contain only clean examples.
-However, real-world data is usually noisy.
-Thus, the models may perform entirely differently on real-world datasets.
-Studies should quantify not just the performance of the models, but also their robustness, as fine-tuned models are prone to word-based adversarial attacks (Woloszyn et al., 2021).
+Moreover, we see that many studies report differences in performance of less than 1%.
+This difference might be due to the choice of model, but it could also be due to the other factors.
+Quantifying the uncertainty induced by those factors helps understand if a difference is significant.
+This could be done using confidence intervals, p-values or simply reporting the standard deviation.
+However, we found only seven papers that quantify the uncertainty of the model performance  [nicolas_webersinke_climatebert_2021, varini_climatext_2020, bingler2023cheaptalkspecificitysentiment, Schimanski2024nature, schimanski_bridging_2023, LEE2023119726, tobias_schimanski_climatebert-netzero_2023].
+Adversarial Attacks and Noisy Text
+While fine-tuning leads to the best performances, it is also prone to word-based adversarial attacks  [vinicius_woloszyn_towards_2021], as the training data are often relatively small, and might not include syntactically diverse examples.
+Noisy texts are an inherent part of real-world applications.
+Yet, most studies mentioned in this review trained and evaluated models on clean datasets (containing only clean examples) and did not quantify the robustness of their approaches to noise.
 Imbalanced Datasets
-The class of interest is relatively rare in most tasks.
-For example, only 5% of text is in the positive class “climate-related” in the study of Yu et al.
-(2024).
+In most of the tasks, the class of interest is relatively rare.
+For example, only 5% of text is in the positive class “climate-related” in the study of yu_climatebug_2024.
 This means that (1) a simple baseline that always says “non-climate-related” will have a very high precision, and (2) a model trained on such a dataset would probably overfit.
 Therefore, most studies mentioned in this review trained and evaluated the models on relatively balanced datasets.
-On the other hand, a model evaluated only on a balanced dataset may fail when tested on real-world, unbalanced datasets (Friederich et al., 2021).
+On the contrary, a model that was evaluated only on a balanced dataset may fail spectacularly when tested on real-world, unbalanced datasets  [Friederich_climate_risk_disclosure].
 For example, a model reaching an F1-score of 90% on a balanced dataset would significantly drop on an unbalanced dataset with 5% of text in the positive class (to a binary F1-score of 47%, and a macro F1-score of 71%).
 This is due to the large number of false positives.
-Thus, studies should consider both the balanced and the unbalanced case.
-Precision-Recall Trade-off
-In most methods (and most tasks), higher precision can be traded for smaller recall.
-In the case of greenwashing, both metrics can be useful: a higher recall is desirable when providing leads for human analysis, so as to avoid missing potential greenwashing cases.
+For tasks related to greenwashing detection, annotating data is a time-consuming and resource-intensive process, often resulting in an imbalanced dataset that fails to represent all relevant cases adequately.
+This imbalance complicates the model’s performance trade-off: a higher recall is desirable when providing leads for human analysis to avoid missing potential greenwashing cases.
 However, if the goal is to accurately measure greenwashing prevalence, minimizing false positives is critical, as false accusations can have serious repercussions.
-Most studies, however, are conducted without the use case in mind.
-6.3. Data Sources
+6.3 Data Sources
 Close-source data and reproducibility
-The studies mentioned in our review use various data sources, most of which are publicly available: annual and sustainable reports  (Doran and Quinn, 2009; Kölbel et al., 2021; Chou et al., 2023; Huang et al., 2020; Kraus et al., 2023; Kheradmand et al., 2021; Rouen et al., 2023; Lai and Chen, 2023; Kang and Kim, 2022; Auzepy et al., 2023; Brié et al., 2022; Webersinke et al., 2022; Luccioni et al., 2020; Kheradmand et al., 2021; Bronzini et al., 2023; Lee and Kim, 2023; Polignano et al., 2022; Kraus et al., 2023), social media (Woloszyn et al., 2021; Luo et al., 2020; Vaid et al., 2022; Coan et al., 2021; Oppong-Tawiah and Webster, 2023; Sharma et al., 2024), specialized websites (Morio and Manning, 2023; Schimanski et al., 2023a; Jin et al., 2022; Leippold and Varini, 2020; Coan et al., 2021).
-However, multiple works rely on closed-source data: ESG scores, carbon emissions, and machine-readable reports that are collected through private providers such as Refinitiv, Reuters, RavenPack, Trucost, Bloomberg, RobecoSAM, Compustat or Wharton (Sautner et al., 2023; Bingler et al., 2023; Stammbach et al., 2023; Friederich et al., 2021; Brié et al., 2022; Lee and Raschke, 2023; Webersinke et al., 2022; Li et al., 2020; Clarkson et al., 2020; Li et al., 2020; Sautner et al., 2023; Mehra et al., 2022; Sharma et al., 2024; Clarkson et al., 2020; Schimanski et al., 2023b; Vinella et al., 2023; Kraus et al., 2023).
-Relying on private providers limits the reproducibility of these works, for both academic work and applications.
+The studies mentioned in our review use various data sources, most of which are publicly available (annual and sustainable reports used in 18 studies)social media (6 studies)specialized website (5 studies).
+However, multiple works rely on closed-source data, this include ESG score, carbon emissions, machine-readable reports that are collected through private providers such as Refinitiv, Reuters, RavenPack, Trucost, Bloomberg, RobecoSAM, Compustat or Wharton (17 studies).
+Relying on private providers limits the reproducibility of those works, for both academic work and applications.
 No greenwashing datasets
-The most fundamental shortcoming of current work on greenwashing is that there is no comprehensive real-world dataset with instances of greenwashing.
-This limits the conclusions that can be drawn from current studies drastically.
-The lack of a gold standard dataset can be explained by several factors:
-  • The definitions of greenwashing are often vague and ambiguous, leaving room for interpretation, and making it difficult to annotate examples of greenwashing.
-  • Green claims are often not self-contained, but spread across multiple paragraphs in a same document, across multiple documents, or even across different types of sources (such as tables, databases, etc.).
-Therefore, just collecting all the elements required to correctly identify greenwashing is a challenge on its own.
-  • Understanding the claim and annotating greenwashing requires expertise, and sometimes even debate, hence producing greenwashing annotations is a difficult task.
-  • Greenwashing is a strong accusation that might impact a company reputation, and that can also, if not substantiated by evidence, be considered a defamation.
-To build a dataset of real-world greenwashing cases, several data collection strategies can be considered.
-To avoid relying on a subjective annotation, researchers can rely on a third-party judgment.
-For example, using reputable news articles that explicitly discuss greenwashing to find actual examples, drawing on prior research focused on detecting media mentions of greenwashing (Gourier and Mathurin, 2024).
-Another valuable resource are regulatory records from authorities such as the Advertising Standards Authority (ASA) or databases like greenwash.com.
-Missing link to regulatory texts
-Existing work often relies on its own definitions of greenwashing.
-This risks introducing subjectivity and incompatibility with other works.
-A safer choice is to use regulatory texts as a reference, as this is done, e.g., in Stammbach et al.
-(2023).
-Recently, the European Commission adopted the corporate sustainability reporting directive (CSRD) (Union, 2022), making environmental disclosure requirements explicit, and the directive on green claims (US Environmental Protection Agency, 2023), which defines requirements when making green claims.
-These could be good starting points for defining greenwashing in future research.
-Such regulations help not just with the definition of greenwashing, but also with its judgment.
-Instead of emitting a judgment, a method can link a statement to a regulatory text.
-For example, the works that predict TCFD simply link a statement to a recommendation (Ding et al., 2023; Auzepy et al., 2023; Brié et al., 2022; Bingler et al., 2022; Sampson et al., 2022).
-This approach makes the methods less prone to ad hoc definitions of scores and judgment strategies, and lets a human take the final decision.
+While theoretical frameworks have provided a starting point for defining greenwashing, a critical gap remains: the lack of a comprehensive, real-world dataset.
+This can be explained by several factors:
+  • Greenwashing definitions are often vague and ambiguous, leaving room for interpretation, making it difficult to use for annotating examples of greenwashing
+  • Green claims are often not self-contained, meaning that the full content of information required to judge if a claim is misleading or not is spread: across multiple paragraphs in a same document, or even across multiple documents.
+Therefore collecting all the elements required to correctly identify greenwashing is a challenges on its own.
+  • Understanding the claim and annotating greenwashing requires expertise, and sometimes debates to make sure the communication is misleading, making the annotation task even harder.
+  • Greenwashing is a strong accusation, that might impact a company reputation, and that can also, if not substantiated with evidence, be considered as defamation.
+To address these challenges, several potential data collection strategies can be considered.
+To avoid relying on a subjective annotation, researchers can rely on a third-party judgement.
+For example, using news articles that explicitly discuss greenwashing, drawing on prior research focused on detecting media mentions of greenwashing  [gourierGreenwashingIndex2024].
+Another valuable resource could include regulatory records from authorities such as the Advertising Standards Authority (ASA) or databases like greenwash.com.
+Alternatively, previous works tried to produce signals that could potentially indicate greenwashing, future research might rely on such indicators to first filter-out texts that are potentially containing greenwashing, and annotate those selected texts.
+Intermediary tasks are also important for other reasons.
+First, decomposing tasks often results in better performance, as supported by the principles behind Chain-of-Thought.
+Second, when AI systems are designed to include human oversight, it is crucial for these intermediary tasks to remain understandable and accessible to users.
+Thirdly, the design of energy-efficient AI systems can benefit from employing simpler models for tasks where high precision is not required, such as document retrieval for RAG.
+Finally, relying on intermediary tasks for the evaluation of models’ performances is necessary for both understanding better the models, their limits, and to propose auditable pipelines with humans in the loop.
+Relying on regulatory texts
+As mentioned above, two of the challenges in building a dataset of examples of greenwashing are using definitions that are subjective and mitigating risks of defamation.
+Using regulatory texts as reference helps mitigate those two issues.
+Firstly, it can help build a proper definition of the label.
+For example, stammbach_environmental_2023 rely on the definition of green claims by the European Commission to build their definition of environmental claims.
+Secondly, instead of emitting a judgment, the method can link a statement to a regulatory text.
+For example, all works focusing on TCFD simply link a statement to a recommendation [dingCarbonEmissionsTCFD2023, auzepy_evaluating_2023, bjarne_brie_mandatory_2022, bingler_cheap_2021, sampson_tcfd-nlp_nodate].
+Recently, the European Commission adopted the CSRD which defines more precisely greenwashing.
+Future research could start by finding content related to parts of the regulatory texts related to greenwashing.
 Quantifying communication vs Quantifying information
-With the exception of Bronzini et al.
-(2023), the studies reviewed in this work all measure how often a given text contains climate-related content, TCFD disclosures, or specific topics of interest.
-However, a company can repeatedly emphasize the same objective or repeatedly highlight a single initiative.
-With this, they skew the quantity of climate-related content in their favor, while they may still fail to implement any meaningful action.
-This phenomenon has been called “cheap talk” (Bingler et al., 2023).
-Future research should thus quantify not only how often some pieces of information are conveyed, but also how many different pieces there are.
-7. Conclusion
+With the exception of bronzini_glitter_2023, the studies reviewed in this work primarily focus on quantifying textual content related to climate, TCFD disclosures, or specific topics of interest.
+However, as highlighted by bingler2023cheaptalkspecificitysentiment, companies increasingly engage in ”cheap talk,” where statements lack substantive commitments or actionable details.
+This underscores the need to assess not only the volume of climate-related communication but also the informational content conveyed.
+For instance, a company might repeatedly emphasize its objectives or highlight a single initiative, creating the illusion of widespread organizational change, while failing to implement meaningful action.
+Future research should focus on quantifying the information instead of the amount of text related to a topic.
+Restricting the definition of greenwashing
+Except for obvious lies, what counts as “misleading communication” is intrinsically subjective.
+For some, disproportionately highlighting a company’s green initiatives is misleading (LEE_greenwashing); others interpret the same behavior as employees’ pride in early achievements or as a strategic shift to engage stakeholders around more ambitious climate action.
+Likewise, hyewon_kang_analyzing_2022 frame greenwashing as overly positive texts, yet such positivity can be a stylistic choice rather than deceptive intent.
+Moving from a broad to a narrower definition can improve the objectivity of what is considered greenwashing.
+The trade-off with narrower definitions is twofold: (i) they may exclude other forms of greenwashing, and (ii) the definition itself can then be attacked for not adequately defining greenwashing.
+Still, this is preferable to relying on a broad but vague definition, which would produce unreliable and subjective annotations.
+Future work should further explore these and other narrow but practical definitions, as they remain useful starting points for uncovering greenwashing practices.
+7 Conclusion
 In this survey, we have provided a comprehensive overview of current research initiatives on detection of greenwashing in textual corporate communications.
@@
 This makes the annotation of actual examples of greenwashing challenging, and indeed, there is currently no large-scale dataset of instances of greenwashing.
-For this reason, the detection of greenwashing has remained on a theoretical level without empirical validation (Moodaley and Telukdarie, 2023a).
+For this reason, the detection of greenwashing has remained on a theoretical level without empirical validation [moodaley_conceptual_2023].
 Most existing works target intermediate tasks such as identifying climate-relatedness or green claims.
 This work also serves its purposes: First, determining whether a text can contain greenwashing at all helps restricting attention to the relevant resources.
-Second, decomposing greenwashing into more manageable steps can lead to an overall better performance, as supported by the principles behind the Chain-of-Thought (Wei et al., 2022).
+Second, decomposing greenwashing into more manageable steps can lead to an overall better performance, as supported by the principles behind the Chain-of-Thought [cot].
 Third, using frugal approaches helps reduce costs and energy consumption.
@@
 We hope that future research will help identify and combat greenwashing more reliably, thereby fostering transparency and accountability in climate communication.
-Limitations
-This survey provides a comprehensive overview of studies on automatic detection of misleading climate-related corporate communications and its related tasks as of early 2025.
-A number of limitations remain: Given the extensive scope of the field, our review concentrated on NLP-based research, providing only limited discussion of topics like ESG prediction, which are predominantly explored through quantitative methods.
-Furthermore, our review relied primarily on keyword searches on arXiv, Google Scholar, and ResearchGate, complemented by references from identified papers.
-In this, we may have overlooked relevant studies that use alternative terminology or are published in less accessible venues.
-Finally, much of greenwashing happens not on the text level, but on the factual level (with claims that contradict data) or on the multi-modal level (across images or videos).
-These remain out of scope for our survey.
-Acknowledgment
-We would like to express our sincere gratitude to Nina Godart and AEF Info for their invaluable contributions and knowledge on the topic.
-We also extend our thanks to Takaya Sekine CFA for his insightful comments.
-This work was partially supported by the grant ANR-20-CHIA-0012-01 (“NoRDF”).
-Appendix
+[…]
+Acknowledgements
+We are grateful to Takaya Sekine for his insightful feedback on the manuscript.
+We further acknowledge Amundi’s climate analysts, notably Aaron Macdougall, and the journalists at AEF Info, in particular Nina Godart, for their assistance in researching the topic and refining definitions.
+Declarations
+Funding
+This work is funded by Amundi Technology, Inria, Télécom Paris, and ANRT through a CIFRE partnership program.
+Conflicts of interest/Competing interests
+Aside from the authors’ employers—Amundi Technology, Inria, and Télécom Paris—there are no conflicts of interest related to the present work.
+Ethics approval and consent to participate
+Not applicable.
+Consent for publication
+Not applicable.
+Data availability
+Not applicable.
+Materials availability
+Not applicable.
+Code availability
+Not applicable.
+Author contribution
+This work was primarily conducted by Tom Calamai, particularly the research efforts (collection and analysis of papers). All authors contributed to the writing of the manuscript.
+Preprint
+An earlier version of this work is available as a preprint on arXiv [calamai2025corporategreenwashingdetectiontext].
 Appendix A TCFD recommendations
@@
   • Describe the targets used by the organization to manage climate-related risks and opportunities and performance against targets.
-A.1. Data sources
+Appendix B Data sources
 source papers
-10-K reports (Doran and Quinn, 2009; Kölbel et al., 2021; Chou et al., 2023; Huang et al., 2020; Kraus et al., 2023)
-Manually collected reports (Kheradmand et al., 2021; Rouen et al., 2023; Lai and Chen, 2023; Kang and Kim, 2022; Auzepy et al., 2023; Brié et al., 2022; Webersinke et al., 2022; Luccioni et al., 2020; Kheradmand et al., 2021; Bronzini et al., 2023; Lee and Kim, 2023; Polignano et al., 2022; Kraus et al., 2023)
-Refinitiv (Sautner et al., 2023; Bingler et al., 2023; Stammbach et al., 2023; Friederich et al., 2021; Brié et al., 2022; Lee and Raschke, 2023; Webersinke et al., 2022)
-Reuters (Li et al., 2020; Clarkson et al., 2020)
-ESG score private providers (Li et al., 2020; Sautner et al., 2023; Mehra et al., 2022; Sharma et al., 2024; Clarkson et al., 2020; Schimanski et al., 2023b)
-CDP (Spokoyny et al., 2023)
-Open for Good Data (Vinella et al., 2023)
-ClimateWatch (Kraus et al., 2023)
-Twitter/Reddit (Woloszyn et al., 2021; Luo et al., 2020; Vaid et al., 2022; Coan et al., 2021; Oppong-Tawiah and Webster, 2023; Sharma et al., 2024)
-LobbyMap (Morio and Manning, 2023)
-Net-zero tracker (Schimanski et al., 2023a)
-Fact-checking websites (Jin et al., 2022; Leippold and Varini, 2020; Coan et al., 2021)
-Table 10. Data Sources
+10-K reports [doran_risk_disclosure, kolbel_ask_2021, chou_ESG, huangFinBERTLargeLanguage2020, krausEnhancingLargeLanguage2023]
+Manually collected reports [kheradmand2021a, rouenEvolutionESGReports2023, lai_using_2023, hyewon_kang_analyzing_2022, auzepy_evaluating_2023, bjarne_brie_mandatory_2022, nicolas_webersinke_climatebert_2021, luccioni_analyzing_2020, kheradmand2021a, bronzini_glitter_2023, LEE2023119726, marco_polignano_nlp_2022, krausEnhancingLargeLanguage2023]
+Refinitiv [SAUTNER_cliamte_change_exp, bingler2023cheaptalkspecificitysentiment, stammbach_environmental_2023, Friederich_climate_risk_disclosure, bjarne_brie_mandatory_2022, LEE_greenwashing, nicolas_webersinke_climatebert_2021]
+Reuters [liCorporateClimateRisk2020, clarkson_nlp_us_csr]
+ESG score private providers [liCorporateClimateRisk2020, SAUTNER_cliamte_change_exp, Mehra_2022, Greenscreen, clarkson_nlp_us_csr, schimanski_bridging_2023]
+CDP [spokoyny2023answering]
+Open for Good Data [avalon_vinella_leveraging_2023]
+ClimateWatch [krausEnhancingLargeLanguage2023]
+Twitter/Reddit [vinicius_woloszyn_towards_2021, luo_detecting_2020, vaid-etal-2022-towards, coanComputerassistedClassificationContrarian2021, divinus_oppong-tawiah_corporate_2023, Greenscreen]
+LobbyMap [morio2023an]
+Net-zero tracker [tobias_schimanski_climatebert-netzero_2023]
+Fact-checking websites [jin-etal-2022-logical, varini_climatext_2020, coanComputerassistedClassificationContrarian2021]
+Table 12: Data Sources
 Publicly available Reports
-Many studies rely on 10-K reports (Doran and Quinn, 2009; Kölbel et al., 2021; Chou et al., 2023; Huang et al., 2020; Kraus et al., 2023) as they are easily accessible (EDGAR), uniformly structured, and mandatory for companies listed on US financial markets.
+Many studies rely on 10-K reports [doran_risk_disclosure, kolbel_ask_2021, chou_ESG, huangFinBERTLargeLanguage2020, krausEnhancingLargeLanguage2023] as they are easily accessible (EDGAR), uniformly structured, and mandatory for companies listed on US financial markets.
 Other reports, such as sustainable reports, are not centralized and mandatory, therefore they need to be collected from other sources.
-They are often collected manually from the company’s website (Kheradmand et al., 2021; Rouen et al., 2023; Lai and Chen, 2023; Kang and Kim, 2022; Auzepy et al., 2023; Brié et al., 2022; Webersinke et al., 2022), or from other websites (Luccioni et al., 2020; Kheradmand et al., 2021; Bronzini et al., 2023; Lee and Kim, 2023; Polignano et al., 2022; Kraus et al., 2023) such as from the GRI, SASB, TCFD, ResponsibilityReports, AnnualReports, Korean Standards Association.
-Polignano et al.
-(2022) relied on Sustainability Disclosure Database from GRI.
-To the best of our knowledge, it is not available any longer.
-The reports are openly available, however they are most of the time in a PDF format, and must be parsed to produce machine-readable versions.
-While Khan (2024) proposed a machine-readable dataset for the 10-K reports, it is often not the case for other types of reports.
-(Bronzini et al., 2023; Polignano et al., 2022; Kang and Kim, 2022) provide a sample of processed reports.
+They are often collected manually from the company’s website [kheradmand2021a, rouenEvolutionESGReports2023, lai_using_2023, hyewon_kang_analyzing_2022, auzepy_evaluating_2023, bjarne_brie_mandatory_2022, nicolas_webersinke_climatebert_2021], or from other websites [luccioni_analyzing_2020, kheradmand2021a, bronzini_glitter_2023, LEE2023119726, marco_polignano_nlp_2022, krausEnhancingLargeLanguage2023] such as from the GRI, SASB, TCFD, ResponsibilityReports, AnnualReports, Korean Standards Association.
+marco_polignano_nlp_2022 relied on Sustainability Disclosure Database from GRI.
+To the best of our knowledge, it is no longer available.
+The reports are openly available; however, they are most of the time in a PDF format, and must be parsed to produce machine-readable versions.
+While financial-reports-sec proposed a machine-readable dataset for the 10-K reports, it is often not the case for other types of reports.
+[bronzini_glitter_2023, marco_polignano_nlp_2022, hyewon_kang_analyzing_2022] provide a sample of processed reports.
 Private Providers
-Machine-readable reports, as well as earning call transcripts are often provided through a private data provider such as Refinitiv (Sautner et al., 2023; Bingler et al., 2023; Stammbach et al., 2023; Friederich et al., 2021; Brié et al., 2022; Lee and Raschke, 2023; Webersinke et al., 2022) or Reuters (Li et al., 2020; Clarkson et al., 2020).
+Machine-readable reports, as well as earning call transcripts are often provided through a private data provider such as Refinitiv [SAUTNER_cliamte_change_exp, bingler2023cheaptalkspecificitysentiment, stammbach_environmental_2023, Friederich_climate_risk_disclosure, bjarne_brie_mandatory_2022, LEE_greenwashing, nicolas_webersinke_climatebert_2021] or Reuters [liCorporateClimateRisk2020, clarkson_nlp_us_csr].
 Similarly, financial data (e.g.
-ESG score) is often collected through data providers such as RavenPack or S&P Global Trucost or Wharton’s research platform WRDS, Reuters, Bloomberg, Refinitiv, RobecoSAM and Compustat (Li et al., 2020; Sautner et al., 2023; Mehra et al., 2022; Sharma et al., 2024; Clarkson et al., 2020; Schimanski et al., 2023b)
+ESG score) is often collected through data providers such as RavenPack or S&P Global Trucost or Wharton’s research platform WRDS, Reuters, Bloomberg, Refinitiv, RobecoSAM and Compustat [liCorporateClimateRisk2020, SAUTNER_cliamte_change_exp, Mehra_2022, Greenscreen, clarkson_nlp_us_csr, schimanski_bridging_2023]
 Carbon emission data
-Carbon emissions of companies can be collected through the private data providers, or through CDP (Spokoyny et al., 2023), Open for Good Data (Vinella et al., 2023) or ClimateWatch website (Kraus et al., 2023).
+Carbon emissions of companies can be collected through the private data providers, or through CDP [spokoyny2023answering], Open for Good Data [avalon_vinella_leveraging_2023] or ClimateWatch website [krausEnhancingLargeLanguage2023].
 Media and Social Media
-As companies communicate with the public mainly through marketing and social media, other data sources include Twitter and Reddit (Woloszyn et al., 2021; Luo et al., 2020; Vaid et al., 2022; Coan et al., 2021; Oppong-Tawiah and Webster, 2023; Sharma et al., 2024).
+As companies communicate with the public primarily through marketing and social media, other data sources include Twitter and Reddit [vinicius_woloszyn_towards_2021, luo_detecting_2020, vaid-etal-2022-towards, coanComputerassistedClassificationContrarian2021, divinus_oppong-tawiah_corporate_2023, Greenscreen].
 Climate expert websites
-Finally, projects such as LobbyMap or Net Zero Tracker, that track corporate policy and engagements, can help build annotated datasets (Morio and Manning, 2023; Schimanski et al., 2023a).
-Similarly, relying on a fact-checking website such as https://climatefeedback.org/ or www.skepticalscience.com is an efficient way of building a dataset (Jin et al., 2022; Leippold and Varini, 2020; Coan et al., 2021)
-Appendix B Detailed model list
-Table 11. Solutions proposed in each paper
-Paper Dataset Baselines BERT Sent. Sim. Other
-Pretraining Models on Climate-Related Text
-Webersinke et al. (2022) Domain-specific corpus - Pre-Training ClimateBERT
-Vaghefi et al. (2022) Deep Climate Change - Pre-Training ClimateGPT-2
-Mullappilly et al. (2023) Deep Climate Change - Pre-Training stsb-xlm-r-multilingual Arabic Mini-ClimateGPT, Vicuna, Alpaca, Dolly V2, Alpaca ar
-Schimanski et al. (2023b) Domain-specific corpus - Pre-Training EnvironmentalBERT
-Yu et al. (2024) climateBUG-data - Pre-Training climateBUG-LM
-Thulke et al. (2024) Domain-specific corpus - Pre-Training ClimateGPT-7/13/70B
-Thulke et al. (2024) ClimaBENCH - Evaluation Stability-3B, Pythia-6.9B, Falcon-7B, Mistral-7B, Llama-2-7B, Jais-13B, Jais-13B-Chat, Llama-2-7/13/70B, Llama-2-Chat-7/13/70B, ClimateGPT-7/13/70B
-Climate-related topic detection
-Doran and Quinn (2009) 10-K filings - climate Keywords
-Engle et al. (2020) WSJ - climate TF-IDF cosine similarity
-Leippold and Varini (2020) climatext NB, Keywords BERT
-Garrido-Merchán (Garrido-Merchán et al., 2023) climatext ClimateBERT, BERT
-Spokoyny et al. (2023) climatext Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
-Webersinke et al. (2022) climate detection distilRoBERTa, climateBERT
-Bingler et al. (2023) climate detection NB, SVM + BoW, SVM + ELMo climateBERT
-Brié et al. (2022) climate detection Random Forest climateBERT, FinBERT
-Yu et al. (2024) climateBUG data BERT, FinBERT, ClimateBERT, climateBUG-LM
-Lin et al. (2023b) sustainable signals review
-Climate-related subtopic detection
-Kang and Kim (2022) Sustainable Reports - topics sentence-BERT
-Mishra and Mittal (2021) sciDCC
-Spokoyny et al. (2023) sciDCC Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
-Vaid et al. (2022) climateEng BERT, RoBERTa, DistilBERT FastText
-Spokoyny et al. (2023) climateEng Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
-Jain et al. (2023) ledger - topics TF-IDF BERT, ClimateBERT and RoBERTa. sentence-BERT Word2Vec-based Random Forest
-Schimanski et al. (2024a) ESGBERT-nature Keywords ClimateBERT, RoBERTa, DistilRoBERTa, EnvironmentalBERT
-Spokoyny et al. (2023) ClimaTOPIC Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
-Detecting climate-related financial disclosure
-Ding et al. (2023) Reports - Disclosure TF-IDF cosine similarity
-Auzepy et al. (2023) Reports - TCFD BART+MNLI
-Brié et al. (2022) Sustainable Reports structural topic model (STM)
-Bingler et al. (2022) climate tcfd rec. TF-IDF + random forest RoBERTa, ClimateBERT (RoBERTa+LogReg) Sentence Enc.
-Sampson et al. (2022) Reports - TCFD TF-IDF ClimateBERT, distilRoBERTa kNN (sentence-transformers embeddings), stacked models
-In-depth Disclosure: Climate Risk Classification
-Li et al. (2020) Reports - Risk dictionaries (risk)
-Chou et al. (2023) 10-K filings - Risk 3 dictionaries (risk: Climate, Impact, Litigious)
-Kheradmand et al. (2021) Annual/Sustainability Reports - Risk dictionaries (risk)
-Sautner et al. (2023) earning calls - Risk dictionaries (risk/opportunity)
-Polignano et al. (2022) Sustainability Reports - Sentiment TextBlob, Sentita
-Kang and Kim (2022) Sustainability Reports - Sentiment DistilBERT (sentiment analysis)
-Kölbel et al. (2021) 10-K filings - Risk BoW, TF-IDF BERT
-Friederich et al. (2021) Reports - Risk SVM distilBERT, RoBERTa
-Bingler et al. (2023) climate sentiment NB, SVM + BoW, SVM + ELMo distilRoBERTa, climateBERT
-Webersinke et al. (2022) climate sentiment distilRoBERTa, climateBERT
-Xiang and Fujii (2023) Web of Science, Scopus - Sentiment BERT, Domain-specific BERT, TinyBERT, DistilBERT, ClimateBERT CCLA+Max-Pooling, Bi-LSTM-Attention, POS-Bi-LSTM-Attention, DARE
-Detecting environmental, social, and governance disclosure
-Rouen et al. (2023) Sustainability Reports - ESG industry-topic dictionaries
-Mehra et al. (2022) Sustainability Reports - ESG sentence BERT, Universal Sentence Encoder, and Deep Averaging Network
-Huang et al. (2020) Sustainability Report, 10-K filings - ESG NB, SVM, RF FinBERT, BERT CNN, LSTM
-Schimanski et al. (2023b) ESGBERT E,S,G BERT, RoBERTa, DistilRoBERTa, E/S/G-RoBERTa, E/S/G-DistilRoBERTa
-Lee and Kim (2023) Korean Reports - ESG Multilingual BERT, KoBERT, DistilKoBERT, KB-ALBERT, KLUE-RoBERTa
-Bronzini et al. (2023) Sustainability Reports - ESG INSTRUCTOR-xl
-Green claim detection
-Woloszyn et al. (2021) green claims RoBERTa, BERTweet Flair
-Stammbach et al. (2023) environmental claims Random, Majority, TF-IDF + SVM, Character n-gram SVM DistilBERT, ClimateBERT, RoBERTa ClaimBuster RoBERTa, Pledge Detection RoBERTa
-Lin et al. (2023b) Product Description - Claim environmental claim detection (finetuned distilBERT)
-Green claim characteristics
-Bingler et al. (2023) climate-commitments-actions NB, SVM + BoW, SVM + ELMo distilRoBERTa, climateBERT
-Schimanski et al. (2023a) netzero/reduction ClimateBERT, DistilRoBERTa, RoBERTa GPT-3.5-turbo
-Woloszyn et al. (2021) green claims 3 RoBERTa, BERTweet Flair
-Bingler et al. (2023) climate specificity NB, SVM + BoW, SVM + ELMo distilRoBERTa, climateBERT
-Green stance detection
-Morio (Morio and Manning, 2023) LobbyMap Majority, TF-IDF BERT, RoBERTa, ClimateBERT, Longformer
-Luo et al. (2020) DeSMOG
-Vaid et al. (2022) ClimateStance BERT, RoBERTa, DistilBERT FastText
-Lai and Chen (2023) Reports - Stance BERT, RoBERTa, ClimateBERT
-Spokoyny et al. (2023) ClimateStance Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
-Spokoyny et al. (2023) DeSMOG Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
-Leippold (Leippold and Diggelmann, 2020) ClimateFEVER ALBERT(FEVER)
-Wang et al. (2021) ClimateFEVER RoBERTa (ClimateFEVER) RoBERTa (FEVER), RoBERTa (ClimateFEVER) + UDA
-Spokoyny et al. (2023) ClimateFEVER Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
-Vaghefi et al. (2022) ClimateFEVER GPT-2, climateGPT-2
-Xiang and Fujii (2023) ClimateFEVER BERT, RoBERTa, DistilRoBERTa, ClimateBERT, POS-Bi-LSTM-Attention
-Webersinke et al. (2022) ClimateFEVER distilRoBERTa, climateBERT
-Vaid et al. (2022) ClimateFEVER BERT, RoBERTa, DistilBERT FastText
-Question Answering
-Luccioni et al. (2020) Reports - QA RoBERTa
-Spokoyny et al. (2023) ClimaQA BM25 MiniLM
-Spokoyny et al. (2023) ClimaINS Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
-Rony et al. (2022) CCMRC - QA pairs ALBERT (trained on SQuAD then finetuned) Sentence-BERT
-Schimanski et al. (2023a) Net-Zero Tracker - Target RoBERTa Quad2
-Ni et al. (2023) Reports - RAG ChatGPT
-Vaghefi et al. (2023) Reports - RAG - Human Eval GPT-4
-Mullappilly et al. (2023) Clima500-Instruct - QA instructions Vicuna (Mini-ClimateGPT)
-Thulke et al. (2024) Instruction following corpus - Human Eval ClimateGPT-70B, ClimateGPT-7B, ClimateGPT-FSC-7B
-Thulke et al. (2024) ClimRetrieve ColBERTv2, DRAGON, GTE-base, text-embedding-ada-002, text-embedding-3-small, text-embedding-3-large BM25
-Deceptive Techniques
-Oppong (Oppong-Tawiah and Webster, 2023) Tweets - Fake News linguistic indicators
-Coan et al. (2021) contrarian claims Logistic, SVM RoBERTa ULMFiT (over sample, weighted, focal Loss, logisitic, SVM), RoBERTa + Logistic
-Jin et al. (2022) logicClimate ELECTRA, BERT lectra-StructureAware
-Bhatia et al. (2021) climate change scepticism - Neutralization SVM BERT MTEXT
-Environmental performance prediction
-Jain et al. (2023) Ledger - Scope 3 Emissions Alpaca, WizardLM, Flan-T5, ChatGPT
-Bronzini et al. (2023) Sustainable Reports - ESG OLS
-Mehra et al. (2022) Reports 10-Qs - ESG Majority BERT, ESGBERT
-Clarkson et al. (2020) CSR Reports - CSR performer linguistic features + random forest, XGBoost, SVM
-Lin et al. (2023b) Product Description / Review - Finch Score Lasso, Gradient Boosting NOCATE + transformer, SUSTAINABLESIGNALS + transformer
-Sharma et al. (2024) Multimodel Tweet - ESG CLIP+MLP, Sentence-BERT+MLP
+Finally, projects such as LobbyMap or Net Zero Tracker, which track corporate policy and engagements, can help build annotated datasets [morio2023an, tobias_schimanski_climatebert-netzero_2023].
+Similarly, relying on a fact-checking website such as https://climatefeedback.org/ or www.skepticalscience.com is an efficient way of building a dataset [jin-etal-2022-logical, varini_climatext_2020, coanComputerassistedClassificationContrarian2021]
 Appendix C Reported Performances
@@
 We reported one table per dataset, which might contain multiple tasks, but multiple studies can tackle the same dataset and are then grouped to be able to compare the results.
-C.1. Climate-related Topic Classification
-Source Model F1
-Webersinke et al. (2022) distilRoBERTa 98.6% ±$\pm 1$
-Webersinke et al. (2022) ClimateBERT 99.1% ±$\pm 0.8$
-Bingler et al. (2023) Naive Bayes 87%
-Bingler et al. (2023) SVM + BoW 87%
-Bingler et al. (2023) SVM + ELMo 89%
-Bingler et al. (2023) ClimateBERT 97%
-Brié et al. (2022) ClimateBERT 98.59%
-Brié et al. (2022) FinBERT 96.67%
-Brié et al. (2022) Random Forest 91.65%
-(a) F1-scores (weighted average) on ClimateBERT’s climate detection dataset (Bingler et al., 2023).
-For Webersinke et al.
-(2022) scores are displayed as mean ±$\pm std$.
-Bingler et al.
-(2023) also reported an uncertainty measure (p-value) which is not reported in this table for simplicity.
-Source Model F1
-Yu et al. (2024) BERT 90.81%
-Yu et al. (2024) FinBERT 90.82%
-Yu et al. (2024) ClimateBERT 91.07%
-Yu et al. (2024) climateBUG-LM 91.36%
-(b) F1-scores (macro average) on ClimateBUG-data (Yu et al., 2024).
+Climate-related Topic Classification
 Source Model Wikipedia 10-K Claims
-F1 (bin.) F1 (bin.) F1 (bin.)
-Leippold and Varini (2020) best BERT 80% ±$\pm 6$ 95% $\pm$ 2 83% $\pm$ 1
-Leippold and Varini (2020) best NB 60% ±$\pm 7$ 90% ±$\pm 3$ 72% ±$\pm 2$
-Leippold and Varini (2020) best Keywords 67% ±$\pm 7$ 88% ±$\pm 3$ 70% ±$\pm 1$
-Leippold and Varini (2020) BERT-AL-wiki 69% ±$\pm 6$ 71% ±$\pm 4$ 81% ±$\pm 1$
-Garrido-Merchán (Garrido-Merchán et al., 2023) BERT-AL-wiki - 91% ±$\pm 0.6$
-Garrido-Merchán (Garrido-Merchán et al., 2023) climateBERT-AL-wiki - 93% ±$\pm 0.7$
-Combined - F1 (macro)
-Spokoyny et al. (2023) Majority 42.08%
-Spokoyny et al. (2023) Random 46.86%
-Spokoyny et al. (2023) SVM 83.39%
-Spokoyny et al. (2023) BERT 87.04%
-Spokoyny et al. (2023) RoBERTa 85.97%
-Spokoyny et al. (2023) DistilRoBERTa 86.06%
-Spokoyny et al. (2023) Longformer 87.80%
-Spokoyny et al. (2023) SciBERT 83.29%
-Spokoyny et al. (2023) ClimateBERT 85.14%
-(c) F1-scores on ClimaText (Leippold and Varini, 2020). For Leippold and Varini (2020) scores are displayed as mean ±$\pm std$.
-Table 12. Reported performances for the datasets on climate-related topic detection (which include climate, sustainability and environment).
-C.2. Thematic Analysis
-Source Model Prec.
-Bingler et al. (2022) TF-IDF+random forest 24%
-Bingler et al. (2022) Sentence Enc. 23%
-Bingler et al. (2022) RoBERTa Para. 22%
-Bingler et al. (2022) RoBERTa Sent. 75%
-Bingler et al. (2022) ClimateBERT* 84%
-Bingler et al. (2022) ClimateBERT* blahblF1: 84%
-(a) Average Precisions and F1-score (macro average) on TCFD recommendations classification (Bingler et al., 2022). (*They refer to a finetuned RoBERTa with a Logistic Regression as ClimateBERT)
+binary F1 F1 F1
+varini_climatext_2020 best BERT 80% $\pm$ 6 95 % $\pm$ 2 83% $\pm$ 1
+varini_climatext_2020 best NB 60% $\pm$ 7 90 % $\pm$ 3 72 % $\pm$ 2
+varini_climatext_2020 best Keywords 67% $\pm$ 7 88% $\pm$ 3 70% $\pm$ 1
+varini_climatext_2020 BERT-AL-wiki 69% $\pm$ 6 71% $\pm$ 4 81 % $\pm$ 1
+garridomerchán2023finetuning BERT-AL-wiki - 91% $\pm$ 0.6 -
+garridomerchán2023finetuning climateBERT-AL-wiki - 93% $\pm$ 0.7 -
+Combined
+F1 (macro)
+spokoyny2023answering Majority 42.08%
+spokoyny2023answering Random 46.86%
+spokoyny2023answering SVM 83.39%
+spokoyny2023answering BERT 87.04%
+spokoyny2023answering RoBERTa 85.97%
+spokoyny2023answering DistilRoBERTa 86.06%
+spokoyny2023answering Longformer 87.80%
+spokoyny2023answering SciBERT 83.29%
+spokoyny2023answering ClimateBERT 85.14%
+(a) Reported performance on ClimaText [varini_climatext_2020] (annotated datasets).
+Source Model F1 (macro)
+yu_climatebug_2024 BERT 90.81%
+yu_climatebug_2024 FinBERT 90.82%
+yu_climatebug_2024 ClimateBERT 91.07%
+yu_climatebug_2024 climateBUG-LM 91.36%
+(b) Reported Performances on ClimateBUG-data yu_climatebug_2024.
+Source Model F1 (weighted average)
+nicolas_webersinke_climatebert_2021 distilRoBERTa 98.6% $\pm$ 1
+nicolas_webersinke_climatebert_2021 ClimateBERT 99.1% $\pm$ 0.8
+bingler2023cheaptalkspecificitysentiment Naive Bayes 87%
+bingler2023cheaptalkspecificitysentiment SVM + BoW 87%
+bingler2023cheaptalkspecificitysentiment SVM + ELMo 89%
+bingler2023cheaptalkspecificitysentiment ClimateBERT 97%
+bjarne_brie_mandatory_2022 ClimateBERT 98.59%
+bjarne_brie_mandatory_2022 FinBERT 96.67%
+bjarne_brie_mandatory_2022 Random Forest 91.65%
+(c) Performances on ClimateBERT’s climate detection dataset bingler2023cheaptalkspecificitysentiment.
+bingler2023cheaptalkspecificitysentiment also reported an uncertainty measure (p-value) which is not reported in this table for simplicity.
+Table 13: Reported performance for the datasets on climate-related topic detection (which include climate, sustainability and environment)
+Thematic Analysis
+Source Model Average P F1 (macro)
+bingler_cheap_2021 TF-IDF+random forest 24 % -
+bingler_cheap_2021 Sentence Enc. 23% -
+bingler_cheap_2021 RoBERTa Para. 22% -
+bingler_cheap_2021 RoBERTa Sent. 75% -
+bingler_cheap_2021 ClimateBERT (RoBERTa+LogReg) 84% 84%
+(a) Performance on TCFD recommendations classification [bingler_cheap_2021] with 5 labels
 Source Model ROC-AUC
-Sampson et al. (2022) TF-IDF (1) 0.867
-Sampson et al. (2022) kNN (2) 0.847
-Sampson et al. (2022) ClimateBERT (3) 0.852
-Sampson et al. (2022) distilRoBERTa (4) 0.819
-Sampson et al. (2022) (1) + (2) 0.884
-Sampson et al. (2022) (1) + (3) 0.865
-Sampson et al. (2022) (1) + (4) 0.857
-(b) ROC-AUC scores on TCFD classification task (Sampson et al., 2022)
-Source Model F1
-Auzepy et al. (2023) BART+MNLI 56.68%
-(c) F1-scores (macro average) on TCFD classification (Auzepy et al., 2023)
-Table 13. Reported performances for the datasets on TFCD classification.
-Source Model F1
-Huang et al. (2020) FinBERT 89.6%
-Huang et al. (2020) BERT 86.9%
-Huang et al. (2020) NB 75.0%
-Huang et al. (2020) SVM 75.0%
-Huang et al. (2020) RF 78.3%
-Huang et al. (2020) CNN 81.9%
-Huang et al. (2020) LSTM 86.0%
-(a) F1-scores (micro average) on ESG classification (Huang et al., 2020)
-Source Model Accuracy
-Lee and Kim (2023) Multiling. BERT 84.09% ±$\pm 0.092$
-Lee and Kim (2023) KoBERT 85.98% ±$\pm 0.163$
-Lee and Kim (2023) DistilKoBERT 84.89% ±$\pm 0.081$
-Lee and Kim (2023) KB-ALBERT 85.62% ±$\pm 0.026$
-Lee and Kim (2023) KLUE (base) 86.05% ±$\pm 0.125$
-Lee and Kim (2023) KLUE (large) 86.66% ±$\pm 0.099$
-(b) Accuracy on ESG classification (Lee and Kim, 2023). Accuracy reported as mean ±$\pm std$
-Environment F1
-Schimanski et al. (2023b) BERT 91.85% ±$\pm 1.25$
-Schimanski et al. (2023b) RoBERTa 92.35% ±$\pm 2.29$
-Schimanski et al. (2023b) DistilRoBERTa 90.97% ±$\pm 2.00$
-Schimanski et al. (2023b) EnvRoBERTa 93.19% ±$\pm 1.40$
-Schimanski et al. (2023b) EnvDistilRoBERTa 92.35% ±$\pm 1.65$
-Social
-Schimanski et al. (2023b) BERT 89.84% ±$\pm 0.79$
-Schimanski et al. (2023b) RoBERTa 89.87% ±$\pm 1.35$
-Schimanski et al. (2023b) DistilRoBERTa 90.59% ±$\pm 1.03$
-Schimanski et al. (2023b) SocRoBERTa 91.90% ±$\pm 1.79$
-Schimanski et al. (2023b) SocDistilRoBERTa 91.24% ±$\pm 1.86$
-Governance
-Schimanski et al. (2023b) BERT 76.86% ±$\pm 1.77$
-Schimanski et al. (2023b) RoBERTa 77.03% ±$\pm 1.82$
-Schimanski et al. (2023b) DistilRoBERTa 76.65% ±$\pm 2.39$
-Schimanski et al. (2023b) GovRoBERTa 78.48% ±$\pm 2.62$
-Schimanski et al. (2023b) GovDistilRoBERTa 78.86% ±$\pm 1.59$
-(c) F1-scores (unspecified) on environment, social, and governance datasets (Schimanski et al., 2024a). F1-scores are reported as mean ±$\pm std$
-Table 14. Reported performances for the datasets on ESG classification
-Source Model Water Forest Biodiv. Nature
-Schimanski et al. (2024a) EnvironmentalBERT 94.47% ±$\pm 1.37$ 95.37% ±$\pm 0.92$ 92.76% ±$\pm 1.01$ 94.19% ±$\pm 0.81$
-Schimanski et al. (2024a) ClimateBERT 95.10% ±$\pm 1.13$ 95.34% ±$\pm 0.94$ 92.49% ±$\pm 1.03$ 93.50% ±$\pm 0.64$
-Schimanski et al. (2024a) RoBERTa 94.55% ±$\pm 0.86$ 94.78% ±$\pm 0.48$ 92.46% ±$\pm 1.54$ 93.97% ±$\pm 0.26$
-Schimanski et al. (2024a) DistilRoBERTa 94.98% ±$\pm 1.16$ 95.29% ±$\pm 0.65$ 92.29% ±$\pm 1.23$ 93.55% ±$\pm 0.72$
-Schimanski et al. (2024a) Keywords - - 63.03% 61.00%
-(a) F1-scores (unspecified) on Nature-related topic classification (Schimanski et al., 2024a). F1-scores are reported as mean ±$\pm std$.
-Source Model F1
-Spokoyny et al. (2023) Majority 0.79%
-Spokoyny et al. (2023) Random 5.05%
-Spokoyny et al. (2023) SVM 48.02%
-Spokoyny et al. (2023) BERT 54.74%
-Spokoyny et al. (2023) RoBERTa 52.90%
-Spokoyny et al. (2023) DistilRoBERTa 51.13%
-Spokoyny et al. (2023) Longformer 54.79%
-Spokoyny et al. (2023) SciBERT 51.83%
-Spokoyny et al. (2023) ClimateBERT 52.97%
-(b) F1-scores (macro average) on SciDCC (Mishra and Mittal, 2021).
-Source Models F1
-Spokoyny et al. (2023) Majority 3.65%
-Spokoyny et al. (2023) Random 6.45%
-Spokoyny et al. (2023) SVM 58.34%
-Spokoyny et al. (2023) BERT 64.64%
-Spokoyny et al. (2023) RoBERTa 65.22%
-Spokoyny et al. (2023) DistilRoBERTa 63.61%
-Spokoyny et al. (2023) Longformer 64.03%
-Spokoyny et al. (2023) SciBERT 63.62%
-Spokoyny et al. (2023) ClimateBERT 64.24%
-(c) F1-scores (macro average) on the climaTOPIC (Spokoyny et al., 2023).
-Source Model F1
-Jain et al. (2023) all-mpnet-base-v2 43.7%
-Jain et al. (2023) TF-IDF 69%
-Jain et al. (2023) Word2Vec 72%
-Jain et al. (2023) RoBERTa 87.19%
-Jain et al. (2023) BERT 87.14%
-Jain et al. (2023) ClimateBERT 85.20%
-(d) F1-scores (weighted average) on ledger transaction classification (Jain et al., 2023).
-Source Model F1
-Vaid et al. (2022) FastText 63.8%
-Vaid et al. (2022) BERT-Base 69.6%
-Vaid et al. (2022) BERT-Large 69.5%
-Vaid et al. (2022) RoBERTa-Base 73.4%
-Vaid et al. (2022) RoBERTa-Large 73.5%
-Vaid et al. (2022) DistilBERT 69.4%
-Spokoyny et al. (2023) Majority 13.83%
-Spokoyny et al. (2023) Random 16.71%
-Spokoyny et al. (2023) SVM 51.81%
-Spokoyny et al. (2023) BERT 71.78%
-Spokoyny et al. (2023) RoBERTa 74.58%
-Spokoyny et al. (2023) DistilRoBERTa 72.33%
-Spokoyny et al. (2023) Longformer 72.28%
-Spokoyny et al. (2023) SciBERT 70.50%
-Spokoyny et al. (2023) ClimateBERT 71.83%
-(e) F1-scores (macro average) on ClimateEng (Vaid et al., 2022).
-Table 15. Reported performances for the datasets on climate-related topic detection (other topics.)
-C.3. In-depth Disclosure: Climate Risk Classification
+sampson_tcfd-nlp_nodate TF-IDF (1) 0.867
+sampson_tcfd-nlp_nodate kNN (2) 0.847
+sampson_tcfd-nlp_nodate FT ClimateBERT (3) 0.852
+sampson_tcfd-nlp_nodate FT distilRoBERTa (4) 0.819
+sampson_tcfd-nlp_nodate (1) + (2) 0.884
+sampson_tcfd-nlp_nodate (1) + (3) 0.865
+sampson_tcfd-nlp_nodate (1) + (4) 0.857
+(b) Performance on TCFD classification task [sampson_tcfd-nlp_nodate] with 11 labels
+Source Model F1 (macro)
+auzepy_evaluating_2023 BART+MNLI 56.68%
+(c) Reported performance of the evaluation of the zero-shot approach for TCFD classification (into 11 categories) [auzepy_evaluating_2023]
+Table 14: Reported performance for the datasets on TFCD classification
+Source Model F1 (micro)
+huangFinBERTLargeLanguage2020 FinBERT 89.6%
+huangFinBERTLargeLanguage2020 BERT 86.9%
+huangFinBERTLargeLanguage2020 NB 75.0%
+huangFinBERTLargeLanguage2020 SVM 75.0%
+huangFinBERTLargeLanguage2020 RF 78.3%
+huangFinBERTLargeLanguage2020 CNN 81.9%
+huangFinBERTLargeLanguage2020 LSTM 86.0%
+(a) Performance for ESG classification [huangFinBERTLargeLanguage2020]
+Source Model Acc Stdev
+LEE2023119726 Multilingual BERT 84.09 % 0.092
+LEE2023119726 KoBERT 85.98 % 0.163
+LEE2023119726 DistilKoBERT 84.89 % 0.081
+LEE2023119726 KB-ALBERT 85.62 % 0.026
+LEE2023119726 KLUE-RoBERTa-base 86.05 % 0.125
+LEE2023119726 KLUE-RoBERTa-large 86.66 % 0.099
+(b) Performance on ESG classification with 4 labels LEE2023119726
+Environment Models F1 (unspecified)
+schimanski_bridging_2023 BERT 91.85% $\pm$ 1.25
+schimanski_bridging_2023 RoBERTa 92.35% $\pm$ 2.29
+schimanski_bridging_2023 DistilRoBERTa 90.97% $\pm$ 2.00
+schimanski_bridging_2023 EnvRoBERTa 93.19% $\pm$ 1.40
+schimanski_bridging_2023 EnvDistilRoBERTa 92.35% $\pm$ 1.65
+Social Models
+schimanski_bridging_2023 BERT 89.84% $\pm$ 0.79
+schimanski_bridging_2023 RoBERTa 89.87% $\pm$ 1.35
+schimanski_bridging_2023 DistilRoBERTa 90.59% $\pm$ 1.03
+schimanski_bridging_2023 SocRoBERTa 91.90% $\pm$ 1.79
+schimanski_bridging_2023 SocDistilRoBERTa 91.24% $\pm$ 1.86
+Governance Models
+schimanski_bridging_2023 BERT 76.86% $\pm$ 1.77
+schimanski_bridging_2023 RoBERTa 77.03% $\pm$ 1.82
+schimanski_bridging_2023 DistilRoBERTa 76.65% $\pm$ 2.39
+schimanski_bridging_2023 GovRoBERTa 78.48% $\pm$ 2.62
+schimanski_bridging_2023 GovDistilRoBERTa 78.86% $\pm$ 1.59
+(c) F1-scores (mean $\pm$ std) for environment, social, and governance models under the standard hyperparameter setup.
+Table 15: Reported performance for the datasets on ESG classification
+Source Model F1 (macro)
+spokoyny2023answering Majority 0.79%
+spokoyny2023answering Random 5.05%
+spokoyny2023answering SVM 48.02%
+spokoyny2023answering BERT 54.74%
+spokoyny2023answering RoBERTa 52.90%
+spokoyny2023answering DistilRoBERTa 51.13%
+spokoyny2023answering Longformer 54.79%
+spokoyny2023answering SciBERT 51.83%
+spokoyny2023answering ClimateBERT 52.97%
+(a) Performance on SciDCC mishra2021neuralnere
+Source Model F1 (weighted)
+jain_supply_2023 all-mpnet-base-v2 43.7%
+jain_supply_2023 TF-IDF 69%
+jain_supply_2023 Word2Vec 72%
+jain_supply_2023 RoBERTa 87.19%
+jain_supply_2023 BERT 87.14%
+jain_supply_2023 ClimateBERT 85.20%
+(b) Ledger classification [jain_supply_2023]
+Source Model Water (%) Forest (%) Biodiversity (%) Nature (%)
+Schimanski2024nature EnvironmentalBERT 94.47 $\pm$ 1.37 95.37 $\pm$ 0.92 92.76 $\pm$ 1.01 94.19 $\pm$ 0.81
+Schimanski2024nature ClimateBERT 95.10 $\pm$ 1.13 95.34 $\pm$ 0.94 92.49 $\pm$ 1.03 93.50 $\pm$ 0.64
+Schimanski2024nature RoBERTa 94.55 $\pm$ 0.86 94.78 $\pm$ 0.48 92.46 $\pm$ 1.54 93.97 $\pm$ 0.26
+Schimanski2024nature DistilRoBERTa 94.98 $\pm$ 1.16 95.29 $\pm$ 0.65 92.29 $\pm$ 1.23 93.55 $\pm$ 0.72
+Schimanski2024nature Keywords - - 63.03 61.00
+(c) Performance (unspecified F1-score) on Nature related topic classification [Schimanski2024nature]
+Source Models F1 (macro)
+spokoyny2023answering Majority 3.65
+spokoyny2023answering Random 6.45
+spokoyny2023answering SVM 58.34
+spokoyny2023answering BERT 64.64
+spokoyny2023answering RoBERTa 65.22
+spokoyny2023answering DistilRoBERTa 63.61
+spokoyny2023answering Longformer 64.03
+spokoyny2023answering SciBERT 63.62
+spokoyny2023answering ClimateBERT 64.24
+(d) Performance on the CLIMA-TOPIC classification [spokoyny2023answering]
+Source Model F1 (macro)
+vaid-etal-2022-towards FastText 63.8%
+vaid-etal-2022-towards BERT-Base 69.6%
+vaid-etal-2022-towards BERT-Large 69.5%
+vaid-etal-2022-towards RoBERTa-Base 73.4%
+vaid-etal-2022-towards RoBERTa-Large 73.5%
+vaid-etal-2022-towards DistilBERT 69.4%
+spokoyny2023answering Majority 13.83%
+spokoyny2023answering Random 16.71%
+spokoyny2023answering SVM 51.81%
+spokoyny2023answering BERT 71.78%
+spokoyny2023answering RoBERTa 74.58%
+spokoyny2023answering DistilRoBERTa 72.33%
+spokoyny2023answering Longformer 72.28%
+spokoyny2023answering SciBERT 70.50%
+spokoyny2023answering ClimateBERT 71.83%
+(e) Results for the Fine-grained Classification using ClimateEng dataset.
+Table 16: Reported performance for the datasets on climate-related topic detection (other topics)
+In-depth Disclosure: Climate Risk Classification
+Source Model F1 (macro)
+xiang_dare_2023 BERT-Base 93.1%
+xiang_dare_2023 BERT-Base (Domain pretrained) 95.5%
+xiang_dare_2023 TinyBERT 87.0%
+xiang_dare_2023 DistilBERT 89.9%
+xiang_dare_2023 ClimateBERT 87.5%
+xiang_dare_2023 CCLA+Max-Pooling 82.9%
+xiang_dare_2023 Bi-LSTM-Attention 71.9%
+xiang_dare_2023 POS-Bi-LSTM-Attention 88.2%
+xiang_dare_2023 DARE 89.4%
+(a) Performance on sentiment analysis (Risk/Opportunity) [xiang_dare_2023]
+Source Model F1 (unspecified)
+kolbel_ask_2021 Bag-of-Words 83.63%
+kolbel_ask_2021 TF-IDF 80.05%
+kolbel_ask_2021 BERT 94.66%
+(b) Performance for risk classification (General/Transition risk/Physical Risk) [kolbel_ask_2021]
 Binary 2 Classes 5 Classes
 source Model REALISTIC REALISTIC REALISTIC HARD NEG. DISCRIMINATORY
-Friederich et al. (2021) SVM 29.0% 35.1% 20.4% 45.7% 59.9%
-Friederich et al. (2021) distilBERT 44.4% 49.7% 24.1% 43.1% 55.8%
-Friederich et al. (2021) RoBERTa 49.6% 44.6% 35.6% 52.8% 59.6%
-(a) F1-scores (macro average) for climate risk disclosure Friederich et al. (2021).
-Source Model F1
-Xiang and Fujii (2023) BERT-Base 93.1%
-Xiang and Fujii (2023) BERT-Base (Domain-spe.) 95.5%
-Xiang and Fujii (2023) TinyBERT 87.0%
-Xiang and Fujii (2023) DistilBERT 89.9%
-Xiang and Fujii (2023) ClimateBERT 87.5%
-Xiang and Fujii (2023) CCLA+Max-Pooling 82.9%
-Xiang and Fujii (2023) Bi-LSTM-Attention 71.9%
-Xiang and Fujii (2023) POS-Bi-LSTM-Attention 88.2%
-Xiang and Fujii (2023) DARE 89.4%
-(b) F1-scores (macro average) on Risk/Opportunity classification (Xiang and Fujii, 2023).
-Source Model F1
-Webersinke et al. (2022) distilRoBERTa 82.5% $\pm$ 4.6
-Webersinke et al. (2022) climateBERT 83.8% $\pm$ 3.6
-Bingler et al. (2023) Naive Bayes 72%
-Bingler et al. (2023) SVM + BoW 72%
-Bingler et al. (2023) SVM + ELMo 75%
-Bingler et al. (2023) climateBERT 80%
-(c) F1-scores (weighted average) on risk/opportunity detection Webersinke et al.
-(2022); Bingler et al.
-(2023).
-Webersinke et al.
-(2022)’s F1-scores are reported as mean ±$\pm std$.
-Bingler et al.
-(2023) also reported an uncertainty measure (p-value) which is not shown in this table for simplicity.
-Source Model F1
-Kölbel et al. (2021) Bag-of-Words 83.63%
-Kölbel et al. (2021) TF-IDF 80.05%
-Kölbel et al. (2021) BERT 94.66%
-(d) F1-scores (unspecified) for risk classification (General/Transition risk/Physical Risk) (Kölbel et al., 2021).
-Table 16. Reported performances for the datasets on risk detection and classification.
-C.4. Green Claim Detection
-Source Model F1
-Stammbach et al. (2023) Majority baseline 00.0%
-Stammbach et al. (2023) Random baseline 33.5%
-Stammbach et al. (2023) ClaimBuster RoBERTa 33.9%
-Stammbach et al. (2023) Pledge Detection RoBERTa 26.4%
-Stammbach et al. (2023) TF-IDF SVM 69.1%
-Stammbach et al. (2023) Character n-gram SVM 70.9%
-Stammbach et al. (2023) DistilBERT 83.7%
-Stammbach et al. (2023) ClimateBERT 83.8%
-Stammbach et al. (2023) RoBERTa_base 82.4%
-Stammbach et al. (2023) RoBERTa_large 84.9%
-(a) F1-scores (binary) on environmental claim detection (Stammbach et al., 2023).
-Source Model F1
-Woloszyn et al. (2021) Bin RoBERTa Bal 92.08%
-Woloszyn et al. (2021) Bin BERTweet Bal 91.48%
-Woloszyn et al. (2021) Bin Flair Bal 88.52%
-Woloszyn et al. (2021) Bin RoBERTa Unbal 88.88%
-Woloszyn et al. (2021) Bin BERTweet Unbal 84.30%
-Woloszyn et al. (2021) Bin Flair Unbal 81.55%
-(b) F1-scores (binary) on green claim detection (Woloszyn et al., 2021).
-Table 17. Reported performances for the datasets on green claim detection.
-C.5. Green Claim Characteristics
-Source Model F1
-Woloszyn et al. (2021) Mult RoBERTa Bal. 81.45%
-Woloszyn et al. (2021) Mult BERTweet Bal. 75.14%
-Woloszyn et al. (2021) Mult Flair Bal. 69.96%
-Woloszyn et al. (2021) Mult RoBERTa Unbal. 76.82%
-Woloszyn et al. (2021) Mult BERTweet Unbal. 59.71%
-Woloszyn et al. (2021) Mult Flair Unbal. 38.15%
-(a) F1-scores (unspecified) on Implicit/Explicit green claim detection (Woloszyn et al., 2021).
-Source Model Accuracy
-Schimanski et al. (2023a) ClimateBERT 96.2% ±$\pm 0.4$
-Schimanski et al. (2023a) DistilRoBERTa 94.4% ±$\pm 0.7$
-Schimanski et al. (2023a) RoBERTa-base 95.8% ±$\pm 0.6$
-Schimanski et al. (2023a) GPT-3.5-turbo 92.0%
-(b) Accuracy on net-zero/reduction classification (Schimanski et al., 2023a). Accuracy is reported as mean ±$\pm std$.
-Source Model Commit. & Act. Specificity
-Bingler et al. (2023) Naive Bayes 75% 75%
-Bingler et al. (2023) SVM + BoW 76% 75%
-Bingler et al. (2023) SVM + ELMo 79% 76%
-Bingler et al. (2023) climateBERT 81% 77%
-(c) F1-scores (weighted average) on Commitment&Action and Specificity classification (Bingler et al., 2023).
-Bingler et al.
-(2023) also reported an uncertainty measure (p-value) which is not reported in this table for simplicity.
-Table 18. Reported performances for the datasets on claim characteristics classification.
-C.6. Green Stance Detection
-Source Model F1
-Vaghefi et al. (2022) GPT-2 67%
-Vaghefi et al. (2022) climateGPT-2 72%
-Claim
-Leippold and Diggelmann (2020) ALBERT(FEVER) 32.85%
-Xiang and Fujii (2023) BERT 80.7%
-Xiang and Fujii (2023) RoBERTa 72.3%
-Xiang and Fujii (2023) DistilRoBERTa 72.0%
-Xiang and Fujii (2023) ClimateBERT 76.8%
-Xiang and Fujii (2023) POS-Bi-LSTM-Att. 79.3%
-Webersinke et al. (2022) DistilRoBERTa 74.8% ±$\pm 3.6$
-Webersinke et al. (2022) ClimateBERT 75.7% ±$\pm 4.4$
-Subset 95 examples
-Wang et al. (2021) RoBERTa (FEVER) 49.2%
-Wang et al. (2021) + ClimateFEVER 64.2%
-Wang et al. (2021) + UDA ratio 18 71.8%
-Evidence
-Spokoyny et al. (2023) Majority 26.08%
-Spokoyny et al. (2023) Random 30.62%
-Spokoyny et al. (2023) BERT 62.47%
-Spokoyny et al. (2023) RoBERTa 60.74%
-Spokoyny et al. (2023) DistilRoBERTa 61.54%
-Spokoyny et al. (2023) Longformer 60.82%
-Spokoyny et al. (2023) SciBERT 62.68%
-Spokoyny et al. (2023) ClimateBERT 61.54%
-(a) F1-score (macro average) on ClimateFEVER Leippold and Diggelmann (2020).
-The prediction can be made at Claim level or for each Claim-Evidence pair (Evidence).
-Webersinke et al.
-(2022)’s F1-scores are reported as mean ±$\pm std$.
-Source Model Acc. F1
-Luo et al. (2020) Majority class 0.43 17%
-Luo et al. (2020) Linear 0.62 60%
-Luo et al. (2020) BERT 0.75 73%
-Luo et al. (2020) Human 0.71 -
-(b) Accuracy and F1-scores (macro average) on GWSD (Luo et al., 2020).
-Source Model F1
-Vaid et al. (2022) FastText 34.3%
-Vaid et al. (2022) BERT-Base 46.4%
-Vaid et al. (2022) BERT-Large 48.9%
-Vaid et al. (2022) RoBERTa-Base 51.0%
-Vaid et al. (2022) RoBERTa-Large 48.9%
-Vaid et al. (2022) DistilBERT 44.8%
-Spokoyny et al. (2023) Majority 29.68%
-Spokoyny et al. (2023) Random 25.52%
-Spokoyny et al. (2023) SVM 42.92%
-Spokoyny et al. (2023) BERT 55.37%
-Spokoyny et al. (2023) RoBERTa 59.69%
-Spokoyny et al. (2023) DistilRoBERTa 52.51%
-Spokoyny et al. (2023) Longformer 34.68%
-Spokoyny et al. (2023) SciBERT 48.67%
-Spokoyny et al. (2023) ClimateBERT 52.84%
-(c) F1-scores (macro average) on ClimateStance (Vaid et al., 2022).
-Source Model macro bin.
-Baselines
-Lai and Chen (2023) BERT - 58%
-Lai and Chen (2023) RoBERTa - 89%
-Lai and Chen (2023) ClimateBERT 90% 90%
-SeqGAN
-Lai and Chen (2023) BERT - 83%
-Lai and Chen (2023) RoBERTa - 87%
-Lai and Chen (2023) ClimateBERT 90% 90%
-MaliGAN
-Lai and Chen (2023) BERT - 74%
-Lai and Chen (2023) RoBERTa - 88%
-Lai and Chen (2023) ClimateBERT 91% 91%
-RankGAN
-Lai and Chen (2023) BERT - 82%
-Lai and Chen (2023) RoBERTa - 87%
-Lai and Chen (2023) ClimateBERT 91% 90%
-(d) F1-scores (binary and macro average) on stance on climate change remediation efforts) (Lai and Chen, 2023).
-Table 19. Reported performance for the datasets on stance detection (Part 1)
+Friederich_climate_risk_disclosure SVM 29.0% 35.1% 20.4% 45.7% 59.9%
+Friederich_climate_risk_disclosure distilBERT 44.4% 49.7% 24.1% 43.1% 55.8%
+Friederich_climate_risk_disclosure RoBERTa 49.6% 44.6% 35.6% 52.8% 59.6%
+(c) Performance (F1 macro average) for climate risk disclosure Tasks from Friederich_climate_risk_disclosure
+Source Model F1 (unspecified)
+huangFinBERTLargeLanguage2020 FinBERT 87.8%
+huangFinBERTLargeLanguage2020 BERT 84.2%
+huangFinBERTLargeLanguage2020 NB 71.1%
+huangFinBERTLargeLanguage2020 SVM 69.6%
+huangFinBERTLargeLanguage2020 RF 66.8%
+huangFinBERTLargeLanguage2020 CNN 72.5%
+huangFinBERTLargeLanguage2020 LSTM 73.3%
+(d) Performance on sentiment Analysis with FinBERT huangFinBERTLargeLanguage2020.
+Source Model F1 (weighted average)
+nicolas_webersinke_climatebert_2021 distilRoBERTa 82.5 % $\pm$ 4.6
+nicolas_webersinke_climatebert_2021 climateBERT 83.8 % $\pm$ 3.6
+bingler2023cheaptalkspecificitysentiment Naive Bayes 72 %
+bingler2023cheaptalkspecificitysentiment SVM + BoW 72 %
+bingler2023cheaptalkspecificitysentiment SVM + ELMo 75 %
+bingler2023cheaptalkspecificitysentiment climateBERT 80 %
+(e) Performance on risk/opportunity detection task from nicolas_webersinke_climatebert_2021 and bingler2023cheaptalkspecificitysentiment.
+bingler2023cheaptalkspecificitysentiment also reported an uncertainty measure (p-value) which is not reported in this table for simplicity.
+Table 17: Reported performance for the datasets on sentiment analysis / risk detection
+Green Claim Detection
+Source Model F1 (binary)
+stammbach_environmental_2023 Majority baseline 00.0%
+stammbach_environmental_2023 Random baseline 33.5%
+stammbach_environmental_2023 ClaimBuster RoBERTa 33.9%
+stammbach_environmental_2023 Pledge Detection RoBERTa 26.4%
+stammbach_environmental_2023 TF-IDF SVM 69.1%
+stammbach_environmental_2023 Character n-gram SVM 70.9%
+stammbach_environmental_2023 DistilBERT 83.7%
+stammbach_environmental_2023 ClimateBERT 83.8%
+stammbach_environmental_2023 RoBERTa_base 82.4%
+stammbach_environmental_2023 RoBERTa_large 84.9%
+(a) Performance on environmental claim detection from stammbach_environmental_2023
+Source Model F1 (binary)
+vinicius_woloszyn_towards_2021 Bin RoBERTa Bal 92.08%
+vinicius_woloszyn_towards_2021 Bin BERTweet Bal 91.48%
+vinicius_woloszyn_towards_2021 Bin Flair Bal 88.52%
+vinicius_woloszyn_towards_2021 Bin RoBERTa Unbal 88.88%
+vinicius_woloszyn_towards_2021 Bin BERTweet Unbal 84.30%
+vinicius_woloszyn_towards_2021 Bin Flair Unbal 81.55%
+(b) Performance on green claim detection from vinicius_woloszyn_towards_2021
+Table 18: Reported performance for the datasets on claim detection
+Green Claim Characteristics
+Source Model Commitments & Actions Specificity
+F1 (weighted average) F1 (weighted average)
+bingler2023cheaptalkspecificitysentiment Naive Bayes 75 % 75 %
+bingler2023cheaptalkspecificitysentiment SVM + BoW 76 % 75 %
+bingler2023cheaptalkspecificitysentiment SVM + ELMo 79 % * 76 %
+bingler2023cheaptalkspecificitysentiment climateBERT 81 % 77 %
+(a) Performance on downstream task from bingler2023cheaptalkspecificitysentiment.
+citetbingler2023cheaptalkspecificitysentiment also reported an uncertainty measure (p-value) which is not reported in this table for simplicity.
+Source Model F1(unspecified)
+vinicius_woloszyn_towards_2021 Mult RoBERTa Bal 81.45 %
+vinicius_woloszyn_towards_2021 Mult BERTweet Bal 75.14 %
+vinicius_woloszyn_towards_2021 Mult Flair Bal 69.96 %
+vinicius_woloszyn_towards_2021 Mult RoBERTa Unbal 76.82 %
+vinicius_woloszyn_towards_2021 Mult BERTweet Unbal 59.71 %
+vinicius_woloszyn_towards_2021 Mult Flair Unbal 38.15 %
+(b) Performance on Implicit/Explicit green claim detection from vinicius_woloszyn_towards_2021
+Source Model Accuracy Std
+tobias_schimanski_climatebert-netzero_2023 ClimateBERT 96.2 % 0.4
+tobias_schimanski_climatebert-netzero_2023 DistilRoBERTa 94.4 % 0.7
+tobias_schimanski_climatebert-netzero_2023 RoBERTa-base 95.8 % 0.6
+tobias_schimanski_climatebert-netzero_2023 GPT-3.5-turbo 92.0 % -
+(c) Performances on net-zero/reduction detection task from tobias_schimanski_climatebert-netzero_2023
+Table 19: Reported performance for the datasets on claim characteristics classification tasks
+Green Stance Detection
+Source Model ClimateFEVER
+vaghefi2022deep GPT-2 67 %
+vaghefi2022deep climateGPT-2 72 %
+Claim F1 (macro)
+diggelmann_climate-fever_2020 ALBERT(FEVER) 32.85%
+xiang_dare_2023 BERT 80.7%
+xiang_dare_2023 RoBERTa 72.3%
+xiang_dare_2023 DistilRoBERTa 72.0%
+xiang_dare_2023 ClimateBERT 76.8%
+xiang_dare_2023 POS-Bi-LSTM-Attention 79.3%
+nicolas_webersinke_climatebert_2021 DistilRoBERTa 74.8% $\pm$ 3.6
+nicolas_webersinke_climatebert_2021 ClimateBERT 75.7% $\pm$ 4.4
+subset 95 examples F1 (macro)
+Wang2021EvidenceBA RoBERTa Large (FEVER) 49.2%
+Wang2021EvidenceBA + CLIMATE-FEVER 64.2%
+Wang2021EvidenceBA + UDA ratio 8 71.7%
+Wang2021EvidenceBA + UDA ratio 18 71.8%
+Evidence F1 (macro)
+spokoyny2023answering Majority 26.08%
+spokoyny2023answering Random 30.62%
+spokoyny2023answering BERT 62.47%
+spokoyny2023answering RoBERTa 60.74%
+spokoyny2023answering DistilRoBERTa 61.54%
+spokoyny2023answering Longformer 60.82%
+spokoyny2023answering SciBERT 62.68%
+spokoyny2023answering ClimateBERT 61.54%
+(a) Performance on ClimateFEVER diggelmann_climate-fever_2020.
+The dataset contains pairs of claim and evidences, and an aggregated label for the claim.
+The evaluation is computed at the claim level (aggregated).
+Except for spokoyny2023answering, they computed the score at the evidence level (for each pair).
+$\pm$ value reports the standard deviation.
+Source Model Acc F1 (macro)
+luo_detecting_2020 Majority class 0.43 17%
+luo_detecting_2020 Linear 0.62 60%
+luo_detecting_2020 BERT 0.75 73%
+luo_detecting_2020 Human 0.71 -
+(b) Performance on GWSD[luo_detecting_2020] (stance detection on global warming)
+Source Model F1 (macro)
+vaid-etal-2022-towards FastText 34.3%
+vaid-etal-2022-towards BERT-Base 46.4%
+vaid-etal-2022-towards BERT-Large 48.9%
+vaid-etal-2022-towards RoBERTa-Base 51.0%
+vaid-etal-2022-towards RoBERTa-Large 48.9%
+vaid-etal-2022-towards DistilBERT 44.8%
+spokoyny2023answering Majority 29.68%
+spokoyny2023answering Random 25.52%
+spokoyny2023answering SVM 42.92%
+spokoyny2023answering BERT 55.37%
+spokoyny2023answering RoBERTa 59.69%
+spokoyny2023answering DistilRoBERTa 52.51%
+spokoyny2023answering Longformer 34.68%
+spokoyny2023answering SciBERT 48.67%
+spokoyny2023answering ClimateBERT 52.84%
+(c) Performance on ClimateStance, the stance classification task of vaid-etal-2022-towards (3 labels: Favor, Against, Ambiguous)
+Model F1 (macro) F1 (binary)
+Baseline Models
+BERT - 58%
+RoBERTa - 89%
+ClimateBERT 90% 90%
+SeqGAN Approach
+BERT - 83%
+RoBERTa - 87%
+ClimateBERT 90% 90%
+MaliGAN Approach
+BERT - 74%
+RoBERTa - 88%
+ClimateBERT 91% 91%
+RankGAN Approach
+BERT - 82%
+RoBERTa - 87%
+ClimateBERT 91% 90%
+(d) Performance on stance on climate change remediation efforts (support/attack) [lai_using_2023]
 Document Page overlap Strict
 Source Model P Q S P Q S P Q S
-Morio and Manning (2023) Most-frequent 46.7% 52.6% 36.8% 51.8% 25.6% 19.8% 41.2% 19.6% 17.5%
-Morio and Manning (2023) Linear 66.0% 61.9% 50.3% 71.4% 44.5% 36.1% 52.0% 31.2% 27.0%
-Morio and Manning (2023) BERT-base 71.0% 63.5% 51.6% 73.6% 48.1% 37.2% 50.2% 31.9% 25.8%
-Morio and Manning (2023) ClimateBERT 71.8% 64.0% 52.8% 74.4% 48.9% 39.0% 50.2% 32.2% 26.8%
-Morio and Manning (2023) RoBERTa-base 71.6% 64.5% 53.1% 73.8% 49.6% 38.3% 50.4% 33.4% 26.6%
-Morio and Manning (2023) Longformer-base 73.7% 66.9% 54.6% 75.9% 53.0% 40.8% 52.5% 36.1% 28.6%
-Morio and Manning (2023) Longformer-large 73.9% 68.8% 57.3% 76.5% 55.0% 44.1% 53.6% 38.7% 31.5%
-(a) Custom metrics on LobbyMap (Morio and Manning, 2023).
-Table 20. Reported performance for the datasets on stance detection (Part 2)
-C.7. Question Answering
-Source embeddings Quest. Def. Concepts Gen. inf_3 inf_all
-Schimanski et al. (2024b) random 0.037 0.037 0.037 0.037 0.037 0.037
-Schimanski et al. (2024b) BM25 0.113 0.114 0.126 0.139 0.172 0.174
-Schimanski et al. (2024b) ColBERTv2 0.109 0.094 0.112 0.137 0.130 0.137
-Schimanski et al. (2024b) DRAGON+ 0.139 0.121 0.161 0.160 0.161 0.160
-Schimanski et al. (2024b) GTE-base 0.161 0.153 0.171 0.154 0.171 0.174
-Schimanski et al. (2024b) text-embedding-ada-002 0.163 0.143 0.152 0.163 0.176 0.176
-Schimanski et al. (2024b) text-embedding-3-small 0.163 0.143 0.152 0.163 0.179 0.179
-Schimanski et al. (2024b) text-embedding-3-large 0.167 0.143 0.152 0.163 0.179 0.179
-(a) F1-score (custom) of the different embedding models and information retrieval approaches (question, definition, concepts, generic, inf_3, inf_all) (Schimanski et al., 2024b)
-Test In-house annotation
-Source Model F1 BLEU METEOR F1 BLEU METEOR
-Rony et al. (2022) ALBERT 81.6% 0.678 0.808 66.1% 0.416 0.694
-(b) F1-score, BLEU and METEOR on CCMRC (Rony et al., 2022)
-Source Model F1
-Luccioni et al. (2020) large 85.5%
-Luccioni et al. (2020) base 82%
-(c) F1-scores (macro average) of RoBERTa on question answering (Luccioni et al., 2020).
-Source Model F1
-Spokoyny et al. (2023) Majority 4.11%
-Spokoyny et al. (2023) Random 12.14%
-Spokoyny et al. (2023) SVM 86.00%
-Spokoyny et al. (2023) BERT 84.57%
-Spokoyny et al. (2023) RoBERTa 85.61%
-Spokoyny et al. (2023) DistilRoBERTa 84.38%
-Spokoyny et al. (2023) Longformer 84.35%
-Spokoyny et al. (2023) SciBERT 84.43%
-Spokoyny et al. (2023) ClimateBERT 84.80%
-(d) F1-scores (macro average) on ClimaINS Spokoyny et al. (2023)
-Source Model Question Answer
-R@1 R@5 R@1 R@5
-Thulke et al. (2024) bge-base-en-v1.5 54.8 71.5 81.8 92.1
-Thulke et al. (2024) bge-large-en-v1.5 55.8 73.6 83.3 93.1
-Thulke et al. (2024) gtr-t5-large 48.8 67.4 79.6 90.1
-Thulke et al. (2024) gtr-t5-xxl 47.6 66.3 79.2 89.7
-Thulke et al. (2024) gte-large 50.7 68.2 80.9 91.4
-Thulke et al. (2024) ember-v1 49.5 68.6 79.7 91.1
-Thulke et al. (2024) instructor-large 50.0 68.2 81.7 91.8
-Thulke et al. (2024) instructor-xl 53.3 69.7 83.3 92.1
-(e) Recall (R@1 and R@5) of retrieving the correct document given the question or the answer from the IFT dataset (Thulke et al., 2024).
-Model CITIES STATES CORP
+morio2023an Most-frequent 46.7% 52.6% 36.8% 51.8% 25.6% 19.8% 41.2% 19.6% 17.5%
+morio2023an Linear 66.0 % 61.9 % 50.3 % 71.4 % 44.5 % 36.1 % 52.0 % 31.2 % 27.0 %
+morio2023an BERT-base 71.0 % 63.5 % 51.6 % 73.6 % 48.1 % 37.2 % 50.2 % 31.9 % 25.8 %
+morio2023an ClimateBERT 71.8 % 64.0 % 52.8 % 74.4 % 48.9 % 39.0 % 50.2 % 32.2 % 26.8 %
+morio2023an RoBERTa-base 71.6 % 64.5 % 53.1 % 73.8 % 49.6 % 38.3 % 50.4 % 33.4 % 26.6 %
+morio2023an Longformer-base 73.7 % 66.9 % 54.6 % 75.9 % 53.0 % 40.8 % 52.5 % 36.1 % 28.6 %
+morio2023an Longformer-large 73.9 % 68.8 % 57.3 % 76.5 % 55.0 % 44.1 % 53.6 % 38.7 % 31.5 %
+(e) Performance on LobbyMap [morio2023an] using custom metrics described in the study.
+Table 20: Reported performance for the datasets on stance detection
+Question Answering
+embeddings question definition concepts generic inf_3 inf_all
+random 0.037 0.037 0.037 0.037 0.037 0.037
+BM25 0.113 0.114 0.126 0.139 0.172 0.174
+ColBERTv2 0.109 0.094 0.112 0.137 0.130 0.137
+DRAGON+ 0.139 0.121 0.161 0.160 0.161 0.160
+GTE-base 0.161 0.153 0.171 0.154 0.171 0.174
+text-embedding-ada-002 0.163 0.143 0.152 0.163 0.176 0.176
+text-embedding-3-small 0.163 0.143 0.152 0.163 0.179 0.179
+text-embedding-3-large 0.167 0.143 0.152 0.163 0.179 0.179
+(a) Results for the F1-score (custom) of the different embedding models and information retrieval approaches (question, definition, concepts, generic, inf_3, inf_all) aggregated across all top-k values (5, 10, 15).
+[schimanski-etal-2024-climretrieve]
+Source Model Test In-house annotation test
+F1-score BLEU METEOR F1-score BLEU METEOR
+cliamtebot_2022 ALBERT 81.6 % 0.678 0.808 66.1 % 0.416 0.694
+(b) Performance on CCMRC cliamtebot_2022
+Task Data Accuracy (% of data)
+net zero target year raw/optimal 0.95 (1.0)
+confidence-tuned 0.97 (0.5)
+reduction target year raw 0.84 (1.0)
+optimal 0.90 (0.97)
+confidence-tuned 0.90 (0.8)
+reduction base year raw 0.68 (1.0)
+optimal 0.80 (0.84)
+confidence-tuned 0.82 (0.62)
+reduction percentage raw 0.88 (1.0)
+optimal 0.91 (0.96)
+confidence-tuned 0.93 (0.35)
+(c) Accuracy of the Q&A approach in assessing the quantitative properties of the net zero and reduction targets [tobias_schimanski_climatebert-netzero_2023]
+Source Model F1 (macro)
+spokoyny2023answering Majority 4.11 %
+spokoyny2023answering Random 12.14 %
+spokoyny2023answering SVM 86.00 %
+spokoyny2023answering BERT 84.57 %
+spokoyny2023answering RoBERTa 85.61 %
+spokoyny2023answering DistilRoBERTa 84.38 %
+spokoyny2023answering Longformer 84.35 %
+spokoyny2023answering SciBERT 84.43 %
+spokoyny2023answering ClimateBERT 84.80 %
+(d) Performance on ClimaINS spokoyny2023answering
+CDP-CITIES CDP-STATES CDP-CORP
+Model MRR@10 MRR@10 MRR@10
 No Finetuning on CDP
@@
 MiniLM 0.099 0.120 0.320
-Finetuned on CDP (In-Domain)
+Finetuned on CDP
+In-Domain In-Domain In-Domain
 ClimateBERT 0.331 0.422 0.753
@@
 MiniLM 0.352 0.489 0.745
-(f) MRR@10 scores on the three subsets of CLIMA-QA (Spokoyny et al., 2023). Models finetuned and evaluated on the same subset fall under In-Domain.
-Task Raw Optimal Confidence
-Net-zero target year 0.95 0.95 0.97
-Reduction target year 0.84 0.90 0.90
-Reduction base year 0.68 0.80 0.82
-Reduction percentage 0.88 0.91 0.93
-(g) Accuracy of the Q&A approach in assessing the quantitative properties of the net zero and reduction targets (Schimanski et al., 2023a)
-Table 21. Reported performances for the datasets on Question/Answering and Retrieval tasks.
+(e) MRR@10 scores for BM25, ClimateBERT and MSMARCO-MiniLM on the three subsets of CLIMA-QA. Models finetuned and evaluated on same subset fall under In-Domain.
+Source Model F1-score (macro)
+luccioni_analyzing_2020 RoBERTa-large 85.5 %
+luccioni_analyzing_2020 RoBERTa-base 82 %
+(f) Performance on question answering [luccioni_analyzing_2020] (2 labels: answer, not) mapping a sentence to one of the TCFD questions
+Table 21: Reported performance for the datasets on Question/Answering (Retrieval) Tasks
+Source Model Accuracy
+s_vaghefi_chatclimate_2023 hybrid ChatClimate 4.38
+s_vaghefi_chatclimate_2023 ChatClimate 4.15
+s_vaghefi_chatclimate_2023 GPT-4 2.62
+(a) ”Accuracy” on a set of 10 questions, human annotations s_vaghefi_chatclimate_2023
+Source Model F1 (macro)
+spokoyny2023answering Majority 20.10%
+spokoyny2023answering Random 24.09%
+spokoyny2023answering SVM - %
+spokoyny2023answering BERT 70.57%
+spokoyny2023answering RoBERTa 71.14%
+spokoyny2023answering DistilRoBERTa 69.27%
+spokoyny2023answering Longformer 67.72%
+spokoyny2023answering SciBERT 68.45%
+spokoyny2023answering ClimateBERT 69.44%
+(b) Performance on ClimaBench spokoyny2023answering
 Source Model ClimaBench Pira 2.0 MCQ Exeter Misinf. Weight. Avg.
-Acc. Acc. F1 (macro)
-Thulke et al. (2024) Stability-3B 71.4% 48.7% 52.6% 62.8%
-Thulke et al. (2024) Pythia-6.9B 63.6% 22.9% 29.9% 48.9%
-Thulke et al. (2024) Falcon-7B 62.9% 19.8% 39.9% 48.3%
-Thulke et al. (2024) Mistral-7B 73.1% 19.0% 63.7% 53.7%
-Thulke et al. (2024) Llama-2-7B 68.5% 51.1% 59.4% 62.6%
-Thulke et al. (2024) Jais-13B 66.9% 26.4% 54.2% 54.4%
-Thulke et al. (2024) Jais-13B-Chat 65.8% 66.4% 61.3% 64.6%
-Thulke et al. (2024) Llama-2-Chat-7B 67.8% 72.0% 64.3% 68.5%
-Thulke et al. (2024) Llama-2-Chat-13B 68.6% 79.3% 68.6% 71.4%
-Thulke et al. (2024) Llama-2-Chat-70B 72.7% 88.2% 72.5% 77.8%
-Thulke et al. (2024) ClimateGPT-7B 75.3% 86.6% 65.9% 77.1%
-Thulke et al. (2024) ClimateGPT-13B 75.0% 89.0% 70.0% 78.0%
-Thulke et al. (2024) ClimateGPT-70B 72.4% 89.9% 72.5% 77.7%
-Thulke et al. (2024) ClimateGPT-FSC-7B 59.3% 17.2% 45.1% 46.2%
-Thulke et al. (2024) ClimateGPT-FSG-7B 53.1% 17.4% 41.5% 42.1%
-(a) F1 scores (macro average) and Accuracy on multiple climate benchmarks (Thulke et al., 2024).
-Source Model F1
-Spokoyny et al. (2023) Majority 20.10%
-Spokoyny et al. (2023) Random 24.09%
-Spokoyny et al. (2023) SVM -%
-Spokoyny et al. (2023) BERT 70.57%
-Spokoyny et al. (2023) RoBERTa 71.14%
-Spokoyny et al. (2023) DistilRoBERTa 69.27%
-Spokoyny et al. (2023) Longformer 67.72%
-Spokoyny et al. (2023) SciBERT 68.45%
-Spokoyny et al. (2023) ClimateBERT 69.44%
-(b) F1-scores (macro average) on ClimaBench (Spokoyny et al., 2023)
-Source Model Acc.
-Vaghefi et al. (2023) hybrid ChatClimate 4.38
-Vaghefi et al. (2023) ChatClimate 4.15
-Vaghefi et al. (2023) GPT-4 2.62
-(c) Accuracy on a set of 10 questions, human annotations (Vaghefi et al., 2023).
-Table 22. Reported performance for the datasets on Question/Answering (few-shot and RAG) Tasks.
-C.8. Deceptive Technique
-Source Model F1
-Trained on Logic applied on LogicCLimate
-Jin et al. (2022) Electra-SA 22.72%
-Jin et al. (2022) Electra 27.23%
+F1 (macro) F1 (unspecified) F1 (macro)
+thulke2024climategpt Stability-3B 71.4 48.7 52.6 62.8
+thulke2024climategpt Pythia-6.9B 63.6 22.9 29.9 48.9
+thulke2024climategpt Falcon-7B 62.9 19.8 39.9 48.3
+thulke2024climategpt Mistral-7B 73.1 19.0 63.7 53.7
+thulke2024climategpt Llama-2-7B 68.5 51.1 59.4 62.6
+thulke2024climategpt Jais-13B 66.9 26.4 54.2 54.4
+thulke2024climategpt Jais-13B-Chat 65.8 66.4 61.3 64.6
+thulke2024climategpt Llama-2-Chat-7B 67.8 72.0 64.3 68.5
+thulke2024climategpt Llama-2-Chat-13B 68.6 79.3 68.6 71.4
+thulke2024climategpt Llama-2-Chat-70B 72.7 88.2 72.5 77.8
+thulke2024climategpt ClimateGPT-7B 75.3 86.6 65.9 77.1
+thulke2024climategpt ClimateGPT-13B 75.0 89.0 70.0 78.0
+thulke2024climategpt ClimateGPT-70B 72.4 89.9 72.5 77.7
+thulke2024climategpt ClimateGPT-FSC-7B 59.3 17.2 45.1 46.2
+thulke2024climategpt ClimateGPT-FSG-7B 53.1 17.4 41.5 42.1
+(c) Results on climate benchmarks[thulke2024climategpt]
+Table 22: Reported performance for the datasets on Question/Answering (few-shot and RAG) Tasks.
+Deceptive Technique
+Source Model F1 (micro)
+Trained on general domain
+jin-etal-2022-logical Electra-StructureAware 22.72 %
+jin-etal-2022-logical Electra 27.23 %
 Trained on LogicClimate
-Jin et al. (2022) Electra-SA 23.71%
-Jin et al. (2022) Electra 29.37%
-Baseline: Trained on Logic applied on Logic
-Jin et al. (2022) Electra-SA 58.77%
-Jin et al. (2022) Electra 53.31%
-Jin et al. (2022) BERT 45.80%
-(a) F1-scores (micro average) on fallacy detection (Jin et al., 2022). Electra-StructureAware is abbreviated Electra-SA
-Source Model F1
-Bhatia et al. (2021) BERT 59%
-Bhatia et al. (2021) BERT* 62%
-Bhatia et al. (2021) MTEXT 66%
-Bhatia et al. (2021) MTEXT* 67%
-Bhatia et al. (2021) MTEXTmulti 68%
-Bhatia et al. (2021) MTEXT*multi 67%
-Bhatia et al. (2021) SVM 49%
-Bhatia et al. (2021) Human 70%
-(b) F1-score (micro average) on neutralization techniques classification (Bhatia et al., 2021)
-Source Model F1 (macro)
-Coan et al. (2021) Logistic (unweighted) 68%
-Coan et al. (2021) Logistic (weighted) 72%
-Coan et al. (2021) SVM (unweighted) 66%
-Coan et al. (2021) SVM (weighted) 72%
-Coan et al. (2021) ULMFiT 72%
-Coan et al. (2021) ULMFiT (weighted) 65%
-Coan et al. (2021) ULMFiT (over sample) 55%
-Coan et al. (2021) ULMFiT (focal Loss) 61%
-Coan et al. (2021) ULMFiT-logistic 75%
-Coan et al. (2021) ULMFiT-SVM 71%
-Coan et al. (2021) RoBERTa 77%
-Coan et al. (2021) RoBERTa-logistic 79%
-Coan et al. (2021) Coder (Human) blAcc (macro): 87%
-Thulke et al. (2024) Llama-2-Chat-70Bbla F1 (Bin.): 72.5%
-Thulke et al. (2024) ClimateGPT-70Bblah F1 (Bin.): 72.5%
-(c) F1-score (macro average and Binary) and Accuracy (macro average) on contrarian claim classification (Coan et al., 2021).
-Table 23. Deceptive techniques
-C.9. Environmental Performance Prediction
+jin-etal-2022-logical Electra-StructureAware 23.71 %
+jin-etal-2022-logical Electra 29.37 %
+Baseline: Performance on Logic
+jin-etal-2022-logical Electra-StructureAware 58.77 %
+jin-etal-2022-logical Electra 53.31 %
+jin-etal-2022-logical BERT 45.80 %
+(a) Performances on jin-etal-2022-logical fallacy detection (13 labels)
+Source Model F1 (micro)
+bhatia_automatic_2021-1 BERT 59 %
+bhatia_automatic_2021-1 BERT* 62 %
+bhatia_automatic_2021-1 MTEXT 66 %
+bhatia_automatic_2021-1 MTEXT* 67 %
+bhatia_automatic_2021-1 MTEXTmulti 68 %
+bhatia_automatic_2021-1 MTEXT*multi 67 %
+bhatia_automatic_2021-1 SVM 49 %
+bhatia_automatic_2021-1 Human 70 %
+(b) Performance on bhatia_automatic_2021-1
+Source Model F1-score (macro)
+coanComputerassistedClassificationContrarian2021 Logistic (unweighted) 68 %
+coanComputerassistedClassificationContrarian2021 Logistic (weighted) 72 %
+coanComputerassistedClassificationContrarian2021 SVM (unweighted) 66 %
+coanComputerassistedClassificationContrarian2021 SVM (weighted) 72 %
+coanComputerassistedClassificationContrarian2021 ULMFiT 72 %
+coanComputerassistedClassificationContrarian2021 ULMFiT (weighted) 65 %
+coanComputerassistedClassificationContrarian2021 ULMFiT (over sample) 55 %
+coanComputerassistedClassificationContrarian2021 ULMFiT (focal Loss) 61 %
+coanComputerassistedClassificationContrarian2021 ULMFiT-logistic 75 %
+coanComputerassistedClassificationContrarian2021 ULMFiT-SVM 71 %
+coanComputerassistedClassificationContrarian2021 RoBERTa 77 %
+coanComputerassistedClassificationContrarian2021 RoBERTa-logistic 79 %
+Accuracy (macro)
+coanComputerassistedClassificationContrarian2021 Coder (Human) 87 %
+Binary Classification
+[thulke2024climategpt] Llama-2-Chat-70B 72.5 %
+[thulke2024climategpt] ClimateGPT-70B 72.5 %
+(c) Performance on contrarian claim classification [coanComputerassistedClassificationContrarian2021]
+Table 23: Deceptive techniques
+Environmental Performance Prediction
 Source Model MSE
-Lin et al. (2023b) Baseline 1.173
-Lin et al. (2023b) Lasso 0.848
-Lin et al. (2023b) Gradient Boosting 0.818
-Lin et al. (2023b) NOCATE(RB) 0.776
-Lin et al. (2023b) NOCATE(CB) 0.763
-Lin et al. (2023b) NOCATE(DB) 0.756
-Lin et al. (2023b) SUSTAINABLESIGNALS(RB) 0.762
-Lin et al. (2023b) SUSTAINABLESIGNALS(CB) 0.753
-Lin et al. (2023b) SUSTAINABLESIGNALS(DB) 0.736
-(a) MSE on Finch Score prediction (Finch Score range between 0 and 10) (Lin et al., 2023b)
-Source Model F1
-Clarkson et al. (2020) Random Forest 89.48%
-Clarkson et al. (2020) XGBoost 95.73%
-(b) F1-score (macro average) on good/bad CSR performer classification (Clarkson et al., 2020).
+linSUSTAINABLESIGNALSijcai2023 Baseline 1.173
+linSUSTAINABLESIGNALSijcai2023 Lasso 0.848
+linSUSTAINABLESIGNALSijcai2023 Gradient Boosting 0.818
+linSUSTAINABLESIGNALSijcai2023 NOCATE(RB) 0.776
+linSUSTAINABLESIGNALSijcai2023 NOCATE(CB) 0.763
+linSUSTAINABLESIGNALSijcai2023 NOCATE(DB) 0.756
+linSUSTAINABLESIGNALSijcai2023 SUSTAINABLESIGNALS(RB) 0.762
+linSUSTAINABLESIGNALSijcai2023 SUSTAINABLESIGNALS(CB) 0.753
+linSUSTAINABLESIGNALSijcai2023 SUSTAINABLESIGNALS(DB) 0.736
+(a) Performance on the task of Finch Score prediction (Finch Score range between 0 and 10) linSUSTAINABLESIGNALSijcai2023
 Source Model RMSE wMAPE
-Bronzini et al. (2023) OLS 7.76 7.9%
+bronzini_glitter_2023 OLS 7.76 7.9 %
 Uniform Random 40 66 %
 Normal Average 20 31 %
-(c) RMSE and wMAPE on ESG score prediction (ESG score range between 0 and 100) (Bronzini et al., 2023).
-We added the average RMSE and wMAPE for a uniform distribution of score with a random classifier; and for a normal distribution (mean=50, scale=20) with the average value as prediction.
+(b) Performance on ESG score prediction (ESG score range between 0 and 100) bronzini_glitter_2023.
+We added the average RMSE and wMAPE for a uniform distribution of score with a random classifier; and for a normal distribution (mean=50, scale=20) with the average value a prediction.
+Source Model F1 (macro)
+clarkson_nlp_us_csr Random Forest 89.48 %
+clarkson_nlp_us_csr XGBoost 95.73 %
+(c) clarkson_nlp_us_csr (2 classes: good performer, bad performer)
 Source Model MAE RMSE
-Sharma et al. (2024) CLIP+MLP 0.54 0.88
-Sharma et al. (2024) Sent.-BERT+MLP 0.57 0.91
+Greenscreen CLIP+MLP 0.54 0.88
+Greenscreen Sentence-BERT+MLP 0.57 0.91
 Uniform Random 3.97 4.88
 Gaussian Average 1.59 1.99
-(d) MAE and RMSE on ESG unmanaged risk score prediction (ESG risk between 0 and 13.5) (Sharma et al., 2024).
+(d) Performance on ESG unmanaged risk score prediction from Greenscreen (ESG risk between 0 and 13.5).
 We added the average RMSE and MAE for a uniform distribution of score with a random classifier; and for a normal distribution (mean=7, scale=2) with the average value a prediction.
 Source Model Accuracy (Change) Accuracy (Direction of change)
-Mehra et al. (2022) majority class 0.5791 0.5682
-Mehra et al. (2022) BERT 0.5985 0.4317
-Mehra et al. (2022) ESGBERT 0.6709 0.793
-(e) Accuracy on ESG-risk change classification (Mehra et al., 2022)
-Table 24. Environmental Performance Prediction
+Mehra_2022 majority class 0.5791 0.5682
+Mehra_2022 BERT 0.5985 0.4317
+Mehra_2022 ESGBERT 0.6709 0.793
+(e) Performance on Mehra_2022’s tasks of ESG risk change classification
+Table 24: Environmental Performance Prediction